### v1

# Linying Lv

Abstract
I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports. Using embeddings from state-of-the-art large language models, I show that textual information in analyst reports explains 10.19% of contemporaneous stock returns out-of-sample, a value that is economically more significant than quantitative forecasts. I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns. The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power. Expressing these findings in economic terms, I estimate that early acquisition of analysts’ reports can yield significant profits. Analysts’ information value peeks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data. JEL Classification: G11, G14, G24
Keywords: Sell-side Analysts, Value of Information, Large Language Model, Explainable AI
“No one ever made a decision because of a number. They need a story.”
― Daniel Kahneman
1 Introduction
Sell-side analysts play a crucial role in financial markets by processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. While prior studies have demonstrated the incremental information value of report text, these analyses have predominantly centered on the sentiment dimension (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information perceived as at least as important as the quantifiable outputs. For example, according to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is currently limited, and it remains unclear which specific component of analyst reports holds the most value for investors.
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning, capturing more granular and rich information than before. This advancement facilitates a systematic analysis of the comprehensive outputs from sell-side analysts. For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the ”earnings analysis” topic, the second provides additional contextual information that LLMs can interpret and analyze.
I quantify the value of information in analyst reports in three general steps: analyst outputs representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices. For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs. Each report is mapped into a structured embedding space using the LLMs’ transformer architecture, allowing these to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2023).
I evaluate the information content of analyst reports using an R-squared measure to quantify how well the information set explains stock returns. A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings. To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients. This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns. Additionally, I assess the explanatory power of both quantitative and qualitative measures for abnormal stock returns using out-of-sample R-squared values with a recursively expanding window.
I compile a comprehensive corpus of analyst reports from the Mergent Investext database, merging these written reports with detailed analyst forecast summaries from the I/B/E/S database. The final sample comprises 122,252 analyst reports on S&P 100 constituent firms spanning from 2000 to 2023. Using this sample, the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually. Economically, a one standard deviation increase in report text favorableness raises the realized CAR by 120 basis points, compared to 100 basis points for forecast revisions. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content winthin analyst report text.
The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations. My analysis reveals that both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions. Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%. In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
Since a notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al. (2014), Kim and Song (2015), Lobo et al. (2017), and Barron et al. (2017), I examine the information content of analyst reports both within and beyond the earnings announcement window. This investigation reveals that the information content of analyst reports peaks in the first week following earnings announcements, supporting Huang et al. (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients. Using separately trained models on samples within and beyond earnings announcement windows, I find that the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods. This significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
To address potential concerns that analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models. By doing so, I can isolate the unique information of analyst reports. The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights, highlighting the different information sets provided by the two sources.
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question. To quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers. This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
To translate the statistic measure of R-squared into economic terms, I quantify the dollar value of analyst information to strategic investors using Kadan and Manela (forthcoming)’s framework, which builds on Back et al. (2000) imperfect competition insider trading model. The information value can be interpreted as the total expected profits for strategic investors either because some investors have early access to analyst views or because they conduct independent research and derive insights comparable to those of analysts. This multi-client framework provides a realistic foundation for analyzing the complex dynamics of information dissemination in analyst reports. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
The findings reveal significant value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports are \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million from both combined. The annualized aggregate information value is estimated at \$6.89 million, based on an average of 15 days with reports released per year. According to Kadan and Manela (forthcoming), this estimate represents a conservative lower bound. Notably, these values increase for large stocks, bold analysts, and reports released promptly following earnings announcements.
This paper relates to three strands of literature. First, This paper relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights. In contrast, my work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations. A direct comparison with the Naive Bayes tone classification used in Huang et al. (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone. To quantify these gains, I employ a BERT-based tone measure, which increases the out-of-sample $R^{2}$ significantly by 2%. Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure to above 10%. The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information, offering a more nuanced and comprehensive understanding of analyst narratives.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2023), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in financial contexts. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI. This approach enables a precise assessment of the relative importance of the topics discussed in analyst reports.
Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010). Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends prior work by quantifying the economic dollar value of such information, accounting for trading costs.
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation. Section 3 summarizes data construction and empirical results. Section 4 concludes.
2 Methodology
In this section, I introduce the large language models used for topic classification and text embeddings, the information content estimation of out-of-sample $R^{2}$, the strategic information value estimation of analyst reports, and the decomposition metric used to evaluate the information content of specific topics. [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
2.1 Large Language Models
Large Language Models such as Bidirectional Encoder Representations from Transformers (BERT), Generative Pre-trained Transformer (GPT), and Large Language Model Meta AI (LLaMA) are trending in literature for understanding textual data. These models represent a paradigm shift in NLP, offering capabilities that go beyond traditional methods like word-to-vec (W2V) or Latent Dirichlet Allocation (LDA) topic models. In this study, I harness the power of these advanced generative AI models to conduct a comprehensive analysis of analyst reports. The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors. This approach not only enriches information extraction but also provides a nuanced understanding of subtle linguistic cues and thematic elements that may influence market reactions.
2.1.1 ChatGPT Prompting
To effectively categorize the topics discussed in analyst reports, I employ ChatGPT prompting to extract exclusive and consistent topics and avoid ad-hoc classification. By leveraging AI, my goal is to identify the primary topics covered in these reports and assign each sentence to a single, specific topic category. This approach ensures a systematic and coherent classification of the content within analyst reports.
In the first step, to derive manageable and meaningful topics, I input a random sample of 100 analyst reports into ChatGPT, requesting an analysis of the content to identify high-level categories.
Prompt 1:Please read the provided text file of sell-side analyst reports carefully. What are high-level mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
It ends up with 13 categories defined as follows.
  • Executive Summary: Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
  • Company Overview: Offers a comprehensive description of the company, including its history, business model, and key products or services; discusses the company’s organizational structure, management team, and corporate governance; analyzes the company’s mission, vision, and strategic objectives.
  • Industry Analysis: Provides an in-depth analysis of the industry in which the company operates; includes information on market size, growth trends, and key drivers; discusses the regulatory environment, technological advancements, and other external factors affecting the industry; analyzes the industry’s competitive dynamics and the company’s position within the industry.
  • Competitive Landscape: Identifies the company’s main competitors and their market share; compares the company’s products, services, and pricing strategies with those of its competitors; analyzes the strengths and weaknesses of the company and its competitors; discusses potential new entrants and substitutes that could disrupt the competitive landscape.
  • Financial Analysis: Analyzes the company’s revenue, expenses, profitability, assets, liabilities, shareholders’ equity, liquidity, solvency, capital structure, cash flows, and key financial ratios; compares the company’s financial performance with that of its competitors and industry benchmarks; discusses trends in the company’s financial performance over time.
  • Business Segments: Provides a detailed analysis of the company’s various business segments or divisions; discusses the financial performance, growth prospects, and challenges of each segment; analyzes the contribution of each segment to the company’s overall revenue and profitability.
  • Growth Strategies: Discusses the company’s strategies for driving future growth, such as organic growth initiatives, product innovations, and geographic expansions; analyzes the company’s mergers and acquisitions (M&A) strategy and potential targets; examines the company’s investments in research and development (R&D) and marketing.
  • Risk Factors: Identifies and analyzes the key risks facing the company, such as market risks, operational risks, financial risks, and legal/regulatory risks; discusses the potential impact of these risks on the company’s financial performance and growth prospects; examines the company’s risk management strategies and mitigation measures.
  • Management and Governance: Provides an overview of the company’s management team, including their experience, expertise, and track record; analyzes the company’s corporate governance practices, such as board composition, executive compensation, and shareholder rights; discusses the company’s succession planning and key person risks.
  • Environmental, Social, and Governance (ESG) Factors: Analyzes the company’s performance and initiatives related to environmental sustainability, social responsibility, and corporate governance; discusses the potential impact of ESG factors on the company’s reputation, risk profile, and financial performance; examines the company’s compliance with ESG regulations and industry standards.
  • Valuation: Estimates the intrinsic value of the company’s shares using various valuation methodologies, such as discounted cash flow (DCF) analysis, relative valuation multiples, and sum-of-the-parts analysis; compares the company’s valuation with that of its peers and historical benchmarks; discusses the key assumptions and sensitivities underlying the valuation analyses.
  • Investment Thesis: Summarizes the key reasons for investing (or not investing) in the company’s shares; discusses the potential catalysts and risks that could impact the company’s valuation and stock price performance; provides a target price or price range for the company’s shares based on the valuation analyses and investment thesis.
  • Appendices and Disclosures: Includes additional supporting materials, such as financial statements, ratio calculations, and detailed segment data; provides important disclosures, such as the analyst’s rating system, potential conflicts of interest, and disclaimers; discusses the limitations and uncertainties of the analysis and the need for further due diligence by investors.
The Financial Analyses category is further divided into four subcategories: Income Statement Analyses, Balance Sheet Analyses, Cash Flow Analyses, and Financial Ratios. Additionally, a category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the predefined topics. In total, this process generates 17 distinct topics for classification. The distribution of topics is relatively stable across years, as illustrated in Figure 1. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, a random sample of 17,028 sentences from 100 analyst reports is selected from analyst reports. I employ the GPT-4-Turbo model through the ChatGPT API to categorize each sentence in this sample into a single, relevant topic.
Prompt 2: Please read the following sentence from the sell-side analyst report of the company {firm} ({ticker}) carefully. Determine which category of information it belongs to in the following 17 categories: {categories}. Pay attention to the sentence in the context of the report.
Output your response in JSON format.
Here is the sentence from the analyst report: {sentence}.
This training sample serves as the foundation for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Conversely, Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level. Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in focus on these issues.
2.1.2 Text Embeddings
Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary. This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand. [^2] BERT uses 30,000 tokens, OpenAI GPT models use around 50,257 tokens, and LLaMA models use approximately 32,000 tokens.
To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”. This granular approach allows for nuanced interpretation of complex terms. For instance, ‘next-gen workloads’ becomes five separate tokens: ‘next’, ‘-’, ‘gen’, ‘work’, and ‘loads’.
At the heart of LLMs lies the transformer architecture. Introduced by Vaswani et al. (2017), the transformer architecture revolutionized the field of NLP with its novel approach to sequence-to-sequence tasks. Unlike recurrent neural networks (RNNs) and long short-term memory networks (LSTMs), transformers use a self-attention mechanism that processes input sequences in parallel. This approach allows the model to weigh the importance of different words in relation to each other, effectively capturing context regardless of word distance. The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance in various NLP tasks.
Text embedding, a core output of the transformer architecture, transforms language into a format suitable for computational analysis. These embeddings are dense, continuous vectors that capture semantic and syntactic information. In models like BERT, GPT, and LLaMA, self-attention mechanisms generate these embeddings, encoding contextual information for each token. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. The embedding dimensions vary: BERT uses 768 or 1,024, GPT-2 Small uses 768, and LLaMA-2-13B model uses 5,120. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The tokenizer processes the text data, converting it into tokens that the model can understand. By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports. These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and text classification. The reason I employ LLaMA2 as a baseline Language Model is due to its 4,096-token input limit, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.
2.2 Information Content
I use out-of-sample $R^{2}$ as a measure for the information content of analyst reports. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
$CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (1)
[…]
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ is the structured representation of information from analyst report $j$ on stock $i$ released at day $t$. I use the $60\%$ of data as the initial training sample. Therefore, the out-of-sampel test starts in 2015 and ends in 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ as
$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum CAR_{[-1,+1],it}^{2}},$ (3)
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation of the model. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
To evaluate the statistical significance of the $R_{\mathrm{oos}}^{2}$, I employ a modified version of the Diebold and Mariano (2002) test for comparing predictive accuracy between two models. Given the strong cross-sectional dependence in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional average of prediction errors from each model rather than individual returns.
2.3 Information Value
I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal based on the theoretical framework of Back et al. (2000):
$\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, multiplied by the closing price of the day prior to the [t-1, t+1] window. [^3] Due to risk neutrality, the equilibrium of the continuous-time Kyle model holds if $\tilde{v}$ is an unbiased signal of the asset price. The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
Specifically, $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ denotes the number of analyst reports about stock $i$ on day $t$. To estimate the price impact $\lambda_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals. The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^4] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression model used is
$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
where $\widehat{\lambda}_{it}$ represents the estimated price impact (or Kyle’s lambda). In this context, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$. Thus, each regression has $1170(=390*3)$ observations.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. This setting is particularly apt for analyzing the value of analyst reports, as it accounts for the correlated nature of information and the strategic interactions among market participants. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To tackle this issue, I apply the delta method in estimating the mean and variance of information value over sample $s$. The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
2.4 Shapley Value Decomposation
To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology, as detailed by Lundberg and Lee (2017) and applied to financial contexts by Chen et al. (2022). SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. In my context, this allows me to attribute the overall explanatory power of an analyst report to its constituent topics. Leveraging the additive nature of text embeddings, I design a novel decomposition method based on SHAP values. This approach enables me to systematically assess the relative importance of various topics in analyst reports.
First, LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ is the total number of words in the report. If the tokens can be classified into different topics, the embedding can be fully decomposed into different topic-specific embeddings:
$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
One concern in the architecture of a transformer model is that each token in the input sequence can pay attention to other words through a mechanism called “self-attention” or “scaled dot-product attention.” This mechanism allows the model to weigh the importance of each word relative to every other word in the sequence, thereby capturing contextual information. Consequently, the embedding of a word contains information from other words, which may pertain to different topics. The embedding of the word might predominantly reflect the context information from other topics. To isolate the self-attention across topics, I process the embedding for each sentence in analyst reports and take the token-weighted mean as the report embedding.
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context in analyst reports.
Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Empirically, the Shapley value can be written as:
$\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly. [^5] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fairly distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions to analyst reports.
3 Data and Empirical Results
This section presents the empirical analysis of analyst report information. I begin by examining the information content of these reports and decomposing it across different topics. I then quantify the strategic dollar value of analyst information. Cross-sectional analyses and robustness checks are conducted to validate the main findings.
3.1 Data
Sell-side analyst reports are obtained from the Mergent Investext Database. I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023. The initial reports are in PDF format, and I use Adobe Acrobat to extract structured text from each report. The segmented text is chunked into sentences. To get clean content of analyst discussions, I train a BERT model to remove the boilerplate segments from each report. [^6] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2023), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports. The BERT model trained for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
The Mergent Investext Database provides information on analyst names and the release date for each report. To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al. (2023). I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database. Following Huang et al. (2014), I define the matching window for each report as 2 days before the announcement date to 2 days after the revision date, accounting for multiple reports released between these dates. [^7] In I/B/E/S, each analyst forecast for a specific firm has two time stamps. The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one. The revision date (REVDATS) is the date the forecast becomes invalid or is revised. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and this time window. The numerical measures include earnings forecasts, stock recommendations, and price target information from analyst reports. I also compile firm characteristics from the CRSP and Compustat databases. [^8] See Table A2 and Table A3 for detailed definitions and summary statistics. After excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 firms from 1,305 analysts across 140 unique brokerage firms.
Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts for each year, or across the Fama-French 12 industries. It also reports the average number of pages and tokens per report, which are important indicators of report lengths. Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al. (2023), and the average number of pages per report is consistent with that in Huang et al. (2014). [^9] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline. This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
The daily stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023. To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
In Section 3.2.2, I explore textual information from earnings conference calls and subsequent analyst reports. The earnings call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005101010Available at https://seekingalpha.com/earnings/earnings-call-transcripts. These transcripts provide comprehensive accounts of the calls, including executive presentations and Q&A sessions with analysts, offering valuable insights into manager-analyst interactions. I sampled earnings call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks. This process yielded a final sample of 7,909 firm-quarter-level conference calls.
Given my goal to estimate the information value of analyst outputs, I need to define the information set available to investors. The numerical information and variables used as inputs to machine learning models are as follows:
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ indicates target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $Boldness$ is an indicator variable that takes the value of 1 if revisions are either above both the consensus and previous forecasts or below both. $SR$ represents stock recommendation, following the I/B/E/S rating system, where a rating of 5 denotes a strong sell and a rating of 1 indicates a strong buy. $ERet$ represents the 12-month return forecast, calculated by scaling the 12-month price target by the stock price 50 days before the report release. $Prior\_CAR$ represents the cumulative abnormal return over time period [-10, -2].
Firm-Level Measures: $Size$ is measured by market equity. The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value. Following Huang et al. (2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value. The $Miss$ variable is a binary indicator set to one if $SUE$ is positive. $TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding. The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures the firm’s product market fluidity.
Industry-Level Measures: $IndustryRecession$ is an indicator variable equal to one if the firm’s industry (defined by Fama-French 48 classification) experiences a negative monthly return in the bottom quintile.
Macroeconomic Measures: The $TimeTrend$ variable captures the temporal dimension by recording the number of years elapsed since the beginning of the sample period.
3.2 Information Content of Analyst Reports
In this section, I provide yearly $R^{2}_{\mathrm{OOS}}$ of analyst reports for explaining contemporaneous cumulative abnormal returns to examine the information content of analysts’ quantitative and qualitative outputs. Furthermore, I attribute the information content of analyst report text across topics using Shapley value decomposition.
3.2.1 Information Content of Quantitative and Qualitative Information
Table 2 compares the information content embedded in numerical and textual measures, quantified through their $R^{2}_{\mathrm{OOS}}$. The estimation model used in this table is the Ridge regression, where the objective variable is $CAR_{[-1,+1]}$ centered at the analyst report release date. For reports released on non-trading days, I adjust t=0 to the next trading day as recorded in the CRSP database. The estimation methodology employs an expanding window approach, constructing training sets from 2000 to the year preceding each target year. I estimate separate Ridge regression models for each input configuration, conducting yearly out-of-sample estimations of $\widehat{CAR}_{[-1,+1]}$, allowing for dynamic assessment of model performance over time.
Panel A of the table delineates the performance of four distinct input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) comprehensive numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, the 5,120-dimensional full-context representations of each report generated by LLaMA-2-13B model; and (4) a hybrid input combining revision measures with textual embeddings.
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for the revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions. The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features. I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4). The Diebold-Mariano test follows a standard normal distribution under the null hypothesis of equal predictive accuracy between two competing forecasting models, allowing similar interpretations of t-statistics as in regression analysis (Gu et al., 2020). The models incorporating revision and numerical inputs significantly outperform a zero-estimation benchmark.
Columns (5) and (6) elucidate the performance of text embedding input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20. This underscores the significant information content of the report text in explaining contemporaneous stock returns. The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession. Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs. Language model outputs may incorporate future information that should not be available. Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period. The Diebold-Mariano test comparing revision-only and text-only inputs yields a t-statistic of 1.66, highlighting the substantial information content in the report text.
In columns (7) and (8), I further investigate the performance of combining both revision and text input in the Ridge regression. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model. I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5). The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$. The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by market investors.
To address the concern that the explanatory power might be derived from quantitative information in analyst reports, I conduct robustness checks using content with numbers removed. In Table A4, I exclude numerical data from the written reports before generating text embeddings. The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, still significantly higher than that of quantitative revision measures (t-statistic = 2.40). This suggests that the superior performance of the text-based model is not driven by numerical information within the reports.
LLMs are known to exhibit improved language skills as the size of their parameters increases. I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI GPT-3 embedding model, and LLaMA-3 model. The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between each model and the LLaMA-2 model. The input in Panel B is report text, comparable to column (3) in Panel A. A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark. Overall, the BERT and OpenAI models underperform the LLaMA-2 model at the 1% level for each year. The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model. The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs. The pairwise comparisons between LLMs also shed light on how the scale of the model influences its performance capabilities.
To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al. (2014) and Asquith et al. (2005), respectively. Additionally, I examine the abnormal stock return on the days that reports are released. In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023. In Figure A3, I plot the information value of analysts’ information by industry, revealing a pattern consistent with the information content results.
In Table A6, I present the information content of analyst report text across the Fama-French 12 industries. The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables. Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors. The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value.
To assess economic magnitude, I perform OLS regressions as specified in Equation (10):
$CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (10)
This model incorporates the out-of-sample expanding estimation of $CAR_{[-1,+1]}$ using report text embeddings as input. For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014. The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension. For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
Table 3 summarizes the summary statistics of the sample used and the regression results. The out-of-sample period starts in 2015. I constrain the samples to include only those with valid previous and current stock recommendations, earnings forecasts, and target prices. The final sample for estimating Equation (10) contains 28,837 observations.
Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al. (2014). Column (2) reveals that $\widehat{CAR}{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2. In column (3), the inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% to 14.9%, underscoring the incremental value of text information.
The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by quantitative analyst revisions. Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the three-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics. The $\widehat{CAR}{rev}$ model alone achieves an $R^{2}$ of 8.9%, while the combined input model improves this to 14.7%. Economically, a one standard deviation increase in $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises. The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged. This stability in coefficients suggests that the market-valued analyst information is distinct from, and not merely chasing back on, past returns or recent news events.
3.2.2 Information Content around Earnings Announcements
Some studies have emphasized the role of analysts surrounding corporate earnings announcements (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information events reinforce each other in sequence. I, therefore, further study the information content of sell-side analyst reports around this most significant corporate event.
Following Chen et al. (2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week. Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week. A finding consistent with Huang et al. (2018) is that an overwhelming number of analyst reports are released immediately following the earnings announcements. Consistent with Huang et al. (2018), a large number of analyst reports are released immediately after earnings announcements. Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17. This performance drops sharply in the second week, turns negative in the third and fourth weeks, and does not recover thereafter.
The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports during earnings announcement weeks. To account for potential temporal variation, I train Ridge models using separate samples for periods surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports relative to earnings announcements. Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained outside these periods. Additionally, the $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen. This trend suggests that analysts react swiftly to earnings announcements, with the information content of their reports declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week.
One potential concern in assessing the value of analyst reports is that analysts might merely be repeating the content of earning announcement calls, with the earnings announcement itself driving market reactions. If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls. To address this issue, I use earnings conference call transcripts data to differentiate the information content between the two text sources. The analysis focuses on reports released within a one-day window following earnings announcements, as this is when analysts are most likely to piggyback on the announcements.
In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcripts as input. [^11] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C. This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study. I truncate the text at the length limits to maintain maximum information. The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16. with a Diebold-Mariano (DM) t-statistic of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods of both texts. The input comprised the mean of all report text embeddings on the day following the earnings announcement. This firm-time level Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
Incorporating both analyst report embeddings and earnings conference call embeddings in the Ridge regression resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3). The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports. These findings collectively refute the idea that analysts merely restate information from earnings announcements. Instead, they emphasize the unique and timely value analysts bring to the market.
3.2.3 What Content in Analyst Reports is Valued by Market Investors?
Given the comprehensive discussion in equity reports, what specific content in analyst reports is the most valuable to the market? I employ the Shapley value approach in Section 2.4 to attribute the $R^{2}_{\mathrm{OOS}}$ of the full text to 17 topics. In this section, I use the token-weighted average of each sentence embedding as the report embedding. This sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another. The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings. This reduction roughly quantifies the magnitude of contextual information content in analyst reports.
Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations. These reductions are then averaged into a single importance measure for each topic. By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring a comprehensive attribution of explanatory power. Figure 2 ranks the Shapley values of 17 topics. Topics are ordered from left to right, with the most important topics on the left and the least important on the right. Each topic is represented by a bar, with a color gradient (from lightest to darkest) indicating the magnitude of its importance in terms of Shapley values.
The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports. The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each comprising 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$. Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively. The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
The robustness of these findings is further reinforced by additional analyses. As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses, underlining the consistency of the results.
Figure A5 in the Online Appendix reveals how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time. The annual variations align with the full-context embedding results in Table 2, exhibiting a consistent story across different text embeddings. A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy. Despite this volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable, indicating that analyst discussions of income statements consistently provide valuable information, even during market disruptions.
Figure A6 further supports this by showing a consistent pattern of key topics across industries. This suggests that the market values Income Statement Analyses and Financial Ratio Analyses universally, beyond specific sectors.
To investigate whether this topic ranking is driven by analysts’ expertise in certain topics (supply effect) or investors’ differential responses to different topics (demand effect), Figure A7 compares Shapley values across analysts with varying experience levels and from different-sized brokerage firms. The remarkable consistency in topic importance across all four subsamples (experienced vs. inexperienced analysts, large vs. small brokerages) eliminates the concern that the topic ranking merely reflects analysts’ varying skill levels in different topics.
Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements. Drawing inspiration from the literature, particularly Huang et al. (2018) and Blankespoor et al. (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
To facilitate this analysis, I further categorize sentences within the Income Statement Analyses topic into two sets of sub-topics: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a sophisticated prompt-based classification system, as detailed in Prompt 3:
Prompt 3: Please read the following sentence from a sell-side analyst report for the company {firm} ({ticker}) and classify it based on two criteria:
Information Type:
Information Acquisition (0): Sentences that directly report quantitative financial data, such as earnings, revenue, expenses, or other metrics. Example: ‘Reported EPS of \$1.40 beat consensus of \$0.94 and our estimate of \$1.17.’
Information Interpretation (1): Sentences that provide analysis, context, or interpretation of the financial data, such as discussing trends, comparing performance to forecasts, assessing market impacts, or considering strategic implications. Example: ‘Biogen reported 1Q12 MS franchise sales below our estimate and the Street, as Avonex sales were negatively impacted by unfavorable distribution channel dynamics.’
Time Reference:
Realized Income (0): Sentences referring to concrete, historical results from completed periods. Example: ‘The company’s revenue for the fiscal year 2022 was \$500 million.’
Expected Income (1): Sentences containing subjective predictions, or expectations for future periods. Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
Output your classification as two comma-separated numbers: the first for Information Type (0 for Acquisition, 1 for Integration) and the second for Time Reference (0 for Actual, 1 for Forecast).
Output format: InfoType, TimeRef
Sentence to classify: {report_sent}.
Figure 3 offers a detailed breakdown of the $R^{2}_{\mathrm{OOS}}$ across 18 topics, with the Income Statement category further divided into sub-topics. Specifically, the income interpretation subtopic contributes approximately three times more to the $R^{2}_{\mathrm{OOS}}$ than income acquisition, underscoring the market’s preference for context and explanation over mere data reporting. Additionally, the income realization subtopic also contributes about three times more to the $R^{2}_{\mathrm{OOS}}$ than income expectation, indicating a stronger market preference for analysis of actual financial results over forward-looking projections. These findings suggest that the market values analysts’ interpretation of realized income to a greater extent than their ability to simply report public information or make expectations about future performance.
This insight is reinforced by the findings in section 3.2.2, which show that analyst reports issued promptly after earnings announcements are the most informative. Taken together, these results suggest that the market places the highest value on analysts who can quickly digest financial information from earnings announcements and provide insightful interpretations.
3.2.4 Comparison of Revision Reports and Reiteration Reports
Researchers have documented a robust relationship between market prices and analysts’ forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021). Analysts’ forecast revisions provide a significant and timely source of information to the financial market and may constitute a more valuable signal than reports that merely reiterate previous forecasts.
Therefore, I compare the information content of analyst report text associated with forecast revisions and forecast reiterations. Following the criteria of Huang et al. (2014), I consider a report as a revision report if its release date is within two days of the I/B/E/S forecast announcement date. For each forecast target in recommendations, target prices, and earnings forecasts, I categorize the report independently. Consequently, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 5 summarizes the out-of-sample performance of numerical and textual inputs in both reiteration and revision samples. Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions. When recommendations are revised, the explanatory power of various information sources becomes particularly pronounced. The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%. The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%. The combination of both quantitative and qualitative information sets further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%. This finding aligns with the literature that attributes large stock price changes to analyst recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions, considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), also reveal a clear contrast between revisions and reiterations. While target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ of 20.88%, t-stat = 11.18).
Earnings forecast revisions are the most frequent, comprising 56.57% of the reports. In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively. As Kothari et al. (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output. Interestingly, the $R^{2}_{\mathrm{OOS}}$ for report text in earnings forecast reiteration reports is negative, indicating no explanatory power for contemporaneous stock returns. However, when focusing on analyst earnings forecast revisions,, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant. The combination of these inputs yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
Collectively, Table 5 shows that report content is more informative when it accompanies forecast revisions rather than reiterations. Additionally, analyst reports associated with recommendation revisions demonstrate the highest value.
3.2.5 Comparison with Analyst Report Sentiment
Huang et al. (2014) emphasize the importance of analyst report text by extracting tone using a Naive Bayes machine learning approach. My study extends this research by incorporating more advanced techniques, including BERT-based tone measures and text embeddings.
I construct report-level tone measures using both the Naive Bayes approach and BERT classification. Following Huang et al. (2014), I manually label 10,000 sentences as a training sample for both models. Each sentence is categorized into positive, neutral, or negative using these pre-trained models, then average the tone at the report level to derive $Tone\_NB$ and $Tone\_BERT$ measures. I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
Table 6 presents a comparison of performance between text tone measures and text embeddings. Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures. The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al. (2014). A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, which is comparable with the 41 basis points increase in two-day abnormal return in Huang et al. (2014). Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics. Both coefficients of tone are significantly positive. Column (3) reports regression results with firm-fixed and year-fixed effects, consistent with the findings in column (2).
Columns (4)-(6) present regression results with BERT-based tone measures as explanatory variables. The coefficient estimates’ magnitude increases dramatically, along with significance levels. Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points. When both tone measures are included in an OLS regression in column (7), Naive Bayes-based tone measure coefficients become insignificant, suggesting BERT-based measures capture sentiment more accurately and effectively.
Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of extracted tone measures and embedding representations. The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows. The $R^{2}_{\mathrm{OOS}}$ increases from 9.49% to 12.27% after incorporating analyst revision and text embeddings, confirming the superior performance of state-of-the-art LLMs across various NLP tasks. The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The $R^{2}_{\mathrm{OOS}}$ difference between these models is not statistically distinguishable from zero for both Naive Bayes-based and BERT-based tone measures.
In general, the empirical results in Table 6 support two key findings regarding the superiority of LLMs compared to conventional NLPs. First, overall report text tone measures extracted using small-size LLMs are more explanatory for contemporaneous stock returns than those extracted using the Naive Bayes algorithm. Second, tone measures (extracted with both Naive Bayes and BERT) add no incremental value to text embeddings, suggesting that LLMs provide more comprehensive contextualized representations.
3.2.6 Information Content with Alternative Machine Learning Models
In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions. Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$. To ensure out-of-sample evaluation, I train the models yearly and estimate for the following year’s sample. The $R^{2}_{\mathrm{OOS}}$ is calculated by averaging the estimated and realized $CAR$ values. Additionally, I perform pairwise Diebold-Mariano tests using a zero benchmark to assess predictive accuracy.
I begin with PLS regression, summarized in Table 7. PLS is notable for extracting the most relevant common components from sentiment proxies that align with the target variable (Huang et al., 2015). If there is a strong common component in the text embeddings that is informative of stock returns, the PLS regression architecture should effectively capture this relevant information. Examining the model performance, I observe that PLS regressions produce an overall $R^{2}_{\mathrm{OOS}}$ of 8.26%. This measure of information content ranges from 2.14% in the 2021 sample to 12.23% in the 2015 sample, showing a similar pattern across years as seen in Ridge models. Diebold-Mariano tests indicate the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all periods, demonstrating strong evidence of information content in the analyst report text.
Next, I replicate the analyses with the XGBoost algorithm. Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees. Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information. Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns. Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable. As summarized in Table 7, the XGBoost algorithm underperforms with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
Neural Networks, forming the core of LLMs and the transformer architecture, have shown significant advancements in various NLP tasks (Cao et al., 2024; Li et al., 2023). In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain. Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text. I construct the Neural Network estimations as an ensemble of five models for each NN configuration. As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms. Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al. (2022).
The experiments with machine learning models yield noteworthy findings. First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns. Second, the superior performance of deep learning architectures emphasizes the importance of capturing contextual information and word interdependencies for textual analyses.
3.3 Information Value of Analyst Reports
How can the statistical measure $R^{2}_{\mathrm{OOS}}$ be translated into economic terms? In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3. This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal while considering price impact. In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
I first summarize the value of information from analyst reports in Table 8, covering the period from 2015 to 2023. Both the mean and standard deviation are derived using the delta method. For S&P 100 constituent stocks, the average three-day window information value from sell-side analysts is \$0.47 million, with a standard error of \$0.05 million. The mean information value of text alone is \$0.38 million, and the mean value for revisions is \$0.34 million. Considering that a stock typically has reports released on an average of 15 days each year, the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm is \$6.89 million. The narrow standard errors and confidence intervals across different information values indicate consistent estimates. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, indicating that \$1 million dollar order pushes the stock price up by 0.034%. These statistics illustrate that information from analysts is economically meaningful, with both textual information and revisions contributing significantly.
Figure 4 describes the evolution of information value from analyst reports over time. The solid line represents the strategic value of combined numerical and textual information, while the dashed line shows the value of textual information alone. Both lines exhibit significant fluctuations over the period. Despite these fluctuations, there is a general upward trend in information value. For most of the period, the combined value is higher than the textual value alone, suggesting that numerical information often complements the textual content in analyst reports.
The shaded area represents the 95% confidence interval. After 2020, there appears to be greater variability in the information value, as indicated by the wider confidence intervals and more pronounced peaks and troughs. This period corresponds with the global COVID-19 pandemic, which likely introduced additional uncertainty and volatility in financial markets, impacting the value of information from analyst reports.
These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions and external events. The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value. Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million. Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time. These alternative estimates support my main findings, reinforcing the substantial value of analyst information and the robustness of the approach.
3.3.1 The Value of Analyst Information in the Cross-section of Stocks
It is likely that the information value of sell-side analysts will vary in the cross-section of stocks. Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984). For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995). I plot the scatter of information value and market capitalization by stock in Figure A9. There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values. However, the relationship is not strictly linear, as there are notable exceptions. I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects. I also control for the book-to-market ratio in the regressions.
Table 9 shows the results. Stocks with larger market equity have significantly higher information value from analysts. Economically, column (4) means a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value. I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact for large stocks. This decomposition can be expressed logarithmically as:
$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
Panel B of Table 9 demonstrates that the main reason behind the positive size elasticity of information value is the higher liquidity of large stocks. Overall, analyst information does not appear to explain the higher contemporaneous return volatility of large stocks, yet it remains more valuable due to the higher liquidity of these stocks.
3.3.2 The Value of Analyst Information in the Cross-section of Analysts
I next examine how the information value of analysts differs with herding behavior. Prior research has stated that analysts making bold forecasts incorporate their private information more completely and provide more relevant information to investors than those making herding forecasts (Clement and Tse, 2005).
To investigate the cross-section of analyst characteristics, I regress the stock-analyst-day information value on an indicator of forecast boldness, controlling for firm-fixed effects, analyst-fixed effects, and year-fixed effects. I also control for broker size, as Jacob et al. (1999) suggests that larger brokerage firms provide better research. The analyst-specific information value measure is defined as
$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
Table 10 presents the results. In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts. On average, information from bold reports is 29% more valuable than that from herding reports. The magnitude remains largely unchanged when testing with information value derived from report text. However, the insignificant coefficient estimates for $Brokersize$ suggest that the value of analyst information is not necessarily tied to the size of the brokerage firm.
The subsequent research question is whether the association between boldness and information value is driven by higher information content or lower price impact. In Panel B, I regress the two decompositions of information value on the boldness indicator. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information. In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant. This finding suggests that while bold forecasts contain more valuable information, they do not necessarily lead to a higher price impact compared to herding forecasts.
3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. To expand on this, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 illustrates the information value of analyst reports between earnings announcements on a weekly basis, from the first week after the earnings announcement up to the $13^{th}$ week. The information value of analyst reports is highest immediately following earnings announcements, particularly in the first week, peaking at a sample mean of \$0.84 million. This value sharply declines by the second week and stabilizes thereafter, with minor fluctuations. These patterns suggest that the market values analyst reports most highly immediately after earnings announcements, with diminishing importance in subsequent weeks.
To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$). This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010). The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al. (2024). The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect information value. The model includes stock and year-fixed effects to control for time-invariant stock characteristics and market-wide yearly trends.
Table 11 presents regression results showing the influence of earnings announcements and trading volume on information value. Panel A indicates that the information value is significantly higher for reports released within one week of earnings announcements ($Week$), with positive and significant coefficients across all models. The interaction between $Week$ and $TradingVolume$ also shows a significant positive effect, suggesting that the information value increases with higher trading volumes around earnings announcements. The rationale is that investors may rely more on analyst opinions when there is greater disagreement in the interpretation of earnings information.
Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable. The interaction term ($Week\times TradingVolume$) is significantly positive in columns (1)-(4), indicating that explained return volatility rises with higher trading volumes during post-earnings periods. Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact. These findings suggest that analyst reports provide more valuable insights in the immediate post-earnings announcement period, particularly when investor uncertainty is high. This increased value likely reflects analysts’ ability to interpret complex earnings information, helping investors make sense of the latest financial data during these critical periods.
4 Conclusion
In this paper, I examine both quantitative and qualitative content of analyst reports using $R^{2}_{\mathrm{OOS}}$. The results show that report text contains more market-valued information content than quantifiable forecast revisions economically, especially right following earnings announcements. A Shapley value decomposition method reveals that analysts’ interpretation of realized income statements is the most valuable content within written reports. The text embedding representation extracts more comprehensive information from analyst reports beyond text sentiment.
I further translate the information content into dollar terms to provide a sense of economic magnitude. Using a measure of explained return volatility relative to price impact, I find that the annualized expected profit from early information acquisition of the analyst reports for an average S&P 100 constituent firm is larger than \$6.89 million. While these results are economically meaningful, readers should exercise caution in generalizing beyond the S&P 100 sample. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The significant potential profit from early access to analyst reports raises concerns about selective distribution. Such practices would violate FINRA Rule 2241, which aims to ensure the objectivity of research and prevent the selective distribution of research reports to preferred clients ahead of others. It is crucial to emphasize that this study presents a hypothetical scenario without assessing or endorsing actual market practices. Rather, by quantifying the potential profit of privileged access to analyst information, this research emphasizes the importance of regulations in ensuring fair disclosure.
A. Additional Figures
(a) (1) Executive Summary
(b) (2) Company Overview
(c) (3) Industry Analysis
(d) (4) Competitive Landscape
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (5) Income Statement Analysis
(b) (6) Balance Sheet Analysis
(c) (7) Cash Flow Analysis
(d) (8) Financial Ratios
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (9) Business Segments
(b) (10) Growth Strategies
(c) (11) Risk Factors
(d) (12) Management and Governance
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (13) ESG Factors
(b) (14) Valuation
(c) (15) Investment Thesis
(d) (16) Appendices and Disclosures
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
Figure A2: Large Language Model Architecture
This figure shows an example of how large language models process a sentence from an analyst report on Nvidia. The model breaks this sentence into tokens, and then generates two types of embeddings: token embeddings and position embeddings. These embeddings then pass through the core of the model - a stack of transformer layers. The output of each transformer layer is a series of vectors representing the tokens in their full context.
Figure A3: Analyst Information Value by Industry
This figure shows the estimated information value of analyst reports for Fama-French 12 industries from 2015Q1 to 2023Q4, quantified in millions of dollars. The solid green line marks the sample mean value of information, with green bars showing the 95% confidence intervals. The horizontal dashed red line marks the sample mean of stocks from all industries. Information value estimates are derived using the delta method and adjusted to 2020 levels using the Consumer Price Index (CPI).
Figure A4: Shapley Values Scaled by Length
This figure presents the SHAP values of topics scaled by the total number of sentences (Panel A) and the total number of tokens (Panel B) for topics with more than 100,000 sentences. The scaling process involves two steps: First, normalize the Shapley value of each topic by dividing it by the number of sentences in that topic. Then, adjust the normalized values to ensure the total sum of SHAP values remains consistent. This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
Figure A5: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
This figure shows the $R^{2}_{\mathrm{OOS}}$ of sentence-segmented report embeddings from 2015 to 2023. The color gradient indicates the magnitude of $R^{2}_{\mathrm{OOS}}$. The red line represents the SHAP value of the Income Statements Analyses topic across the years. The shaded area marks the pandemic recession in 2020.
Figure A6: Shapley Values for Topic Importance across Industries
This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value. For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings. The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
Figure A7: Shapley Values for Topic Importance by Analyst Characteristics
This figure illustrates how 17 different topics contribute to the predictive power of analyst reports across four subsamples, measured using SHAP values. The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), and large (Panel C) versus small brokerage firms (Panel D), with subsamples determined by yearly median. For each subsample, the bars show the SHAP values of individual topics, whose sum equals the out-of-sample R-squared ($R^{2}_{\mathrm{OOS}}$) from the sentence-segmented report embeddings. The red line indicates each topic’s relative importance, calculated as the topic’s SHAP value divided by the total $R^{2}_{\mathrm{OOS}}$.
Figure A8: Alternative Analyst Information Value Estimations Over Time
This figure presents the alternative estimations of analyst information value from 2015Q1 to 2024Q4, reported in millions of dollars and adjusted to 2020 dollars. Quarterly mean approximations are calculated using the delta method. TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Figure A9: Analyst Information Value by Stock
This figure shows the estimated information value of analyst reports for individual stocks from 2015Q1 to 2023Q4, quantified in millions of dollars. Estimates are derived using the delta method. Market equity, averaged daily at the time of each report release, is reported in billions of dollars. All values are adjusted to 2020 levels using the Consumer Price Index (CPI).
B. Additional Tables
Table A1: Topic-Specific Sentence Examples from Analyst Reports
Topics Examples
Executive Summary
Our key takeaways from CEO Jeff Immelt’s presentation at EPG were: Outlook for substantial EPS growth over 2010/12 driven by abatement of credit losses (\$8-9bn in 2010E tapering to \$4bn run-rate) and CRE impair. Looking ahead, guidance was tightened, essentially framing the Street, and commentary suggests a strong outlook for the balance of 2007. Key topics: 1) Update on ABTÕs portfolio of COVID-19 tests; 2) Libre trends, Libre 2 launch and expectations for Libre 3; 3) Elective surgery trends exiting Q1, expectations for 2021 and an update on recent and upcoming new product approvals; and 4) Global trends and impact on EPD and Nutrition.
Company Overview
Oracle Corporation, founded in 1977 and headquartered in Redwood Shores, California, is one of the largest and most prominent companies in the software space – and a technology bellwether. As it has grown, Microsoft has expanded into enterprise software with Windows Server, SQL Server, Dynamics CRM, SharePoint, Azure and Lync; hardware with the Xbox gaming/media platform and the Surface tablet; and online services through MSN and Bing. Altria Group, Inc., is the world’s largest producer and marketer of consumer products, and had revenues of \$80 billion in 2002.
Industry Analysis
According to our global Immunology market model, US Psoriasis (PsO) represented a \$7.7B market in 2016 and is expected to grow at a low-teens CAGR to \$11.8B in 2019E and  \$13B by 2021E driven by more highly effective therapies. Despite record prices, oil demand continues to grow, while supply growth lags and spare production and refining capacity is almost nonexistent. Add to that that some new advertising expectations from PricewaterhouseCoopers of a decline in ad spending of 12% worldwide and 15% in the U.S. for 2009 and continue to decline in 2010.
Competitive Landscape
According to Mercury Research, NVIDIA is now the 3rd largest chipset supplier (consisting of desktop and mobile chipsets, and integrated and non-integrated chipsets), shipping 5.4 million units in calendar Q3 for an 8.2% market share, versus Intel’s shipments of 51 million units (62% share), VIA’s shipments of 9.6 million units (14.4% share), SiS’s shipments of 5.3 million units (8% share), and ATI’s shipments of 4.4 million units (6.6% share). Further, competition in the CDK-4/6 space is rising with Verzenio (abemaciclib) & Kisqali launches placing downward pressure on Ibrance trajectory. We believe that Accenture has recognized that web services could compete directly with client/server as the new systems architecture.
Financial Statement Analysis: Income Statement Analysis
We also note that Alico cannot dividend any of its 2010 earnings to AIG, implying that the income recognized this quarter and throughout 2010 will benefit MetLife results. At the respective midpoints, sales of \$8.5 billion would be down 22% annually and 8% sequentially; and non-GAAP EPS would be down 39%. The growth in operating income from Q1 2009 is largely due to recoveries in partnership income.
Financial Statement Analysis: Balance Sheet Analysis
Long-term borrowings of \$6.59 billion at September 30, 2022 were modest compared to shareholders’ equity of \$37.3 billion. The company finished 2Q16 with \$21.4 billion in cash and short-term investments, up from \$15.6 Growth & Valuation Analysis GROWTH ANALYSIS RISK ANALYSIS billion at the end of 4Q15. Inventory turns improved to 4.5x in 4Q from 4.4x in the same period last year.
Financial Statement Analysis: Cash Flow Analysis
UPS generated \$2.3 billion in operating cash for the quarter. Assuming that the company completes a large portion of its current \$1 billion stock buyback plan in 2008, we estimate that cash per share will be about \$6.20 by the end of the year. We view eBays FCF generation as relatively defensible even in the case of a revenue shortfall.
Financial Statement Analysis: Financial Ratios
The company achieves average scores on our three main measures of financial strength: leverage based on debt-to-cap, profitability and interest coverage. P/S is at 0.6x and EV/EBITDA is 7.7x. The index members currently trade at an average of 16.3-times trailing earnings, which is below the five-year average of 19-times.
Business Segments
The company is organized into three businesses, software, representing the majority of the company’s total revenues, hardware systems and services. Results in 2008 also benefited from the absence of the significant level of mark-to-market losses in the company’s Gas Marketing segment in 2007. The company operates five distinct segments: Americas (71% of FY15 profit); Europe, Middle East, and Africa (4% of FY15 profit); China and Asia Pacific (12% of FY15 profit); Channel Development (13% of FY15 profit).
Growth Strategies
Visa has been especially active on the acquisition front over the last several months. Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle. Specific areas where more investment may be needed: Over the past decade, ROK management often claim that Process automation represents the growth opportunity for the company, but its sales in this market have barely grown in recent years, despite its much smaller sales base compared with established incumbents.
Risk Factors
Risks to our BUY thesis have to do with global competition, changing user behavior, global macro uncertainty, and anything else that can affect FB’s relationship with members, its advertisers or its publishing partners. Risks to achieving our price target include: 1) Apple crushing PayPal; 2) increasing competition in the payments space; 3) heavy investment spending on marketing, point of sale, or technology; and 4) legislative action. In addition to the expenses incurred by patent challenges, product liability and other legal suits could occur and lead to additional liabilities and revenue loss, which could substantially change our financial assumptions.
Management and Governance
Top management changes can be unsettling, and the resulting uncertainty has caused 3M shares to decline. Chairman, President, and CEO Charles Ergen beneficially owns about 53.6% of DISH’s equity securities and has 90.5% voting power. Bill Johnson, the present CEO of Progress Energy, will become president and chief executive officer of the new company.
Environmental, Social, and Governance (ESG) Factors
The assessment of ESG (Environmental, Social & Governance) risks by Baptista Research includes a wide range of considerations that pertain to the long-term sustainability of a company. Sustainalytics assesses the degree to which a company’s enterprise business is affected by ESG issues. Failure to adequately address social risks like labor disputes and community relations could jeopardize the company’s social license to operate in certain regions.
Valuation
Our DCF derives an intrinsic value of ¿\$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC. We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment. Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
Investment Thesis
The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road. As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high. Clearly, our Ford Investment Thesis, which was based in large part on our belief that Ford would be able to offset headwinds (slowing cyclical tailwinds in North America, weakness in South America, weak growth in Europe, slowing growth in China, and regulatory cost headwinds), has been thrown into question.
Appendices and Disclosures
Although the information contained in the subject report has been obtained from sources, we believe to be reliable, its accuracy and completeness cannot be guaranteed. For a complete discussion of the risk factors that could affect the market price of a company’s shares, refer to the most recent Form 10-Q or 10-K that a company has filed with the Securities and Exchange Commission. The Benchmark Company, LLC makes every effort to use reliable, comprehensive information, but we make no representation that it is accurate or complete.
Table A2: Numerical Variables Description
This table shows the definitions of numerical measures.
Numerical Measures Definition and/or sources
Panel A: Firm Level Measures
Size The market value equity of the firm ($CSHOQ*PRCCQ$) at the end of the quarter prior to report release.
BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of quarter prior the report release.
Prior_CAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement. Consensus EPS is the median value of 1-year EPS forecast within 90-days window of all analysts following the firm. The unexpected earnings is scaled by price per share at the fiscal quarter end.
AbsSUE Absolue value of SUE, representing the distance between realized EPS and EPS expectation.
Miss Dummy variable that equals one if the actual EPS is less than the last consensus forecast, and 0 otherwise.
Trading volume Trading volume at earnings announcement day (or the first trading day post earnings announcement), calculated as $VOL/SHROUT$.
Distance to default The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
Fluidity A measure of firms’ product market competition introduced by Hoberg et al. (2014). The data is compiled from the Hoberg and Phillips database.
Panel B: Industry Level Measures
Industry recession An indicator variable that equals one if the FF-48 industry return is negative and in the bottom quintile of FF-48 industry returns and zero otherwise.
Panel C: Macroeconomic Measures
Time trend (ttr) The number of years elapsed from the beginning of the sample.
Panel D: Report Level Measures
$REC\_REV$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
$EF\_REV$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
$TP\_REV$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
Boldness An indicator variable that takes the value of 1 if the EPS forecast revision is above both the analyst’s own prior forecast and the consensus forecast, or else below both, and zero otherwise.
SR Stock recommendation from I/B/E/S rating, with 1 being the most bullish (Strong Buy) and 5 being the most bearish (Sell), based on the ratings provided by the Institutional Brokers’ Estimate System (I/B/E/S).
ERet 12-month return forecast by scaling the 12-month price target by the stock price 1-day before release.
Table A3: Sumamry Statistics for Numerical Measures
This table reports summary statistics of numerical measures.
Mean Std p25 P50 P75 N
Miss 0.35 0.48 0.00 0.00 1.00 102776
Trading volume 18.19 21.98 7.24 11.46 20.29 99908
Prior_CAR 0.00 0.05 -0.02 0.00 0.02 122251
Size 2.39 0.11 2.33 2.39 2.46 117534
BtoM 0.46 0.43 0.18 0.32 0.61 117534
Distance to default 0.00 0.01 0.00 0.00 0.00 86693
Fluidity 7.17 3.66 4.40 6.57 9.44 114942
Time trend 13.02 5.71 9.00 13.00 17.00 122252
Industry recession 0.14 0.35 0.00 0.00 0.00 122252
BrokerSize 71.92 56.71 26.00 53.00 113.00 119233
Firm Experience 7.68 6.72 3.00 6.00 11.00 111192
Number of firms 20.08 17.08 14.00 18.00 23.00 119233
REC_REV 0.00 0.16 0.00 0.00 0.00 120673
EF_REV 0.00 0.00 0.00 0.00 0.00 90625
TP_REV 0.01 0.07 0.00 0.00 0.00 84108
ERet 0.19 0.23 0.07 0.18 0.31 84108
Boldness 0.75 0.43 1.00 1.00 1.00 81182
SR 2.77 0.84 2.00 3.00 3.00 122252
Table A4: Information Content of Analyst Reports with Numbers Removed
This table presents the $R^{2}_{\mathrm{OOS}}$ using analyst reports with numbers removed. The input sets include revision numerical measures, text embeddings from number-free report text, and a combination of both. The estimation targets are the three-day $CAR$ centered around the analyst report release date. The machine learning model employed is Ridge regression, with training samples expanding yearly. The $t$-statistic for the $R^{2}_{\mathrm{OOS}}$ is calculated following Gu et al. (2020). In columns (2), (4), and (6), the benchmark estimation is set to zero. Columns (3)-(1) and (5)-(3) present pairwise tests comparing the performance of models using revision-only input versus text-only input, and text-only input versus combined text and revision input, respectively.
Year Revision only t-stat Text only t-stat Revision plus text t-stat t-stat t-stat
(1) (2) (3) (4) (5) (6) (3)- (1) (5)- (3)
2015 10.30% 3.59 13.19% 5.87 11.10% 3.24 3.97 -1.45
2016 15.47% 11.80 13.73% 5.00 17.91% 9.02 -0.81 3.13
2017 9.16% 5.28 11.66% 5.98 12.33% 6.12 10.58 1.56
2018 9.79% 2.98 11.46% 9.12 13.62% 7.97 1.22 4.09
2019 9.76% 15.69 12.61% 14.73 13.90% 15.17 4.26 2.79
2020 5.20% 3.43 4.42% 4.38 6.98% 5.81 -0.70 8.37
2021 5.31% 4.80 8.77% 5.06 10.66% 29.67 1.54 1.56
2022 10.24% 6.16 16.71% 10.47 17.90% 8.94 9.93 2.38
2023 5.09% 3.86 9.27% 3.60 10.39% 5.09 2.60 1.81
Overall 8.93% 8.51 10.95% 7.87 12.51% 8.83 2.40 3.00
Table A5: Cumulative Abnormal Returns with Alternative Windows
This table reports the $R^{2}_{\mathrm{OOS}}$ of using analyst report text embeddings to estimate $CAR$ with alternative windows. The $R^{2}_{\mathrm{OOS}}$ is calculated annually using a training sample from 2000 to the preceding year. $CAR$ measures are centered on the release day of analyst reports ($T_{0}$). Panel A shows results for the full sample, while Panels B and C focus on periods within and beyond 1-day earnings announcement windows, respectively. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
AR[0] CAR[0,+1] CAR[-2,+2]
year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
Panel A: Full sample period
2015 8.03% 8.53 8.24% 8.36 10.87% 9.40
2016 7.88% 8.27 6.09% 4.68 6.60% 3.11
2017 8.75% 14.35 8.56% 9.29 9.00% 8.78
2018 9.09% 14.95 9.39% 11.26 9.75% 10.54
2019 9.89% 99.70 7.68% 9.45 8.66% 18.93
2020 2.33% 2.33 1.76% 3.15 4.13% 2.14
2021 5.75% 7.32 5.80% 6.62 4.05% 3.57
2022 10.54% 12.32 11.23% 18.20 10.91% 13.65
2023 5.38% 15.44 3.39% 7.00 0.53% 1.48
Overall 7.55% 8.07 6.92% 6.77 7.57% 6.79
Panel B: Earnings announcement period
Overall 9.78% 5.58 7.87% 5.10 13.14% 8.02
Panel C: Non-earnings announcement period
Overall 5.77% 5.78 5.62% 5.64 4.96% 4.91
Table A6: Information Content of Analyst Reports across Industries
This table reports the $R^{2}_{\mathrm{OOS}}$ of Fama-French 12 industries, using the definition from Kenneth French’s website. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
year Shops Other Manuf Chems
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 13.27% 6.03 657 17.69% 5.03 585 10.84% 3.19 808 15.05% 1.74 147
2016 10.43% 2.46 556 3.34% 0.48 593 10.37% 5.38 811 23.91% 3.25 111
2017 16.65% 3.49 506 20.45% 2.67 544 25.03% 13.64 722 11.04% 0.59 82
2018 10.86% 4.32 406 21.65% 3.40 405 3.32% 1.55 553 25.80% 1.87 77
2019 21.04% 9.24 394 7.49% 2.37 610 16.25% 12.07 568 5.00% 3.22 106
2020 4.39% 2.55 499 -0.30% -0.53 499 4.06% 1.41 542 2.65% 1.60 115
2021 10.27% 2.13 312 9.70% 3.66 362 8.66% 4.32 367 1.46% 0.30 102
2022 22.75% 3.83 320 15.98% 2.66 437 14.26% 11.46 420 7.00% 2.10 110
2023 17.78% 1.76 353 3.49% 2.00 339 7.15% 2.71 360 19.31% 6.54 76
Overall 13.88% 6.03 4003 11.27% 3.69 4374 11.39% 5.57 5151 11.35% 3.55 926
year Durbl BusEq Hlth NoDur
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 16.83% 5.77 216 15.66% 6.41 1378 5.75% 2.72 1242 0.68% 0.31 241
2016 4.95% 0.79 209 10.24% 6.09 1303 11.07% 2.90 1185 17.14% 5.18 242
2017 -2.17% -0.30 282 10.36% 3.12 1154 3.37% 2.39 1060 -2.14% -0.68 278
2018 18.29% 5.57 208 0.32% 0.31 1002 12.47% 6.58 918 1.29% 0.13 324
2019 15.31% 2.26 222 10.17% 6.13 1036 10.02% 7.68 884 18.51% 2.30 277
2020 5.73% 0.79 168 8.11% 4.14 1100 2.86% 1.84 950 -0.37% -0.28 265
2021 10.01% 5.31 150 11.43% 1.49 731 7.66% 4.58 592 -6.38% -0.60 195
2022 7.99% 2.58 162 12.08% 4.77 822 12.05% 2.71 850 5.24% 1.11 187
2023 12.05% 2.37 140 10.45% 11.80 962 3.52% 1.32 844 -5.69% -0.71 158
Overall 11.02% 3.24 1757 9.68% 6.77 9488 7.53% 5.71 8525 7.50% 1.38 2167
year Money Telcm Enrgy Utils
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 18.13% 4.41 1153 1.68% 0.00 403 -4.21% -1.45 303 18.47% 3.69 401
2016 7.64% 3.53 1108 -10.18% -1.22 329 15.80% 2.83 252 -1.84% -0.23 310
2017 1.77% 0.46 1071 8.60% 5.16 307 -18.18% -8.75 307 -39.46% -6.99 315
2018 13.55% 8.48 879 10.65% 2.46 219 8.38% 1.29 275 -23.76% -5.53 215
2019 2.10% 2.25 977 1.53% 0.52 351 -10.28% -9.07 340 -2.36% -0.27 193
2020 1.42% 0.76 952 -8.84% -2.83 360 5.72% 2.62 434 -7.29% -2.36 165
2021 2.62% 0.47 614 6.15% 3.27 188 -15.04% -6.83 251 -4.01% -1.16 103
2022 4.34% 2.13 667 12.84% 3.50 143 16.14% 4.44 286 8.31% 2.71 132
2023 -0.63% -0.18 569 8.99% 6.89 182 -27.33% -5.03 290 -6.43% -1.12 120
Overall 5.18% 3.31 7990 3.32% 1.39 2482 1.90% 0.31 2738 -1.97% -0.88 1954
Table A7: Summary Statistics for Alternative Analyst Information Value Estimations
The analyst information value is measured as the explained return variance divided by the price impact, with the results reported in millions of dollars. All dollar values are adjusted for inflation to reflect 2020 values using the Consumer Price Index (CPI). The sample comprises analyst reports targeting common constituent firms of the S&P 100 index from 2015Q1 to 2023Q4. Both mean and standard deviation are computed using the delta method.TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. Intraday denotes that $\sigma_{v}^{2}$ does not include overnight return volatility. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Mean SE 95%CI 99%CI N
Information value (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17669
Information value of text (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17672
Information value of revisions (TAQ volitility), \$M 0.42 0.04 [0.34, 0.50] [0.31, 0.53] 17672
Information value (Intraday), \$M 0.42 0.04 [0.34, 0.51] [0.32, 0.53] 17672
Information value of text (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17669
Information value of revisions (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17672
Information value (EMO), \$M 0.34 0.04 [0.27, 0.41] [0.25, 0.43] 17672
Information value of text (EMO), \$M 0.35 0.04 [0.27, 0.42] [0.25, 0.44] 17672
Information value of revisions (EMO), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17669
Information value (CLNV), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17672
Information value of text (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.41] 17672
Information value of revisions (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.42] 17672
C. Machine Learning Models
Ridge Regression
Ridge regression addresses multicollinearity by adding a regularization term to the least squares objective function. The ridge regression estimator is given by:
[…]
where $\alpha$ is the regularization parameter that controls the trade-off between fitting the data and shrinking the coefficients.
To find the optimal value of $\alpha$, cross-validation is used over a grid of values ranging from $10^{-10}$ to $10^{10}$. The cross-validation process ensures that the chosen model generalizes well to unseen data, preventing overfitting while capturing the predictive power of the text embeddings.
Partial Least Square Regression
To mitigate the risk of overfitting inherent in high-dimensional text embeddings, I employ Partial Least Squares (PLS) for dimensionality reduction.
The optimization problem can be expressed as follows:
[…]
where $\Omega$ is a $K\times P$ transformation matrix that reduces the $K$ predictors in $X_{i}$ to $P$ lower-dimensional components.
The extraction of the $j$-th PLS component is guided by the following objective function:
[…]
In essence, this approach sequentially extracts components that maximize the covariance with the outcome variable, while ensuring orthogonality to previously extracted components.
Extreme Gradient Boosting
Tree-based approaches are commonly applied in stock return forecasting literature (see, e.g., Gu et al., 2020; Cao et al., 2024; Bonini et al., 2023). XGBoost is an advanced implementation of tree-based machine learning models that builds an ensemble of decision trees. In XGBoost, each tree is added sequentially to correct the errors of previous trees. The main idea is to combine the outputs of multiple weak learners (decision trees) to create a strong learner. XGBoost incorporates regularization techniques, such as L1 (Lasso) and L2 (Ridge), to prevent overfitting and enhance model generalization.
In comparison to random forests, which build a multitude of independent trees and aggregate their predictions, XGBoost constructs trees sequentially, with each tree designed to correct the errors of the preceding ones. While random forests rely on bagging, a method that combines the predictions of various trees to reduce variance, XGBoost uses boosting, an approach that aims to reduce both bias and variance by focusing on difficult-to-predict instances in subsequent iterations. This boosting approach allows XGBoost to effectively capture and leverage the nuanced information embedded in textual data, leading to a more accurate estimation of stock returns.
Neural Networks
To extend beyond the linear modeling approach, I explore the use of Neural Networks to estimate $CAR$ using text embeddings derived from analyst reports. A Neural Network can capture complex non-linear relationships between the text embeddings and the $CAR$, potentially improving the model’s fitting capabilities. Consider a three-layer Neural Network as the prediction model. The prediction problem can be formulated as follows:
$f(X_{i};\theta)=W_{3}\sigma\left(W_{2}\sigma\left(W_{1}X_{i}+b_{1}\right)+b_{2}\right)+b_{3},$ (16)
where $\sigma(\cdot)$ is the ReLU activation function, $W_{i}$ and $b_{i}$ represent the weights and biases for layer $i$, respectively. The Neural Network architecture consists of an input layer representing the text embeddings, followed by multiple hidden layers. The specific architecture employed includes 32 neurons in the first hidden layer, followed by 16, 8, 4, and 2 neurons in subsequent layers. This structure is flexible and can be adjusted by adding or removing layers as necessary to optimize performance.
The training process involves optimizing the weights and biases to minimize the loss function, which, in this case, is the mean-square loss. To regularize the model and prevent overfitting, early stopping is implemented, halting training once the validation loss ceases to decrease. Additionally, following Gu et al. (2020), the models are retrained five times, and the final estimation is obtained by averaging the outputs of these five models, forming an ensemble estimation.
D. Theory
I provide an intuition for the measure of strategic value by discussing a simple extension of Kyle (1985) model. In the Kyle model, there is one risky asset with a payoff $\tilde{v}\sim N\left(p_{0},\Sigma_{0}\right)$. Three types of traders exist: a strategic trader with insider information, a market maker who sets prices in a perfectly competitive market, and an uninformed trader who trades $\tilde{u}\sim N\left(0,\sigma_{u}^{2}\right)$, where $\tilde{u}$ is independent of $\tilde{v}$. Illiquidity is measured by Kyle’s lambda ($\lambda$). Kyle’s lambda depends on private information and liquidity trading. I extend the model by considering a case where $\phi$ percentage of variance in $\tilde{v}$ is explained by the informed trader’s information.
A Single Auction Model
There are two periods, $t_{0}$ and $t_{1}$. The asset is traded with asymmetric information at period $t_{0}$, and the value $\tilde{v}$ is realized at period $t_{1}$. I assume without loss of generality that
$\tilde{v}=P_{0}+\tilde{s}+\tilde{\epsilon},$
where $\tilde{s}$ is a mean-zero signal observed by the informed trader at $t_{0}$, and $\tilde{\epsilon}$ is the combination of residual information and noise. I assume that the signal $\tilde{s}$ and residual $\tilde{\epsilon}$ are uncorrelated, that is, $\sigma_{v}^{2}=\sigma_{s}^{2}+\sigma_{\epsilon}^{2}$.
Let
$\phi=\frac{\operatorname{var}(\tilde{s})}{\operatorname{var}\left(\tilde{v}\right)}=\frac{\sigma_{s}^{2}}{\sigma_{v}^{2}}.$
This measure $\phi$ is the ”R-square” of projecting $\tilde{v}$ on $\tilde{s}$, or the percentage of explainable variance in $\tilde{v}$ using signal $\tilde{s}$.
After observing $\tilde{s}$, the informed trader submits a market order $\tilde{x}=X(\tilde{s})$, and the uninformed trader trades a zero-mean random variable $\tilde{z}$ that is normally distributed and independent of $\tilde{v}$. The profit of the informed trader is given by: $\tilde{\pi}=(\tilde{v}-\tilde{p})\tilde{v}$. The insider has a rational guess of $P(\tilde{x}+\tilde{u})$ and understands that his order $\tilde{x}$ will move the price against him.
The market maker observes the order flow $\tilde{y}\stackrel{{\scriptstyle\text{def}}}{{=}}\tilde{x}+\tilde{z}$ and determines the equilibrium price $\tilde{p}=P(\tilde{x}+\tilde{u})$ to break even. The assumptions for the market maker are risk neutrality and perfect competition, which drives the profits for market makers to zero.
Equilibrium
An equilibrium is a set of $X$ and $P$ satisfying
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s],$
$\tilde{p}(X,P)=E[\tilde{v}\mid\tilde{x}+\tilde{u}].$
The first condition is profit maximization, stating that given the market maker’s pricing rule, the insider chooses a strategy $X$ maximizing her conditional expected profit, taking into account the pricing rule. The second condition is market efficiency. Given the insiders’ trading strategy, the market maker sets the price to be the expected value of the security.
Conjecture:
$P(\tilde{y})=\mu+\lambda\tilde{y},$
$X(\tilde{s})=\alpha+\beta\tilde{s}.$
The profit of the insider can be written as:
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s]$ $=E[(\tilde{v}-\mu-\lambda(\tilde{u}+x))x\mid\tilde{s}=s]$
$=(P_{0}+s+-\mu-\lambda x)x.$
Traders take into account that her order flow will move the price against her, which serves to restrain her position size.
Solving for optimal profit, I get:
$x^{*}=\frac{P_{0}+s-\mu}{2\lambda}=\alpha+\beta v.$
Hence, I can express $\alpha$ and $\beta$ :
$\begin{gathered}\beta=\frac{1}{2\lambda},\\ \alpha=\frac{P_{0}-\mu}{2\lambda}=(P_{0}-\mu)\beta.\end{gathered}$
When the market maker puts a higher weight on the order flow in setting the price, the trader puts a lower weight on her information.
I now look at the price-setting rule:
$\mu+\lambda y=E\{\tilde{v}\mid\alpha+\beta\tilde{s}+\tilde{u}=y\}.$
Essentially, the market maker observes a normally distributed signal about
$\mu+\lambda y$ $=E[\tilde{v}\mid y]$
$=\bar{v}+\frac{\operatorname{cov}(\tilde{v},\alpha+\beta\tilde{s}+\tilde{u})}{\operatorname{var}(\alpha+\beta\tilde{s}+\tilde{u})}(\beta\tilde{s}-\beta\bar{s}+\tilde{u})$
$=p_{0}+\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}}(y-\alpha).$
Hence, I can express $\lambda$ and $\mu$ :
$\lambda=\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}},$
$\mu=p_{0}-\alpha\lambda.$
There is a unique linear equilibrium given by
$\mu=P_{0},$
$\lambda=\frac{\sigma_{s}}{2\sigma_{u}},$
$\alpha=0,$
$\beta=\frac{\sigma_{u}}{\sigma_{s}}.$
Discussion
Kyle’s Lambda
The parameter $\lambda$ is universally known as Kyle’s lambda. Formally, it is the impact on the equilibrium price of a unit order. Its reciprocal ($1/\lambda$) measures the liquidity (or depth) of the market. If $1/\lambda=\frac{2\sigma_{u}}{\sigma_{s}}$ is larger, then the market is more liquid, either because there is less private information in $\sigma_{s}$ or there is more liquid trade in the sense of $\sigma_{u}$.
Value of Private Information
Notice that the equilibrium strategy of the informed trader is $\tilde{x}=\beta s$. The unconditional expected gain of the informed trader is
$\mathrm{E}[\tilde{x}[\tilde{v}-p(\tilde{x}+\tilde{z})]]$ $=\beta\mathrm{E}[s(\tilde{v}-\mu-\lambda\beta s-\lambda\tilde{u})]$
$=\beta(1-\lambda\beta)\sigma_{s}^{2}$
$=\frac{\sigma_{s}\sigma_{u}}{2}$
$=\frac{\phi\sigma_{v}^{2}}{4\lambda}.$
The expected gain for the informed trader is maximized when she has more private information or when there is more liquidity trading. On the other hand, liquidity traders incur losses equivalent to the informed trader’s gains, but they accept these losses due to other motives for trading. The equilibrium price ensures that market makers do not profit or lose in expectation.
E. Delta Method Approximation
I use the delta method to approximate the expected value and variance of a ratio of two random variables. This method relies on a first-order Taylor expansion. Specifically, consider two random variables $X$ and $Y$ with means $\mu_{X}$ and $\mu_{Y}$, and variances $\sigma_{X}^{2}$ and $\sigma_{Y}^{2}$, respectively. We are interested in the ratio $Z=\frac{X}{Y}$ and want to approximate the mean $E[Z]$ and the variance $\text{Var}(Z)$.
Delta Method Approximation
The key idea is to approximate the function $g(X,Y)=\frac{X}{Y}$ using a Taylor series expansion around the means $\mu_{X}$ and $\mu_{Y}$. For the function $g(X,Y)$, I use a first-order Taylor expansion around $(\mu_{X},\mu_{Y})$:
$g(X,Y)\approx g(\mu_{X},\mu_{Y})+\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}(X-\mu_{X})+\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}(Y-\mu_{Y}).$
The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are:
$\frac{\partial g}{\partial X}=\frac{1}{Y},\quad\frac{\partial g}{\partial Y}=-\frac{X}{Y^{2}}.$
Evaluating these at $(\mu_{X},\mu_{Y})$, I get:
$\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}=\frac{1}{\mu_{Y}},\quad\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}=-\frac{\mu_{X}}{\mu_{Y}^{2}}.$
Substituting the partial derivatives into the Taylor expansion, I obtain:
$\frac{X}{Y}\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}(X-\mu_{X})-\frac{\mu_{X}}{\mu_{Y}^{2}}(Y-\mu_{Y}).$
Taking the expectation on both sides:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}E[X-\mu_{X}]-\frac{\mu_{X}}{\mu_{Y}^{2}}E[Y-\mu_{Y}].$
Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation simplifies to:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}.$
Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by:
$\operatorname{Var}(Z)\approx\left(\frac{\partial g}{\partial X}\right)^{2}\operatorname{Var}(X)+\left(\frac{\partial g}{\partial Y}\right)^{2}\operatorname{Var}(Y)+2\left(\frac{\partial g}{\partial X}\right)\left(\frac{\partial g}{\partial Y}\right)\operatorname{Cov}(X,Y).$
Substituting the partial derivatives:
$\operatorname{Var}\left(\frac{X}{Y}\right)\approx\frac{\operatorname{Var}(X)}{\mu_{Y}^{2}}+\frac{\mu_{X}^{2}\operatorname{Var}(Y)}{\mu_{Y}^{4}}-2\frac{\mu_{X}\operatorname{Cov}(X,Y)}{\mu_{Y}^{3}}.$
Application to Value of Information
In the context of the provided study, I estimate the mean value of information for a subsample $s$ using the delta method.
Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$. Define the mean and variance over subsample $s$ as:
$\mu_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right),$
and the mean price impact per dollar over subsample $s$ as:
$\mu_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}.$
Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by:
$E\widehat{\Omega}_{s}\approx\frac{\mu_{\nu s}}{\mu_{\lambda s}}.$
The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as:
$\text{Var}\left(\widehat{\Omega}_{s}\right)\approx\frac{1}{\mu_{\lambda s}^{2}}\left(\Sigma_{\nu s}+\frac{\mu_{\nu s}^{2}}{\mu_{\lambda s}^{2}}\Sigma_{\lambda s}-2\frac{\mu_{\nu s}}{\mu_{\lambda s}}\Sigma_{\nu\lambda s}\right),$
where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as:
$\Sigma_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(\widehat{\sigma}_{it}^{2}-\mu_{\nu s}\right)^{2},$
$\Sigma_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\left(\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}-\mu_{\lambda s}\right)^{2},$
$\Sigma_{\nu\lambda s}=\text{Cov}\left(\widehat{\sigma}_{it}^{2},\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}\right).$
This provides a first-order approximation for the mean and variance of the value of information over the subsample $s$ using the delta method.

### v2

@@
 JEL Classification: G11, G14, G24
-Keywords: Sell-side Analysts, Value of Information, Large Language Model, Explainable AI
+Keywords: Sell-side Analysts, Value of Information, Large Language Models, Explainable AI
 “No one ever made a decision because of a number. They need a story.”
@@
 Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data.
-While prior studies have demonstrated the incremental information value of report text, these analyses have predominantly centered on the sentiment dimension (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018).
-However, analyst reports are meant to provide contextual information perceived as at least as important as the quantifiable outputs.
-For example, according to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations.
-Yet, empirical evidence supporting this survey feedback is currently limited, and it remains unclear which specific component of analyst reports holds the most value for investors.
-The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning, capturing more granular and rich information than before.
-This advancement facilitates a systematic analysis of the comprehensive outputs from sell-side analysts.
-For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the ”earnings analysis” topic, the second provides additional contextual information that LLMs can interpret and analyze.
-I quantify the value of information in analyst reports in three general steps: analyst outputs representation, econometric modeling, and value decomposition.
+Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018).
+However, analyst reports are meant to provide contextual information beyond simple sentiment dimension and the context has been perceived at least as important as the quantifiable outputs.
+According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations.
+Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
+The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024).
+For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the “earnings analysis” topic, the latter provides additional contextual information and reasoning that can be interpreted and analyzed by LLMs.
+I quantify the value of analyst information in analyst reports through three general steps: analyst output representation, econometric modeling, and value decomposition.
 Analyst outputs include both quantitative information and qualitative information.
@@
 For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs.
-Each report is mapped into a structured embedding space using the LLMs’ transformer architecture, allowing these to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2023).
-I evaluate the information content of analyst reports using an R-squared measure to quantify how well the information set explains stock returns.
+Each report is mapped into a structured embedding space using the LLMs’ transformer architecture.
+With structured text representation, I evaluate how well analyst information explains stock returns using an R-squared measure, which also serves as the terminology of information content.
 A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings.
@@
 This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns.
-Additionally, I assess the explanatory power of both quantitative and qualitative measures for abnormal stock returns using out-of-sample R-squared values with a recursively expanding window.
-I compile a comprehensive corpus of analyst reports from the Mergent Investext database, merging these written reports with detailed analyst forecast summaries from the I/B/E/S database.
-The final sample comprises 122,252 analyst reports on S&P 100 constituent firms spanning from 2000 to 2023.
-Using this sample, the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information.
+Additionally, I evaluate the out-of-sample explanatory power of quantitative and qualitative measures for abnormal stock returns to assess model stability.
+I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information.
 When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually.
-Economically, a one standard deviation increase in report text favorableness raises the realized CAR by 120 basis points, compared to 100 basis points for forecast revisions.
-These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content winthin analyst report text.
+These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content within analyst report text.
 The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations.
-My analysis reveals that both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions.
+Both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions.
 Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%.
 In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
-Since a notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al.
+A notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al.
 (2014), Kim and Song (2015), Lobo et al.
 (2017), and Barron et al.
-(2017), I examine the information content of analyst reports both within and beyond the earnings announcement window.
-This investigation reveals that the information content of analyst reports peaks in the first week following earnings announcements, supporting Huang et al.
+(2017), An investigation on samples within and beyond the earnings announcement windows reveals that the information content of analyst reports peaks in the first week following earnings announcements.
+The finding supports Huang et al.
 (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients.
-Using separately trained models on samples within and beyond earnings announcement windows, I find that the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods.
-This significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
-To address potential concerns that analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models.
-By doing so, I can isolate the unique information of analyst reports.
-The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights, highlighting the different information sets provided by the two sources.
+Using separately trained models on samples within and beyond earnings announcement windows, the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods.
+The significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
+Analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models to isolate the unique information from analyst reports.
+The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights.
 As raised in the review paper of Bradshaw et al.
 (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question.
-To quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports.
+To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports.
 The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers.
 This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
-To translate the statistic measure of R-squared into economic terms, I quantify the dollar value of analyst information to strategic investors using Kadan and Manela (forthcoming)’s framework, which builds on Back et al.
-(2000) imperfect competition insider trading model.
-The information value can be interpreted as the total expected profits for strategic investors either because some investors have early access to analyst views or because they conduct independent research and derive insights comparable to those of analysts.
-This multi-client framework provides a realistic foundation for analyzing the complex dynamics of information dissemination in analyst reports.
+So far, I have measured information value statistically with R-squared.
+To translate it into economic terms, I adopt Kadan and Manela (forthcoming)’s framework to quantify the dollar value of analyst information to strategic investors.
+The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts.
+This multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination.
 Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
@@
 Notably, these values increase for large stocks, bold analysts, and reports released promptly following earnings announcements.
-This paper relates to three strands of literature.
-First, This paper relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024).
+My paper relates to three strands of literature.
+This paper first relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024).
 Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights.
-In contrast, my work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations.
+My work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations.
 A direct comparison with the Naive Bayes tone classification used in Huang et al.
 (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone.
-To quantify these gains, I employ a BERT-based tone measure, which increases the out-of-sample $R^{2}$ significantly by 2%.
-Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure to above 10%.
-The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information, offering a more nuanced and comprehensive understanding of analyst narratives.
+Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure from 3% to above 10%.
+The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information.
+Topic decomposition analysis offers a new dimension for understanding analyst information value by revealing their role in interpreting financial data.
 Second, this work contributes to the growing field of machine learning and artificial intelligence in finance.
@@
 This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI.
-This approach enables a precise assessment of the relative importance of the topics discussed in analyst reports.
 Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010).
@@
 (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%.
-This paper extends prior work by quantifying the economic dollar value of such information, accounting for trading costs.
+This paper extends prior work by quantifying the economic dollar value of such insider information.
 The rest of the paper is organized as follows.
@@
 2 Methodology
-In this section, I introduce the large language models used for topic classification and text embeddings, the information content estimation of out-of-sample $R^{2}$, the strategic information value estimation of analyst reports, and the decomposition metric used to evaluate the information content of specific topics.
+In this section, I introduce how I employ large language models for topic modeling and text embeddings, the information content estimation, the strategic information value estimation, and the decomposition metric used to evaluate the information value of topics.
 [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
 2.1 Large Language Models
-Large Language Models such as Bidirectional Encoder Representations from Transformers (BERT), Generative Pre-trained Transformer (GPT), and Large Language Model Meta AI (LLaMA) are trending in literature for understanding textual data.
-These models represent a paradigm shift in NLP, offering capabilities that go beyond traditional methods like word-to-vec (W2V) or Latent Dirichlet Allocation (LDA) topic models.
-In this study, I harness the power of these advanced generative AI models to conduct a comprehensive analysis of analyst reports.
+I harness the power of advanced generative AI models to conduct comprehensive analyses of analyst reports.
 The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors.
-This approach not only enriches information extraction but also provides a nuanced understanding of subtle linguistic cues and thematic elements that may influence market reactions.
 2.1.1 ChatGPT Prompting
-To effectively categorize the topics discussed in analyst reports, I employ ChatGPT prompting to extract exclusive and consistent topics and avoid ad-hoc classification.
-By leveraging AI, my goal is to identify the primary topics covered in these reports and assign each sentence to a single, specific topic category.
-This approach ensures a systematic and coherent classification of the content within analyst reports.
-In the first step, to derive manageable and meaningful topics, I input a random sample of 100 analyst reports into ChatGPT, requesting an analysis of the content to identify high-level categories.
+To avoid ad-hoc classification, I employ ChatGPT prompting to extract exclusive and meaningful topics.
+The goal is to identify primary topics covered in analyst reports and assign each sentence to a single topic category.
+In the first step, I use a random sample of 100 analyst reports and request ChatGPT to analyze the content and identify high-level categories with the following prompt:
 Prompt 1:Please read the provided text file of sell-side analyst reports carefully.
@@
 Here is the report content: {text}.
-It ends up with 13 categories defined as follows.
+It ends up with 13 categories defined as follows:
   • Executive Summary: Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
@@
 In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences.
-To begin with, a random sample of 17,028 sentences from 100 analyst reports is selected from analyst reports.
-I employ the GPT-4-Turbo model through the ChatGPT API to categorize each sentence in this sample into a single, relevant topic.
-Prompt 2: Please read the following sentence from the sell-side analyst report of the company {firm} ({ticker}) carefully.
+To begin with, I employ the GPT-4-Turbo model to categorize each sentence in a random sample of 17,028 sentences from 100 analyst reports into a single, most relevant topic.
+Prompt 2: Please read the following sentence from a sell-side analyst report of the company {firm} ({ticker}) carefully.
 Determine which category of information it belongs to in the following 17 categories: {categories}.
@@
 Here is the sentence from the analyst report: {sentence}.
-This training sample serves as the foundation for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus.
+This training sample serves as the training sample for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus.
 The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample.
@@
 These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics.
-Conversely, Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
+Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
 Several factors contribute to these observed frequencies.
@@
 The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level.
-Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in focus on these issues.
-2.1.2 Text Embeddings
+Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in ESG discussions.
+2.1.2 Text Representations
+A systematic textual analysis requires a structured representation of analyst reports.
 Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps.
@@
 This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand.
-[^2] BERT uses 30,000 tokens, OpenAI GPT models use around 50,257 tokens, and LLaMA models use approximately 32,000 tokens.
 To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”.
@@
 At the heart of LLMs lies the transformer architecture.
-Introduced by Vaswani et al.
-(2017), the transformer architecture revolutionized the field of NLP with its novel approach to sequence-to-sequence tasks.
-Unlike recurrent neural networks (RNNs) and long short-term memory networks (LSTMs), transformers use a self-attention mechanism that processes input sequences in parallel.
-This approach allows the model to weigh the importance of different words in relation to each other, effectively capturing context regardless of word distance.
-The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance in various NLP tasks.
-Text embedding, a core output of the transformer architecture, transforms language into a format suitable for computational analysis.
+The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance.
+A core output of the transformer architecture is text embeddings.
 These embeddings are dense, continuous vectors that capture semantic and syntactic information.
-In models like BERT, GPT, and LLaMA, self-attention mechanisms generate these embeddings, encoding contextual information for each token.
 Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia.
-The embedding dimensions vary: BERT uses 768 or 1,024, GPT-2 Small uses 768, and LLaMA-2-13B model uses 5,120.
 These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
 To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI.
+The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.
 The tokenizer processes the text data, converting it into tokens that the model can understand.
 By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports.
-These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and text classification.
-The reason I employ LLaMA2 as a baseline Language Model is due to its 4,096-token input limit, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.
+These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and economic modeling.
 2.2 Information Content
-I use out-of-sample $R^{2}$ as a measure for the information content of analyst reports.
+Following Asquith et al.
+(2005), I use $R^{2}$ as a measure for information content.
 To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
@@
 Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ as
-$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum CAR_{[-1,+1],it}^{2}},$ (3)
-where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation of the model.
+$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum_{(i,t)\in\mathcal{T}}CAR_{[-1,+1],it}^{2}},$ (3)
+where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation.
 $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
@@
 2.3 Information Value
-I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal based on the theoretical framework of Back et al.
-(2000):
+I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal:
 $\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
-where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, multiplied by the closing price of the day prior to the [t-1, t+1] window.
-[^3] Due to risk neutrality, the equilibrium of the continuous-time Kyle model holds if $\tilde{v}$ is an unbiased signal of the asset price.
-The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
-Specifically, $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return.
+where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price of day t-2.
+$\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return.
 $N$ denotes the number of analyst reports about stock $i$ on day $t$.
-To estimate the price impact $\lambda_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$).
+To estimate the price impact $\widehat{\lambda}_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$).
 Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals.
@@
 (2007) are also used in robustness tests.
-The regression model used is
+The regression model is
 $r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
-where $\widehat{\lambda}_{it}$ represents the estimated price impact (or Kyle’s lambda).
-In this context, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$.
-The interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$.
+where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda).
+In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$.
+The 3-day interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$.
 Thus, each regression has $1170(=390*3)$ observations.
+Essentially, the measure reflects the ratio of return volatility to price impact.
+It is derived based on the theoretical framework of Back et al.
+(2000), which is a continuous-time version of Kyle (1985) model.
+Due to risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of the asset price.
+The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
 I employ the measure for two key reasons.
 First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information.
-This setting is particularly apt for analyzing the value of analyst reports, as it accounts for the correlated nature of information and the strategic interactions among market participants.
 Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
+[^3] Kadan and Manela (forthcoming) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply to the analyst case.
 Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced.
-To tackle this issue, I apply the delta method in estimating the mean and variance of information value over sample $s$.
+To tackle this issue, I apply the delta method to estimate the mean and variance of information value over sample $s$.
 The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
 2.4 Shapley Value Decomposation
-To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology, as detailed by Lundberg and Lee (2017) and applied to financial contexts by Chen et al.
+To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology.
+SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves.
+It was later applied for machine learning model interpretation by Lundberg and Lee (2017) and Chen et al.
 (2022).
-SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves.
-In my context, this allows me to attribute the overall explanatory power of an analyst report to its constituent topics.
-Leveraging the additive nature of text embeddings, I design a novel decomposition method based on SHAP values.
-This approach enables me to systematically assess the relative importance of various topics in analyst reports.
-First, LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
+In my context, the approach allows me to attribute the overall explanatory power of an analyst report to its constituent topics.
+Leveraging the additive nature of text embeddings, I design a novel decomposition method to systematically assess the relative importance of various topics in analyst reports.
+LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
 $y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
@@
 where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
-One concern in the architecture of a transformer model is that each token in the input sequence can pay attention to other words through a mechanism called “self-attention” or “scaled dot-product attention.” This mechanism allows the model to weigh the importance of each word relative to every other word in the sequence, thereby capturing contextual information.
-Consequently, the embedding of a word contains information from other words, which may pertain to different topics.
-The embedding of the word might predominantly reflect the context information from other topics.
-To isolate the self-attention across topics, I process the embedding for each sentence in analyst reports and take the token-weighted mean as the report embedding.
+One concern raised by the nature of transformer models is that each token in the input sequence can pay attention to other tokens through a mechanism called “self-attention” or “scaled dot-product attention.” This architecture allows the model to weigh the importance of each token relative to every other token and capture contextual information.
+For example, the embedding of token $i$ contains information from token $j$, which may pertain to a different topic.
+To isolate the self-attention across topics, I process the embeddings sentence by sentence to ensure the token embeddings only reflect the context information within the sentence.
+I then take the token-weighted mean of sentence embeddings as the report representation, which is referred to as sentence-segment embeddings.
 $y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
 where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
-The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context in analyst reports.
-Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Empirically, the Shapley value can be written as:
+The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context across sentences.
+Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Specifically, the Shapley value can be written as:
 $\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
@@
 I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation.
-This approach fairly distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions to analyst reports.
+This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions.
 3 Data and Empirical Results
 This section presents the empirical analysis of analyst report information.
-I begin by examining the information content of these reports and decomposing it across different topics.
+I begin by examining the information content and decomposing it across different topics.
 I then quantify the strategic dollar value of analyst information.
-Cross-sectional analyses and robustness checks are conducted to validate the main findings.
+The results emphasize analysts’ income statement analyses following earning announcements.
 3.1 Data
@@
 I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023.
-The initial reports are in PDF format, and I use Adobe Acrobat to extract structured text from each report.
-The segmented text is chunked into sentences.
-To get clean content of analyst discussions, I train a BERT model to remove the boilerplate segments from each report.
+The initial reports are in PDF format, and I use Adobe Acrobat to extract plain text from each report.
+The text is chunked into sentences.
+To get clean content of firm-relevant discussions, I fine-tune a BERT model to remove the boilerplate segments.
 [^5] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions.
@@
 (2024), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports.
-The BERT model trained for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
-The Mergent Investext Database provides information on analyst names and the release date for each report.
+The BERT model fine-tuned for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
+The Mergent Investext Database provides information on analyst names and the release date of each report.
 To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al.
@@
 I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database.
-Following Huang et al.
-(2014), I define the matching window for each report as 2 days before the announcement date to 2 days after the revision date, accounting for multiple reports released between these dates.
+Since multiple reports could be released between a forecast announcement and its revision, I follow Huang et al.
+(2014) and define the matching window for each report as 2 days before the announcement date to 2 days after the revision date.
 [^6] In I/B/E/S, each analyst forecast for a specific firm has two time stamps.
@@
 The revision date (REVDATS) is the date the forecast becomes invalid or is revised.
-Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and this time window.
-The numerical measures include earnings forecasts, stock recommendations, and price target information from analyst reports.
+Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window.
 I also compile firm characteristics from the CRSP and Compustat databases.
-[^8] See Table A2 and Table A3 for detailed definitions and summary statistics.
-After excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 firms from 1,305 analysts across 140 unique brokerage firms.
-Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts for each year, or across the Fama-French 12 industries.
-It also reports the average number of pages and tokens per report, which are important indicators of report lengths.
+[^7] See Table A2 and Table A3 for detailed variable definitions and summary statistics.
+After the matching procedure and excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 stocks from 1,305 analysts across 140 unique brokerage firms.
+Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts across years and the Fama-French 12 industries.
+It also summarizes the average number of pages and tokens per report, which are important indicators of report lengths.
 Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al.
@@
 This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
-The daily stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023.
+The stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023.
 To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
-In Section 3.2.2, I explore textual information from earnings conference calls and subsequent analyst reports.
-The earnings call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005101010Available at https://seekingalpha.com/earnings/earnings-call-transcripts.
-These transcripts provide comprehensive accounts of the calls, including executive presentations and Q&A sessions with analysts, offering valuable insights into manager-analyst interactions.
-I sampled earnings call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat.
+Given my goal to estimate the information value of analyst outputs, I need to control for other information available to investors.
+In Section 3.2.2, I compare textual information from earnings conference calls and subsequent analyst reports.
+Earnings conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Available at https://seekingalpha.com/earnings/earnings-call-transcripts..
+These transcripts provide comprehensive record of the calls, including executive presentations and manager-analyst interactions like Q&A sessions.
+I sample earnings conference call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat.
 The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks.
 This process yielded a final sample of 7,909 firm-quarter-level conference calls.
-Given my goal to estimate the information value of analyst outputs, I need to define the information set available to investors.
-The numerical information and variables used as inputs to machine learning models are as follows:
+The set of numerical variables used as input in machine learning models includes:
 Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
@@
 The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI).
-$Fluidity$ measures the firm’s product market fluidity.
+$Fluidity$ measures firms’ product market fluidity.
 Industry-Level Measures: $IndustryRecession$ is an indicator variable equal to one if the firm’s industry (defined by Fama-French 48 classification) experiences a negative monthly return in the bottom quintile.
@@
 3.2 Information Content of Analyst Reports
-In this section, I provide yearly $R^{2}_{\mathrm{OOS}}$ of analyst reports for explaining contemporaneous cumulative abnormal returns to examine the information content of analysts’ quantitative and qualitative outputs.
-Furthermore, I attribute the information content of analyst report text across topics using Shapley value decomposition.
+In this section, I examine the information content of analysts’ quantitative and qualitative outputs by projecting contemporaneous cumulative abnormal returns on analyst information.
+Furthermore, I attribute the information content of analyst report text across topics using the Shapley value decomposition introduced in Section 2.4.
 3.2.1 Information Content of Quantitative and Qualitative Information
-Table 2 compares the information content embedded in numerical and textual measures, quantified through their $R^{2}_{\mathrm{OOS}}$.
-The estimation model used in this table is the Ridge regression, where the objective variable is $CAR_{[-1,+1]}$ centered at the analyst report release date.
-For reports released on non-trading days, I adjust t=0 to the next trading day as recorded in the CRSP database.
-The estimation methodology employs an expanding window approach, constructing training sets from 2000 to the year preceding each target year.
-I estimate separate Ridge regression models for each input configuration, conducting yearly out-of-sample estimations of $\widehat{CAR}_{[-1,+1]}$, allowing for dynamic assessment of model performance over time.
-Panel A of the table delineates the performance of four distinct input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) comprehensive numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, the 5,120-dimensional full-context representations of each report generated by LLaMA-2-13B model; and (4) a hybrid input combining revision measures with textual embeddings.
-Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for the revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions.
+Table 2 compares the information content embedded in numerical and textual measures, quantified with $R^{2}_{\mathrm{OOS}}$.
+The estimation model applied in this table is the Ridge regression specified in Equations 1 and 2.
+The objective variable is ${CAR}_{[-1,+1]}$ centered at the analyst report release date.
+For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database.
+I estimate separate Ridge regression models for distinct input configurations and employ an expanding window approach, constructing training sets from 2000 to the year preceding each target year.
+Panel A of the table delineates the performance of four types of input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
+Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% with forecast revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions.
 The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features.
 I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4).
-The Diebold-Mariano test follows a standard normal distribution under the null hypothesis of equal predictive accuracy between two competing forecasting models, allowing similar interpretations of t-statistics as in regression analysis (Gu et al., 2020).
-The models incorporating revision and numerical inputs significantly outperform a zero-estimation benchmark.
-Columns (5) and (6) elucidate the performance of text embedding input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20.
-This underscores the significant information content of the report text in explaining contemporaneous stock returns.
+The models with revision and numerical inputs significantly outperform a zero-estimation benchmark.
+Columns (5) and (6) elucidate the performance of Ridge models with text embeddings as input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20, underscoring the significant information content of the report text in explaining contemporaneous stock returns.
 The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession.
+The DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, suggesting substantial information content in analyst report text compared with forecast revisions.
 Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs.
@@
 Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period.
-The Diebold-Mariano test comparing revision-only and text-only inputs yields a t-statistic of 1.66, highlighting the substantial information content in the report text.
-In columns (7) and (8), I further investigate the performance of combining both revision and text input in the Ridge regression.
+In columns (7) and (8), I further investigate the model performance when combining both forecast revisions and text embeddings in the Ridge regression.
 The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model.
@@
 The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$.
-The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by market investors.
-To address the concern that the explanatory power might be derived from quantitative information in analyst reports, I conduct robustness checks using content with numbers removed.
-In Table A4, I exclude numerical data from the written reports before generating text embeddings.
-The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, still significantly higher than that of quantitative revision measures (t-statistic = 2.40).
-This suggests that the superior performance of the text-based model is not driven by numerical information within the reports.
-LLMs are known to exhibit improved language skills as the size of their parameters increases.
-I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI GPT-3 embedding model, and LLaMA-3 model.
+The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by investors.
+The explanatory power might be derived from quantitative information in the numbers mentioned in report text.
+I conduct robustness checks using content with numbers removed.
+To address the concern, I exclude numerical data from the written reports before generating text embeddings and report model performance in Table A4.
+The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, significantly higher than that of quantitative revision measures (t-statistic = 2.40).
+The robustness check suggests that the superior performance of text-based models is not driven by the numbers embedded in text.
+LLMs exhibit improved language skills as the parameter size increases.
+I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI text-embedding-3-small model, and LLaMA-3 model.
 The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models.
-Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between each model and the LLaMA-2 model.
-The input in Panel B is report text, comparable to column (3) in Panel A.
+Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 model.
+The input in Panel B is report text embeddings, comparable to column (3) in Panel A.
 A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark.
-Overall, the BERT and OpenAI models underperform the LLaMA-2 model at the 1% level for each year.
+Overall, the BERT and OpenAI models underperform the LLaMA-2 model.
 The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model.
 The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs.
-The pairwise comparisons between LLMs also shed light on how the scale of the model influences its performance capabilities.
+Pairwise comparisons between LLMs also shed light on how the scale of the model influences performance capabilities.
 To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al.
@@
 In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023.
-In Figure A3, I plot the information value of analysts’ information by industry, revealing a pattern consistent with the information content results.
 In Table A6, I present the information content of analyst report text across the Fama-French 12 industries.
@@
 The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value.
-To assess economic magnitude, I perform OLS regressions as specified in Equation (10):
+In Figure A3, I plot analyst information value by Fama-French 12 industries, revealing a pattern consistent with the information content results.
+To assess the economic magnitude, I perform OLS regressions specified as follows:
 $CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (10)
-This model incorporates the out-of-sample expanding estimation of $CAR_{[-1,+1]}$ using report text embeddings as input.
+This model incorporates the out-of-sample estimation of $CAR_{[-1,+1]}$ using report text embeddings or forecast revisions as input.
 For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014.
@@
 (2014).
-Column (2) reveals that $\widehat{CAR}{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2.
-In column (3), the inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% to 14.9%, underscoring the incremental value of text information.
-The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by quantitative analyst revisions.
-Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the three-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
+Column (2) reveals that $\widehat{CAR}_{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2.
+The inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% in column (1) to 14.9% in column (3), underscoring the incremental value of text information.
+The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by forecast revisions.
+Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the 3-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
 In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics.
-The $\widehat{CAR}{rev}$ model alone achieves an $R^{2}$ of 8.9%, while the combined input model improves this to 14.7%.
-Economically, a one standard deviation increase in $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively.
+The $\widehat{CAR}_{rev}$ model alone achieves an $R^{2}$ of 8.9%, while combined input improves it to 14.7%.
+Economically, a one standard deviation increase in $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively.
 These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
@@
 The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged.
-This stability in coefficients suggests that the market-valued analyst information is distinct from, and not merely chasing back on, past returns or recent news events.
+The stability of the coefficients suggests that market-valued analyst information is distinct from recent earnings news and is not simply a reaction to past returns.
 3.2.2 Information Content around Earnings Announcements
-Some studies have emphasized the role of analysts surrounding corporate earnings announcements (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information events reinforce each other in sequence.
-I, therefore, further study the information content of sell-side analyst reports around this most significant corporate event.
+Some studies have emphasized the relationship between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information sources reinforce each other in sequence.
+I examine the information content of sell-side analyst reports around earnings announcements, the most significant corporate events.
 Following Chen et al.
@@
 Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week.
-A finding consistent with Huang et al.
-(2018) is that an overwhelming number of analyst reports are released immediately following the earnings announcements.
 Consistent with Huang et al.
-(2018), a large number of analyst reports are released immediately after earnings announcements.
+(2018), a large portion of analyst reports are released immediately after earnings announcements.
 Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17.
-This performance drops sharply in the second week, turns negative in the third and fourth weeks, and does not recover thereafter.
-The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports during earnings announcement weeks.
-To account for potential temporal variation, I train Ridge models using separate samples for periods surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports relative to earnings announcements.
+This performance drops sharply in the second week, turns negative in the third and fourth weeks.
+The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports surrounding earnings announcements.
+To account for potential temporal variation, I train Ridge models using separate samples of analyst reports released surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports.
 Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows.
-The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained outside these periods.
-Additionally, the $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen.
-This trend suggests that analysts react swiftly to earnings announcements, with the information content of their reports declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week.
-One potential concern in assessing the value of analyst reports is that analysts might merely be repeating the content of earning announcement calls, with the earnings announcement itself driving market reactions.
+The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained beyond these periods.
+The $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen, declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week.
+This trend suggests that the market values information from analysts who react swiftly to earnings announcements.
+One potential concern is that analysts may simply restate information from earnings calls, and that market reactions are driven by the earnings announcements themselves rather than analyst reports.
 If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls.
-To address this issue, I use earnings conference call transcripts data to differentiate the information content between the two text sources.
-The analysis focuses on reports released within a one-day window following earnings announcements, as this is when analysts are most likely to piggyback on the announcements.
-In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcripts as input.
+To address this issue, I use earnings conference call transcripts to differentiate the information content between the two text sources.
+I analyze reports released within one day after earnings announcements, when analysts are most likely to simply restate earnings information.
+In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcript embeddings as input.
 [^10] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C.
@@
 The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16.
-with a Diebold-Mariano (DM) t-statistic of 5.16.
-Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods of both texts.
+Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods.
 The input comprised the mean of all report text embeddings on the day following the earnings announcement.
-This firm-time level Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
-Incorporating both analyst report embeddings and earnings conference call embeddings in the Ridge regression resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3).
+The Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
+Incorporating both analyst report embeddings and earnings conference call embeddings in Ridge regressions resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3).
 The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports.
@@
 3.2.3 What Content in Analyst Reports is Valued by Market Investors?
-Given the comprehensive discussion in equity reports, what specific content in analyst reports is the most valuable to the market?
-I employ the Shapley value approach in Section 2.4 to attribute the $R^{2}_{\mathrm{OOS}}$ of the full text to 17 topics.
-In this section, I use the token-weighted average of each sentence embedding as the report embedding.
-This sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another.
+After showing that information in report text is significant, a key question arises: what specific content in analyst reports is the most valuable to the market?
+To systematically evaluate topic importance, I employ the Shapley value approach to attribute the text-based $R^{2}_{\mathrm{OOS}}$ to the 17 topics described in Section 2.1.1.
+I use sentence-segmented embeddings as report representation.
+The sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another.
 The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings.
-This reduction roughly quantifies the magnitude of contextual information content in analyst reports.
+This reduction roughly quantifies the magnitude of contextual information content across sentences.
 Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations.
-These reductions are then averaged into a single importance measure for each topic.
-By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring a comprehensive attribution of explanatory power.
-Figure 2 ranks the Shapley values of 17 topics.
-Topics are ordered from left to right, with the most important topics on the left and the least important on the right.
-Each topic is represented by a bar, with a color gradient (from lightest to darkest) indicating the magnitude of its importance in terms of Shapley values.
+These reductions are then averaged over the possible different permutations into a single importance measure.
+By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring an efficient attribution of explanatory power.
 The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports.
-The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each comprising 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$.
+The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each contributing to 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$.
 Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively.
@@
 The robustness of these findings is further reinforced by additional analyses.
-As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses, underlining the consistency of the results.
-Figure A5 in the Online Appendix reveals how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time.
-The annual variations align with the full-context embedding results in Table 2, exhibiting a consistent story across different text embeddings.
+As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses.
+Figure A5 reveals consistent topic rankings across industries.
+Figure A6 in the Online Appendix shows how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time.
 A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy.
-Despite this volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable, indicating that analyst discussions of income statements consistently provide valuable information, even during market disruptions.
-Figure A6 further supports this by showing a consistent pattern of key topics across industries.
-This suggests that the market values Income Statement Analyses and Financial Ratio Analyses universally, beyond specific sectors.
-To investigate whether this topic ranking is driven by analysts’ expertise in certain topics (supply effect) or investors’ differential responses to different topics (demand effect), Figure A7 compares Shapley values across analysts with varying experience levels and from different-sized brokerage firms.
-The remarkable consistency in topic importance across all four subsamples (experienced vs.
+Despite the volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable.
+Analyst discussions of income statements consistently provide valuable information, even during market disruptions.
+The relative importance of topics could stem from analysts’ expertise in certain areas (supply effect) or from investors’ varying responses to different topics (demand effect).
+Figure A7 compares Shapley values across analysts of different experience levels and brokerage firm sizes.
+The consistency in topic importance across all four subsamples (experienced vs.
 inexperienced analysts, large vs.
-small brokerages) eliminates the concern that the topic ranking merely reflects analysts’ varying skill levels in different topics.
+small brokerages) suggests that topic rankings are not driven by differences in analysts’ topical expertise.
 Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements.
-Drawing inspiration from the literature, particularly Huang et al.
+Drawing inspiration from Huang et al.
 (2018) and Blankespoor et al.
 (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
-To facilitate this analysis, I further categorize sentences within the Income Statement Analyses topic into two sets of sub-topics: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income.
-This categorization is achieved through a sophisticated prompt-based classification system, as detailed in Prompt 3:
+To facilitate the analysis, I categorize Income Statement sentences based on two criteria: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income.
+This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
 Prompt 3: Please read the following sentence from a sell-side analyst report for the company {firm} ({ticker}) and classify it based on two criteria:
@@
 Sentence to classify: {report_sent}.
-Figure 3 offers a detailed breakdown of the $R^{2}_{\mathrm{OOS}}$ across 18 topics, with the Income Statement category further divided into sub-topics.
-Specifically, the income interpretation subtopic contributes approximately three times more to the $R^{2}_{\mathrm{OOS}}$ than income acquisition, underscoring the market’s preference for context and explanation over mere data reporting.
-Additionally, the income realization subtopic also contributes about three times more to the $R^{2}_{\mathrm{OOS}}$ than income expectation, indicating a stronger market preference for analysis of actual financial results over forward-looking projections.
-These findings suggest that the market values analysts’ interpretation of realized income to a greater extent than their ability to simply report public information or make expectations about future performance.
+Figure 3 provides a detailed breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, with Income Statement subcategories.
+Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw data.
+The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on analysis of actual results than on forward-looking projections.
+These patterns demonstrate that markets particularly value analysts’ interpretation of realized income, more than their reporting of public information or future expectations.
 This insight is reinforced by the findings in section 3.2.2, which show that analyst reports issued promptly after earnings announcements are the most informative.
@@
 Consequently, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
-Table 5 summarizes the out-of-sample performance of numerical and textual inputs in both reiteration and revision samples.
+Table 5 summarizes the out-of-sample performance of models with numerical and textual inputs using both reiteration and revision samples.
 Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions.
-When recommendations are revised, the explanatory power of various information sources becomes particularly pronounced.
+Earnings forecast revisions are the most frequent, comprising 56.57% of the reports.
+In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively.
+When stock recommendations are revised, the explanatory power of various information sources becomes particularly pronounced.
 The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%.
 The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%.
-The combination of both quantitative and qualitative information sets further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%.
-This finding aligns with the literature that attributes large stock price changes to analyst recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
-Target price revisions, considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), also reveal a clear contrast between revisions and reiterations.
-While target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ of 20.88%, t-stat = 11.18).
-Earnings forecast revisions are the most frequent, comprising 56.57% of the reports.
-In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively.
+The combination of both quantitative and qualitative information further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%.
+This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
+Target price revisions are considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003).
+While reports with target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
 As Kothari et al.
 (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output.
-Interestingly, the $R^{2}_{\mathrm{OOS}}$ for report text in earnings forecast reiteration reports is negative, indicating no explanatory power for contemporaneous stock returns.
-However, when focusing on analyst earnings forecast revisions,, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant.
-The combination of these inputs yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
+For models with text from earnings forecast reiteration reports as input, the $R^{2}_{\mathrm{OOS}}$ is insignificant.
+When focusing on analyst earnings forecast revisions, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant.
+The combination yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
 Collectively, Table 5 shows that report content is more informative when it accompanies forecast revisions rather than reiterations.
@@
 Huang et al.
-(2014) emphasize the importance of analyst report text by extracting tone using a Naive Bayes machine learning approach.
-My study extends this research by incorporating more advanced techniques, including BERT-based tone measures and text embeddings.
-I construct report-level tone measures using both the Naive Bayes approach and BERT classification.
+(2014) emphasize the information content of analyst report text by extracting tone using a Naive Bayes machine learning approach.
+I compare report-level tone measures derived from both Naive Bayes and BERT classification.
 Following Huang et al.
-(2014), I manually label 10,000 sentences as a training sample for both models.
-Each sentence is categorized into positive, neutral, or negative using these pre-trained models, then average the tone at the report level to derive $Tone\_NB$ and $Tone\_BERT$ measures.
+(2014), I manually label 10,000 sentences as training data for both models.
+Each sentence is categorized as positive, neutral, or negative using these pre-trained models, and I average the tone at the report level to create $Tone_{N}B$ and $Tone_{B}ERT$ measures.
 I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
@@
 (2014).
-A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, which is comparable with the 41 basis points increase in two-day abnormal return in Huang et al.
+A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, comparable with the 41 basis points increase in two-day abnormal return reported in Huang et al.
 (2014).
@@
 Both coefficients of tone are significantly positive.
-Column (3) reports regression results with firm-fixed and year-fixed effects, consistent with the findings in column (2).
-Columns (4)-(6) present regression results with BERT-based tone measures as explanatory variables.
+Regression results with firm-fixed and year-fixed effects are consistent, as reported in column (3).
+Columns (4)-(6) summarizes regression results with BERT-based tone measures as explanatory variables.
 The coefficient estimates’ magnitude increases dramatically, along with significance levels.
 Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points.
-When both tone measures are included in an OLS regression in column (7), Naive Bayes-based tone measure coefficients become insignificant, suggesting BERT-based measures capture sentiment more accurately and effectively.
-Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of extracted tone measures and embedding representations.
+When both tone measures are included in an OLS regression in column (7), coefficient estimates of Naive Bayes-based tone measure become insignificant, suggesting that BERT model captures sentiment more accurately and effectively.
+Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of tone measures and embeddings.
 The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows.
-The $R^{2}_{\mathrm{OOS}}$ increases from 9.49% to 12.27% after incorporating analyst revision and text embeddings, confirming the superior performance of state-of-the-art LLMs across various NLP tasks.
-The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The $R^{2}_{\mathrm{OOS}}$ difference between these models is not statistically distinguishable from zero for both Naive Bayes-based and BERT-based tone measures.
-In general, the empirical results in Table 6 support two key findings regarding the superiority of LLMs compared to conventional NLPs.
-First, overall report text tone measures extracted using small-size LLMs are more explanatory for contemporaneous stock returns than those extracted using the Naive Bayes algorithm.
-Second, tone measures (extracted with both Naive Bayes and BERT) add no incremental value to text embeddings, suggesting that LLMs provide more comprehensive contextualized representations.
+The $R^{2}_{\mathrm{OOS}}$ improves to 9.49% and 12.27% after incorporating forecast revisions and text embeddings.
+The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The difference between the $R^{2}_{\mathrm{OOS}}$ of models with and without tone measures is not statistically distinguishable from zero.
+The empirical results in Table 6 highlight two key advantages of LLMs over conventional NLP methods.
+First, tone measures extracted using LLMs better explain contemporaneous stock returns than those from the Naive Bayes algorithm.
+Second, neither BERT nor Naive Bayes tone measures add incremental value beyond text embeddings, suggesting that LLMs capture richer contextual information than simple tone metrics.
 3.2.6 Information Content with Alternative Machine Learning Models
@@
 Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$.
-To ensure out-of-sample evaluation, I train the models yearly and estimate for the following year’s sample.
-The $R^{2}_{\mathrm{OOS}}$ is calculated by averaging the estimated and realized $CAR$ values.
-Additionally, I perform pairwise Diebold-Mariano tests using a zero benchmark to assess predictive accuracy.
-I begin with PLS regression, summarized in Table 7.
-PLS is notable for extracting the most relevant common components from sentiment proxies that align with the target variable (Huang et al., 2015).
-If there is a strong common component in the text embeddings that is informative of stock returns, the PLS regression architecture should effectively capture this relevant information.
-Examining the model performance, I observe that PLS regressions produce an overall $R^{2}_{\mathrm{OOS}}$ of 8.26%.
-This measure of information content ranges from 2.14% in the 2021 sample to 12.23% in the 2015 sample, showing a similar pattern across years as seen in Ridge models.
-Diebold-Mariano tests indicate the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all periods, demonstrating strong evidence of information content in the analyst report text.
+Table 7 presents results from PLS regressions.
+PLS effectively extracts common components that correlate with the target variable.
+If text embeddings contain strong common components that predict stock returns, PLS regression should capture this information.
+The PLS regressions achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to those in Ridge models.
+DM tests confirm that the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all years.
 Next, I replicate the analyses with the XGBoost algorithm.
@@
 Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable.
-As summarized in Table 7, the XGBoost algorithm underperforms with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76).
+As summarized in Table 7, the XGBoost algorithm underperforms Ridge regressions with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76).
 The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
-Neural Networks, forming the core of LLMs and the transformer architecture, have shown significant advancements in various NLP tasks (Cao et al., 2024; Li et al., 2023).
+I then proceed with Neural Network models, which form the core of LLMs and the transformer architecture.
 In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain.
@@
 First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns.
-Second, the superior performance of deep learning architectures emphasizes the importance of capturing contextual information and word interdependencies for textual analyses.
+Second, the superior performance of deep learning architectures emphasizes the importance of contextual information and word interdependencies in textual analyses.
 3.3 Information Value of Analyst Reports
-How can the statistical measure $R^{2}_{\mathrm{OOS}}$ be translated into economic terms?
+How to translate statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms?
 In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3.
-This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal while considering price impact.
+This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal, while accounting for price impact.
 In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
@@
 This period corresponds with the global COVID-19 pandemic, which likely introduced additional uncertainty and volatility in financial markets, impacting the value of information from analyst reports.
-These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions and external events.
+These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions.
 The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
@@
 Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time.
-These alternative estimates support my main findings, reinforcing the substantial value of analyst information and the robustness of the approach.
 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
-It is likely that the information value of sell-side analysts will vary in the cross-section of stocks.
+The information value of sell-side analysts likely vary in the cross-section of stocks.
 Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984).
@@
 Stocks with larger market equity have significantly higher information value from analysts.
-Economically, column (4) means a 1% increase in market equity is associated with a 0.864% increase in analyst information value.
+Economically, results in column (4) indicate a 1% increase in market equity is associated with a 0.864% increase in analyst information value.
 Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value.
-I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact for large stocks.
-This decomposition can be expressed logarithmically as:
+I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact of large stocks.
+The decomposition can be expressed logarithmically as:
 $\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
-Panel B of Table 9 demonstrates that the main reason behind the positive size elasticity of information value is the higher liquidity of large stocks.
-Overall, analyst information does not appear to explain the higher contemporaneous return volatility of large stocks, yet it remains more valuable due to the higher liquidity of these stocks.
+Panel B of Table 9 shows that the higher information value for large stocks stems primarily from their greater liquidity.
+Analyst reports of large stocks do not explain higher contemporaneous return volatility, yet their value increases with stock size due to enhanced liquidity.
 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
-I next examine how the information value of analysts differs with herding behavior.
-Prior research has stated that analysts making bold forecasts incorporate their private information more completely and provide more relevant information to investors than those making herding forecasts (Clement and Tse, 2005).
+Analysts who make bold forecasts incorporate their private information more completely and provide more relevant information to investors than those who herd in their forecasts (Clement and Tse, 2005).
+I examine how information value varies with analysts’ herding behavior.
 To investigate the cross-section of analyst characteristics, I regress the stock-analyst-day information value on an indicator of forecast boldness, controlling for firm-fixed effects, analyst-fixed effects, and year-fixed effects.
@@
 On average, information from bold reports is 29% more valuable than that from herding reports.
-The magnitude remains largely unchanged when testing with information value derived from report text.
-However, the insignificant coefficient estimates for $Brokersize$ suggest that the value of analyst information is not necessarily tied to the size of the brokerage firm.
-The subsequent research question is whether the association between boldness and information value is driven by higher information content or lower price impact.
+This magnitude persists when examining the information value of report text alone.
+The insignificant coefficients for $Brokersize$ suggest that the value of analyst information is not related to brokerage firm size.
+A subsequent research question is whether the relationship between analyst boldness and information value is driven by higher report informativeness or lower price impact.
 In Panel B, I regress the two decompositions of information value on the boldness indicator.
@@
 In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant.
-This finding suggests that while bold forecasts contain more valuable information, they do not necessarily lead to a higher price impact compared to herding forecasts.
 3.3.3 The Value of Analyst Information around Earnings Announcements
 Previous evidence underlines the information content of analyst reports in the first week following earnings announcements.
-To expand on this, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
+Accordingly, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
 Figure 5 illustrates the information value of analyst reports between earnings announcements on a weekly basis, from the first week after the earnings announcement up to the $13^{th}$ week.
@@
 (2024).
-The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect information value.
-The model includes stock and year-fixed effects to control for time-invariant stock characteristics and market-wide yearly trends.
+The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect analyst information value.
+Stock and year-fixed effects are included to control for time-invariant stock characteristics and market-wide trends.
 Table 11 presents regression results showing the influence of earnings announcements and trading volume on information value.
@@
 Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable.
-The interaction term ($Week\times TradingVolume$) is significantly positive in columns (1)-(4), indicating that explained return volatility rises with higher trading volumes during post-earnings periods.
 Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact.
-These findings suggest that analyst reports provide more valuable insights in the immediate post-earnings announcement period, particularly when investor uncertainty is high.
-This increased value likely reflects analysts’ ability to interpret complex earnings information, helping investors make sense of the latest financial data during these critical periods.
+These findings indicate that analyst reports provide more valuable insights in the first week after earnings announcements, especially during periods of high investor uncertainty.
+Investors value analysts’ ability to promptly interpret complex earnings information and make sense of the latest earnings numbers.
 4 Conclusion
@@
 This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
+Figure A5: Shapley Values for Topic Importance across Industries
+This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value.
+For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings.
+The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
 Figure A6: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
@@
 The shaded area marks the pandemic recession in 2020.
-Figure A6: Shapley Values for Topic Importance across Industries
-This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value.
-For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings.
-The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
 Figure A7: Shapley Values for Topic Importance by Analyst Characteristics

### v3

@@
 1 Introduction
-Sell-side analysts play a crucial role in financial markets by processing information for investors.
+Sell-side analysts play a crucial role in financial markets by producing and processing information for investors.
 Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021).
@@
 Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018).
-However, analyst reports are meant to provide contextual information beyond simple sentiment dimension and the context has been perceived at least as important as the quantifiable outputs.
+However, analyst reports are meant to provide contextual information beyond the simple sentiment dimension.
+The context has been perceived at least as important as the quantifiable outputs.
 According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations.
@@
 The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024).
-For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the “earnings analysis” topic, the latter provides additional contextual information and reasoning that can be interpreted and analyzed by LLMs.
-I quantify the value of analyst information in analyst reports through three general steps: analyst output representation, econometric modeling, and value decomposition.
+Consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning.
+The question now becomes: do investors disregard obvious facts and figures in favor of details in lines?
+I investigate the question through three general steps: analyst output representation, econometric modeling, and value decomposition.
 Analyst outputs include both quantitative information and qualitative information.
-The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices.
-For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs.
-Each report is mapped into a structured embedding space using the LLMs’ transformer architecture.
-With structured text representation, I evaluate how well analyst information explains stock returns using an R-squared measure, which also serves as the terminology of information content.
-A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings.
-To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients.
-This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns.
-Additionally, I evaluate the out-of-sample explanatory power of quantitative and qualitative measures for abnormal stock returns to assess model stability.
-I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information.
-When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually.
-These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content within analyst report text.
-The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations.
-Both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions.
-Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%.
-In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
-A notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al.
-(2014), Kim and Song (2015), Lobo et al.
-(2017), and Barron et al.
-(2017), An investigation on samples within and beyond the earnings announcement windows reveals that the information content of analyst reports peaks in the first week following earnings announcements.
-The finding supports Huang et al.
-(2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients.
-Using separately trained models on samples within and beyond earnings announcement windows, the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods.
-The significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
-Analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models to isolate the unique information from analyst reports.
-The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights.
+The quantitative data consists of numerical forecasts like stock recommendations, earnings forecasts, and target prices.
+For qualitative textual information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
+With structured representation, I evaluate how well analyst information explains stock returns as evidence of information content.
+A significant challenge in projecting stock returns onto text embeddings is the risk of in-sample overfitting due to high dimensionality.
+To address this issue, I employ Ridge regressions.
+The simple model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market.
+Additionally, I use out-of-sample R-squared to evaluate model performances.
+I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR_{[-1,+1]}$), surpassing the 9.01% explained by numerical information.
+When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using either information type alone.
+These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
+The informativeness of analyst reports varies substantially with forecast changes.
+Analyst reports accompanied by forecast revisions provide significantly more valuable information to financial markets compared to reiterations.
+For reports containing recommendation revisions, the combination of numerical and textual information generates an out-of-sample $R^{2}$ of 22.63%.
+In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$ is not statistically significant, indicating they provide minimal new information to market participants.
+Previous research emphasizes analysts’ role in information processing around corporate earnings announcements( e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017).
+My analysis on samples during and outside earnings announcement periods reveals that analyst reports’ information content peaks in the first week following earnings releases.
+The finding aligns with Huang et al.
+(2018), who argue that prompt responses to corporate disclosures provide valuable insights to clients.
+Using distinct models trained on samples during and outside earnings announcement windows, the out-of-sample $R^{2}$ approximately doubles in the first week following earnings announcement dates compared to other periods.
+The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings information for investors.
+To address the concern that my results merely capture analysts’ “piggybacking” on earnings conference call information or post-earning-announcement drift (PEAD), I control for both earnings surprises and the latest earnings conference call transcripts using text embeddings.
+The results show that analyst reports provide distinct and valuable insights even after accounting for these alternative explanations.
 As raised in the review paper of Bradshaw et al.
-(2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question.
+(2017), identifying the qualitative attributes of analyst reports that investors find most salient remains largely unsolved.
+I develop a systematic approach to quantify the importance of different report content.
+By exploiting the additive property of text embeddings and implementing a Shapley value decomposition, I attribute the explanatory power to 17 major topics discussed in analyst reports.
 To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports.
-The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers.
-This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
-So far, I have measured information value statistically with R-squared.
-To translate it into economic terms, I adopt Kadan and Manela (forthcoming)’s framework to quantify the dollar value of analyst information to strategic investors.
+The results reveal that Income Statement Analyses, particularly interpretations of realized earnings, account for over half of the reports’ explanatory power in terms of out-of-sample $R^{2}$.
+These results reinforce the critical role of analysts in interpreting financial data and constructing the narratives and context essential for investment decisions.
+To translate statistical measures of information content into economic terms, I adopt Kadan and Manela (2024)’s framework to quantify the dollar value of analyst information for strategic investors.
+The multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination.
 The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts.
-This multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination.
-Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
-The findings reveal significant value in analyst information.
-For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports are \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million from both combined.
-The annualized aggregate information value is estimated at \$6.89 million, based on an average of 15 days with reports released per year.
-According to Kadan and Manela (forthcoming), this estimate represents a conservative lower bound.
-Notably, these values increase for large stocks, bold analysts, and reports released promptly following earnings announcements.
+Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to price impact, representing the cost of trading on this information.
+The analysis reveals substantial economic value in analyst information.
+For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports amount to \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when combined.
+Based on an average of 15 report days annually, the aggregate annualized information value reaches \$6.89 million—a conservative lower bound according to Kadan and Manela (2024).
+This value increases for large-cap stocks, bold analysts, and reports released promptly following earnings announcements.
 My paper relates to three strands of literature.
-This paper first relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024).
-Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights.
-My work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations.
-A direct comparison with the Naive Bayes tone classification used in Huang et al.
-(2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone.
-Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure from 3% to above 10%.
-The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information.
-Topic decomposition analysis offers a new dimension for understanding analyst information value by revealing their role in interpreting financial data.
+It first connects to recent studies of market reactions to security analyst reports (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024).
+Earlier investigations primarily rely on manual text coding and sentiment analysis, which are labor-intensive and may overlook nuanced insights.
+My work demonstrates the superiority of state-of-the-art LLMs in generating contextualized textual representations.
+Compared to Huang et al.
+(2014)’s Naive Bayes tone classification, LLMs offer two main advantages: better capture of overall report sentiment and extraction of richer information beyond tone.
+Moving from simple tone analysis to comprehensive LLM embeddings boosts the $R^{2}$ from 3% to over 10%.
+The substantial improvement in explanatory power highlights LLMs’ potential to transform textual analysis.
+Additionally, the topic decomposition approach provides novel insights into analysts’ value by revealing their role in interpreting financial data.
 Second, this work contributes to the growing field of machine learning and artificial intelligence in finance.
@@
 (2024), and Beckmann et al.
-(2024) have showcased the potential of implementing LLMs in financial contexts.
+(2024) have showcased the potential of implementing LLMs in a financial paradigm.
 However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns.
-This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI.
-Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010).
-Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information.
+By exploiting the additive property of LLM embeddings and developing a Shapley value decomposition approach, my research provides a new framework for explainable AI.
+Third, this paper advances the understanding of analyst tipping - the practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010).
+Tipping allows brokerages to offset research costs by earning commissions from clients who benefit from the privileged information.
 Prior studies, such as those by Green (2006) and Kadan et al.
 (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%.
-This paper extends prior work by quantifying the economic dollar value of such insider information.
+This paper extends by quantifying the economic dollar value of such privileged information.
 The rest of the paper is organized as follows.
-Section 2 introduces methodology, measures of analyst information value, and interpretation.
-Section 3 summarizes data construction and empirical results.
+Section 2 introduces methodology, measures of analyst information value, and interpretation framework.
+Section 3 describes the data and empirical results.
 Section 4 concludes.
 2 Methodology
-In this section, I introduce how I employ large language models for topic modeling and text embeddings, the information content estimation, the strategic information value estimation, and the decomposition metric used to evaluate the information value of topics.
-[^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
+This section describes the key methodologies: the use of large language models for topic modeling and text embeddings, estimation of information content and strategic information value, and decomposition metrics for evaluating topic-specific information value.
+[^1] The details of machine learning models used to explain contemporaneous stock returns are provided in the Online Appendix C.
 2.1 Large Language Models
-I harness the power of advanced generative AI models to conduct comprehensive analyses of analyst reports.
-The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors.
+The analysis of analyst reports leverages three advanced generative AI models.
+A fine-tuned BERT model classifies sentences according to predefined criteria, ChatGPT extracts and synthesizes major topics through prompting, and LLaMA generates high-dimensional vector embeddings to represent report content.
 2.1.1 ChatGPT Prompting
@@
 Here is the report content: {text}.
-It ends up with 13 categories defined as follows:
-  • Executive Summary: Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
-  • Company Overview: Offers a comprehensive description of the company, including its history, business model, and key products or services; discusses the company’s organizational structure, management team, and corporate governance; analyzes the company’s mission, vision, and strategic objectives.
-  • Industry Analysis: Provides an in-depth analysis of the industry in which the company operates; includes information on market size, growth trends, and key drivers; discusses the regulatory environment, technological advancements, and other external factors affecting the industry; analyzes the industry’s competitive dynamics and the company’s position within the industry.
-  • Competitive Landscape: Identifies the company’s main competitors and their market share; compares the company’s products, services, and pricing strategies with those of its competitors; analyzes the strengths and weaknesses of the company and its competitors; discusses potential new entrants and substitutes that could disrupt the competitive landscape.
-  • Financial Analysis: Analyzes the company’s revenue, expenses, profitability, assets, liabilities, shareholders’ equity, liquidity, solvency, capital structure, cash flows, and key financial ratios; compares the company’s financial performance with that of its competitors and industry benchmarks; discusses trends in the company’s financial performance over time.
-  • Business Segments: Provides a detailed analysis of the company’s various business segments or divisions; discusses the financial performance, growth prospects, and challenges of each segment; analyzes the contribution of each segment to the company’s overall revenue and profitability.
-  • Growth Strategies: Discusses the company’s strategies for driving future growth, such as organic growth initiatives, product innovations, and geographic expansions; analyzes the company’s mergers and acquisitions (M&A) strategy and potential targets; examines the company’s investments in research and development (R&D) and marketing.
-  • Risk Factors: Identifies and analyzes the key risks facing the company, such as market risks, operational risks, financial risks, and legal/regulatory risks; discusses the potential impact of these risks on the company’s financial performance and growth prospects; examines the company’s risk management strategies and mitigation measures.
-  • Management and Governance: Provides an overview of the company’s management team, including their experience, expertise, and track record; analyzes the company’s corporate governance practices, such as board composition, executive compensation, and shareholder rights; discusses the company’s succession planning and key person risks.
-  • Environmental, Social, and Governance (ESG) Factors: Analyzes the company’s performance and initiatives related to environmental sustainability, social responsibility, and corporate governance; discusses the potential impact of ESG factors on the company’s reputation, risk profile, and financial performance; examines the company’s compliance with ESG regulations and industry standards.
-  • Valuation: Estimates the intrinsic value of the company’s shares using various valuation methodologies, such as discounted cash flow (DCF) analysis, relative valuation multiples, and sum-of-the-parts analysis; compares the company’s valuation with that of its peers and historical benchmarks; discusses the key assumptions and sensitivities underlying the valuation analyses.
-  • Investment Thesis: Summarizes the key reasons for investing (or not investing) in the company’s shares; discusses the potential catalysts and risks that could impact the company’s valuation and stock price performance; provides a target price or price range for the company’s shares based on the valuation analyses and investment thesis.
-  • Appendices and Disclosures: Includes additional supporting materials, such as financial statements, ratio calculations, and detailed segment data; provides important disclosures, such as the analyst’s rating system, potential conflicts of interest, and disclaimers; discusses the limitations and uncertainties of the analysis and the need for further due diligence by investors.
-The Financial Analyses category is further divided into four subcategories: Income Statement Analyses, Balance Sheet Analyses, Cash Flow Analyses, and Financial Ratios.
-Additionally, a category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the predefined topics.
-In total, this process generates 17 distinct topics for classification.
-The distribution of topics is relatively stable across years, as illustrated in Figure 1.
+It ends up with 16 categories defined in Table LABEL:tab:_topic.
+A category labeled “None of the Above” is included to accommodate sentences that do not fit into predefined topics.
 To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category.
@@
 Here is the sentence from the analyst report: {sentence}.
-This training sample serves as the training sample for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus.
+This labeled sample is used to fine-tune a BERT model, which then classifies all sentences in the corpus.
 The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample.
+The distribution of topics is relatively stable across years, as illustrated in Figure 1.
 Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively.
 These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics.
-Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
+Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least frequent topics.
 Several factors contribute to these observed frequencies.
 The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase.
-The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level.
-Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in ESG discussions.
+Executive Summary sections appear infrequently because they are difficult to identify at the sentence level.
+While ESG discussions are minimal overall, they increase notably after 2020, reflecting growing market interest in sustainability issues.
 2.1.2 Text Representations
-A systematic textual analysis requires a structured representation of analyst reports.
+I need a structured representation of writtrn reports.
 Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps.
@@
 Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters.
-Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary.
-This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand.
+These tokens are then mapped to unique identifiers based on a predefined vocabulary learned during model training.
 To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”.
@@
 At the heart of LLMs lies the transformer architecture.
-The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance.
-A core output of the transformer architecture is text embeddings.
-These embeddings are dense, continuous vectors that capture semantic and syntactic information.
+A core output of the transformer is text embeddings - dense, continuous vectors that capture semantic and syntactic information.
 Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia.
 These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
-To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI.
+For this analysis, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI.
 The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.
-The tokenizer processes the text data, converting it into tokens that the model can understand.
-By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports.
-These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and economic modeling.
+The resulting contextual embeddings capture nuanced report information and are suitable for downstream tasks like sentiment analysis and economic modeling.
 2.2 Information Content
 Following Asquith et al.
-(2005), I use $R^{2}$ as a measure for information content.
-To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
+(2005), I consider $R^{2}$ as a proxy for information content.
+To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations annually using an expanding window:
 $CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (1)
 […]
-where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ is the structured representation of information from analyst report $j$ on stock $i$ released at day $t$.
-I use the $60\%$ of data as the initial training sample.
-Therefore, the out-of-sampel test starts in 2015 and ends in 2023.
-Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ as
+where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ for stock $i$ at day $t$.
+The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
+Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ using:
 $R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum_{(i,t)\in\mathcal{T}}CAR_{[-1,+1],it}^{2}},$ (3)
 where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation.
-$R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
-To evaluate the statistical significance of the $R_{\mathrm{oos}}^{2}$, I employ a modified version of the Diebold and Mariano (2002) test for comparing predictive accuracy between two models.
-Given the strong cross-sectional dependence in the stock-level analysis, I follow Gu et al.
-(2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional average of prediction errors from each model rather than individual returns.
+$R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the proportion of return variance attributable to analyst information.
+To assess the statistical significance of the $R_{\mathrm{oos}}^{2}$, I implement a modified version of the Diebold and Mariano (2002) test for predictive accuracy comparison.
+Given the strong cross-sectional dependence inherent in the stock-level analysis, I follow Gu et al.
+(2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional averages of prediction errors.
 2.3 Information Value
-I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal:
+Following Kadan and Manela (2024), I evaluate the value of information using a measure that approximates information value in a strategic setting where multiple informed investors possess imperfect signals:
 $\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
-where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price of day t-2.
-$\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return.
-$N$ denotes the number of analyst reports about stock $i$ on day $t$.
-To estimate the price impact $\widehat{\lambda}_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$).
-Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals.
-The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991).
+where $\widehat{\Omega}_{it}$ represents the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price on day t-2.
+$\widehat{r}_{ijt}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$.
+Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return.
+$N$ represents the number of analyst reports for stock $i$ on day $t$.
+To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$).
+The order flow is derived from changes in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals, with trades classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991).
 [^2] The algorithms from Ellis et al.
@@
 (2007) are also used in robustness tests.
-The regression model is
+The regression specification is:
 $r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
@@
 In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$.
-The 3-day interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$.
-Thus, each regression has $1170(=390*3)$ observations.
-Essentially, the measure reflects the ratio of return volatility to price impact.
-It is derived based on the theoretical framework of Back et al.
-(2000), which is a continuous-time version of Kyle (1985) model.
-Due to risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of the asset price.
-The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
+The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
+The measure reflects the ratio of return volatility to price impact and is grounded in the theoretical framework of Back et al.
+(2000), which is a continuous-time extension of Kyle (1985) model.
+Under risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of asset price.
+The intuition is illustrated with an extended single-auction Kyle model with incomplete information in Online Appendix D.
 I employ the measure for two key reasons.
@@
 Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
-[^3] Kadan and Manela (forthcoming) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply to the analyst case.
+[^3] Kadan and Manela (2024) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply in the analyst context.
 Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced.
-To tackle this issue, I apply the delta method to estimate the mean and variance of information value over sample $s$.
-The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
-2.4 Shapley Value Decomposation
-To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology.
+To address this challenge, I apply the delta method to estimate the mean and variance of information value over sample $s$.
+The derivation of the first-order estimator using the delta method is detailed in Online Appendix E.
+2.4 Shapley Value Decomposition
+To quantify the relative contribution of different topics, I implement the Shapley Additive exPlanations (SHAP) methodology.
 SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves.
-It was later applied for machine learning model interpretation by Lundberg and Lee (2017) and Chen et al.
-(2022).
-In my context, the approach allows me to attribute the overall explanatory power of an analyst report to its constituent topics.
-Leveraging the additive nature of text embeddings, I design a novel decomposition method to systematically assess the relative importance of various topics in analyst reports.
-LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
+It was later applied for machine learning model interpretation (Lundberg and Lee, 2017 and Chen et al., 2022).
+The approach allows for attributing an analyst report’s overall explanatory power to its constituent topics.
+Leveraging the additive nature of text embeddings, I develop a novel decomposition method to systematically evaluate the relative importance of various topics in analyst reports.
+The report-level embeddings are generated as the average of embeddings across all tokens and layers. The report embeddings can be formalized as:
 $y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
-where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ is the total number of words in the report.
-If the tokens can be classified into different topics, the embedding can be fully decomposed into different topic-specific embeddings:
+where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ denotes the total token count of the report.
+Given the tokens can be classified into distinct topics, the embeddings can be fully decomposed into topic-specific components:
 $y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
-where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
-One concern raised by the nature of transformer models is that each token in the input sequence can pay attention to other tokens through a mechanism called “self-attention” or “scaled dot-product attention.” This architecture allows the model to weigh the importance of each token relative to every other token and capture contextual information.
-For example, the embedding of token $i$ contains information from token $j$, which may pertain to a different topic.
-To isolate the self-attention across topics, I process the embeddings sentence by sentence to ensure the token embeddings only reflect the context information within the sentence.
-I then take the token-weighted mean of sentence embeddings as the report representation, which is referred to as sentence-segment embeddings.
+where $P$ denotes the total number of topics, $y_{p}^{emb}$ represents the embedding for topic $p$, $N_{p}$ is the number of tokens classified as topic $p$, and $i_{p}$ indicates the index of tokens within topic $p$.
+A consideration in the transformer architectures is the self-attention mechanism, which allows the model to compute relative importance weights between tokens.
+This architecture enables each token in the input sequence to “talk” with each other.
+Specifically, the embedding of token $i$ incorporates contextual information from token $j$, even when these tokens belong to distinct topics.
+To mitigate this cross-topic attention effect, I process embeddings sentence by sentence to ensure the token embeddings only reflect intra-sentence contextual information.
+I then take the token-weighted mean of sentence embeddings as the report representation, which I designate as sentence-segment embeddings:
 $y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
-where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
-The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context across sentences.
-Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Specifically, the Shapley value can be written as:
+where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embeddings of the $i$-th sentence.
+The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding quantifies the information content of context across sentences.
+Subsequently, I compute the Shapley value across topics using sentence-segmented embeddings. The Shapley value is defined as:
 $\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
-where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set.
-Note that $\binom{n}{k}$ is the binomial coefficient.
-$y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$.
-For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly.
-[^4] Shapley value requires an input for $X$ when it is missing.
+where the sum extends over all subsets $S\subseteq P\backslash\{p\}$, including the empty set.
+$y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$.
+The $R_{\text{oos}}^{2}$ is calculated for each unique combination of topic embeddings.
+[^4] Shapley value requires a defined input for missing $X$ values.
 I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation.
-This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions.
+This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ across topics, revealing their relative contributions to explaining market reactions.
 3 Data and Empirical Results
-This section presents the empirical analysis of analyst report information.
-I begin by examining the information content and decomposing it across different topics.
+In this section, I present the empirical analysis of analyst report information, beginning with an examination of information content and its decomposition across topics.
 I then quantify the strategic dollar value of analyst information.
@@
 3.1 Data
-Sell-side analyst reports are obtained from the Mergent Investext Database.
-I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023.
-The initial reports are in PDF format, and I use Adobe Acrobat to extract plain text from each report.
-The text is chunked into sentences.
-To get clean content of firm-relevant discussions, I fine-tune a BERT model to remove the boilerplate segments.
-[^5] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions.
+I refer to the Mergent Investext database for 223,091 sell-side analyst reports for S&P 100 firms spanning the years 2000 to 2023.
+The reports, initially in PDF format, are converted to plain text using Adobe Acrobat.
+The extracted text is then segmented into sentences.
+To ensure a clean dataset focused on firm-relevant discussions, I fine-tune a BERT model to filter out boilerplate content.
+[^5] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions.
 Following Li et al.
-(2024), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports.
-The BERT model fine-tuned for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
-The Mergent Investext Database provides information on analyst names and the release date of each report.
-To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al.
+(2024), I construct a training dataset by randomly sampling one report from each of the top 20 brokerage firms by report volume and manually label each sentence.
+The BERT model fine-tuned for boilerplate classification achieves a 97.31% accuracy rate on test data.
+The Mergent Investext Database provides analyst identification and report release dates.
+To connect Investext with the I/B/E/S database, I first match lead analyst names to their corresponding Analyst IDs (AMASKCD) following Li et al.
 (2024).
-I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database.
-Since multiple reports could be released between a forecast announcement and its revision, I follow Huang et al.
-(2014) and define the matching window for each report as 2 days before the announcement date to 2 days after the revision date.
-[^6] In I/B/E/S, each analyst forecast for a specific firm has two time stamps.
-The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one.
-The revision date (REVDATS) is the date the forecast becomes invalid or is revised.
+I then match individual reports to corresponding EPS forecasts, price targets, and recommendation files in I/B/E/S.
+Given that multiple reports may be issued between a forecast announcement and its subsequent revision, I adopt Huang et al.
+(2014)’s approach and define the matching window as spanning from two days before the announcement date to two days after the revision date.
+[^6] I/B/E/S records two timestamps for each analyst forecast: the announcement date (ANNDATS) when the new forecast is issued and the old one is revised, and the revision date (REVDATS) when the forecast becomes invalid or undergoes revision.
 Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window.
-I also compile firm characteristics from the CRSP and Compustat databases.
-[^7] See Table A2 and Table A3 for detailed variable definitions and summary statistics.
-After the matching procedure and excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 stocks from 1,305 analysts across 140 unique brokerage firms.
-Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts across years and the Fama-French 12 industries.
-It also summarizes the average number of pages and tokens per report, which are important indicators of report lengths.
-Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al.
-(2023), and the average number of pages per report is consistent with that in Huang et al.
+Additional firm characteristics are sourced from CRSP and Compustat.
+[^7] Detailed variable definitions and summary statistics are provided in Table A2 and Table A3, respectively.
+After the matching procedure and excluding reports released on earnings announcement days, the final sample consists of 122,252 reports covering S&P 100 stocks, produced by 1,305 analysts across 140 distinct brokerage firms.
+Table 2 provides summary statistics of analyst reports, including the number of reports, unique brokerage firms, and analysts across years and the Fama-French 12 industries.
+It also reports average page and token counts per report as measures of report length.
+The temporal distribution of analyst reports aligns with patterns documented in Bonini et al.
+(2023), while the average page count per report corresponds with findings in Huang et al.
 (2014).
-[^8] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline.
-This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
-The stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023.
-To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
-Given my goal to estimate the information value of analyst outputs, I need to control for other information available to investors.
-In Section 3.2.2, I compare textual information from earnings conference calls and subsequent analyst reports.
-Earnings conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Available at https://seekingalpha.com/earnings/earnings-call-transcripts..
-These transcripts provide comprehensive record of the calls, including executive presentations and manager-analyst interactions like Q&A sessions.
-I sample earnings conference call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat.
-The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks.
-This process yielded a final sample of 7,909 firm-quarter-level conference calls.
+[^8] The data reveals a peak in the number of analyst reports in 2013, followed by a gradual decline.
+This trend is likely influenced by the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and MiFID II in Europe (Bradshaw et al., 2017).
+Stock-level price impacts are estimated using intraday data from the NYSE TAQ database for an out-of-sample period spanning January 2, 2015, to December 28, 2023.
+To ensure temporal comparability, all prices are adjusted for inflation using the Consumer Price Index (CPI), normalized to December 2020 as the base period.
+To isolate the information value of analyst outputs, I control for alternative information sources available to investors.
+In Section 3.2.2, I conduct a comparative analysis between earnings conference call content and subsequent analyst reports.
+Conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts..
+These transcripts provide detailed documentation of executive presentations and manager-analyst interactions during Q&A sessions.
+I sample transcripts of S&P 100 constituent firms from 2005 to 2023, matching them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat.
+The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks, yielding a final sample of 7,909 firm-quarter-level conference calls.
 The set of numerical variables used as input in machine learning models includes:
@@
 $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
-$TP\_REV$ indicates target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
-$Boldness$ is an indicator variable that takes the value of 1 if revisions are either above both the consensus and previous forecasts or below both.
-$SR$ represents stock recommendation, following the I/B/E/S rating system, where a rating of 5 denotes a strong sell and a rating of 1 indicates a strong buy.
-$ERet$ represents the 12-month return forecast, calculated by scaling the 12-month price target by the stock price 50 days before the report release.
-$Prior\_CAR$ represents the cumulative abnormal return over time period [-10, -2].
-Firm-Level Measures: $Size$ is measured by market equity.
-The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value.
+$TP\_REV$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 days pre-release.
+$Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both.
+$SR$ denotes stock recommendations following the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell).
+$ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 days pre-release.
+$Prior\_CAR$ measures cumulative abnormal return over [-10, -2].
+Firm-Level Measures: $Size$ is quantified by market equity.
+Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value.
 Following Huang et al.
-(2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value.
-The $Miss$ variable is a binary indicator set to one if $SUE$ is positive.
-$TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding.
-The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI).
+(2018), Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value.
+$Miss$ is a binary indicator for positive $SUE$.
+$TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding.
+Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI).
 $Fluidity$ measures firms’ product market fluidity.
-Industry-Level Measures: $IndustryRecession$ is an indicator variable equal to one if the firm’s industry (defined by Fama-French 48 classification) experiences a negative monthly return in the bottom quintile.
-Macroeconomic Measures: The $TimeTrend$ variable captures the temporal dimension by recording the number of years elapsed since the beginning of the sample period.
+Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the firm’s industry (defined by Fama-French 48 classification) experiences negative monthly returns in the bottom quintile.
+Macroeconomic Measures: $TimeTrend$ is a temporal variable measuring the number of years since the start of the sample period.
 3.2 Information Content of Analyst Reports
-In this section, I examine the information content of analysts’ quantitative and qualitative outputs by projecting contemporaneous cumulative abnormal returns on analyst information.
-Furthermore, I attribute the information content of analyst report text across topics using the Shapley value decomposition introduced in Section 2.4.
-3.2.1 Information Content of Quantitative and Qualitative Information
-Table 2 compares the information content embedded in numerical and textual measures, quantified with $R^{2}_{\mathrm{OOS}}$.
-The estimation model applied in this table is the Ridge regression specified in Equations 1 and 2.
-The objective variable is ${CAR}_{[-1,+1]}$ centered at the analyst report release date.
+This section investigates market reactions to analysts’ quantitative and qualitative outputs.
+I employ the Shapley value decomposition methodology introduced in Section 2.4 to attribute the information content across distinct topics.
+3.2.1 Quantitative and Qualitative Information
+Table 3 compares the information content embedded in numerical and textual measures.
+The analysis employs Ridge regression as specified in Equations 1 and 2, with ${CAR}_{[-1,+1]}$ centered at the report release date serving as the dependent variable.
 For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database.
-I estimate separate Ridge regression models for distinct input configurations and employ an expanding window approach, constructing training sets from 2000 to the year preceding each target year.
-Panel A of the table delineates the performance of four types of input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
-Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% with forecast revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions.
-The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features.
-I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4).
-The models with revision and numerical inputs significantly outperform a zero-estimation benchmark.
-Columns (5) and (6) elucidate the performance of Ridge models with text embeddings as input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20, underscoring the significant information content of the report text in explaining contemporaneous stock returns.
-The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession.
-The DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, suggesting substantial information content in analyst report text compared with forecast revisions.
-Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs.
-Language model outputs may incorporate future information that should not be available.
-Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period.
-In columns (7) and (8), I further investigate the model performance when combining both forecast revisions and text embeddings in the Ridge regression.
-The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model.
-I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5).
-The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$.
-The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by investors.
-The explanatory power might be derived from quantitative information in the numbers mentioned in report text.
-I conduct robustness checks using content with numbers removed.
-To address the concern, I exclude numerical data from the written reports before generating text embeddings and report model performance in Table A4.
-The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, significantly higher than that of quantitative revision measures (t-statistic = 2.40).
-The robustness check suggests that the superior performance of text-based models is not driven by the numbers embedded in text.
-LLMs exhibit improved language skills as the parameter size increases.
-I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI text-embedding-3-small model, and LLaMA-3 model.
-The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models.
-Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 model.
+The estimation utilizes distinct input configurations with an expanding window approach, where training sets span from 2000 to the year preceding each target year.
+Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
+Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for forecast revision input, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst revisions.
+The incorporation of additional 14 numerical features in column (2) yields a modest improvement to 9.08% $R^{2}_{\mathrm{OOS}}$.
+Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the Diebold-Mariano test t-statistics reported in columns (2) and (4).
+Columns (5) and (6) reveal that Ridge models with text embeddings achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text.
+The $R^{2}_{\mathrm{OOS}}$ remains stable from 2015 to 2023, except during the 2020 pandemic recession.
+A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating substantial information content of report text relative to forecast revisions.
+Addressing the look-ahead bias concern raised by Sarkar and Vafa (2024) regarding pre-trained LLMs, I examine the $R^{2}_{\mathrm{OOS}}$ in 2023.
+Since the employed LLMs were trained on pre-2023 data, this period represents true out-of-sample performance, validating the robustness of the main findings.
+Columns (7) and (8) examine the model performance when integrating both forecast revisions and text embeddings.
+This combined approach achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively.
+The DM t-statistics of 3.95 and 3.77 for comparisons between combined input versus revision input and text input demonstrate that report text and forecast revisions contain distinct, complementary information sets valued by investors.
+The explanatory power might stem from quantitative information embedded in the numbers mentioned within the text.
+To address this concern, I conduct robustness checks using content with numbers removed.
+Table A4 presents the performance of models using number-free text embeddings.
+The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% significantly exceeds that of quantitative revision measures (t-statistic = 2.40), suggesting that the superior performance of text-based models is not driven by embedded numerical content.
+I examine the scaling law relationship by comparing LLMs with varying parameters, including BERT, OpenAI text-embedding-3-small, and LLaMA-3 models.
+The LLaMA models feature significantly more parameters than the BERT and OpenAI models.
+Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 benchmark.
 The input in Panel B is report text embeddings, comparable to column (3) in Panel A.
-A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark.
-Overall, the BERT and OpenAI models underperform the LLaMA-2 model.
-The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model.
-The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs.
-Pairwise comparisons between LLMs also shed light on how the scale of the model influences performance capabilities.
-To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al.
+A positive t-statistic indicates superior performance relative to the LLaMA-2 model.
+While smaller models like BERT and OpenAI underperform LLaMA-2, the larger LLaMA-3 model shows only marginally better performance with an $R^{2}_{\mathrm{OOS}}$ of 9.66%, though the improvement is statistically insignificant.
+These results demonstrate the robustness of analyst report text information content across different LLMs and illuminate the relationship between model scale and performance capabilities.
+To validate the robustness of contemporaneous measures, I examine alternative $CAR$ windows, including 2-day and 5-day intervals following Huang et al.
 (2014) and Asquith et al.
 (2005), respectively.
-Additionally, I examine the abnormal stock return on the days that reports are released.
-In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023.
-In Table A6, I present the information content of analyst report text across the Fama-French 12 industries.
-The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables.
-Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors.
-The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value.
-In Figure A3, I plot analyst information value by Fama-French 12 industries, revealing a pattern consistent with the information content results.
-To assess the economic magnitude, I perform OLS regressions specified as follows:
+Table A5 shows consistent $R^{2}_{\mathrm{OOS}}$ magnitudes for $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% during the 2015-2023 out-of-sample period.
+Table A6 presents the cross-industry analysis of analyst report text information content across Fama-French 12 classifications.
+Five industries exhibit $R^{2}_{\mathrm{OOS}}$ exceeding 10%: Shops, Other, Manufacturing, Chemicals, and Durables.
+Conversely, Non-Durables, Telecommunications, Energy, and Utilities sectors show significantly lower $R^{2}_{\mathrm{OOS}}$ and DM t-statistics.
+This pattern, visualized in Figure A3, suggests that industries with complex products and services derive greater value from detailed analyst insights, while those operating in stable or regulated environments benefit less from incremental analyst information.
+To evaluate the economic significance, I estimate OLS regressions specified as follows:
 $CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (10)
-This model incorporates the out-of-sample estimation of $CAR_{[-1,+1]}$ using report text embeddings or forecast revisions as input.
-For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014.
-The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension.
-For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
-Table 3 summarizes the summary statistics of the sample used and the regression results.
-The out-of-sample period starts in 2015.
-I constrain the samples to include only those with valid previous and current stock recommendations, earnings forecasts, and target prices.
-The final sample for estimating Equation (10) contains 28,837 observations.
-Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al.
+This model incorporates out-of-sample $CAR_{[-1,+1]}$ estimations using either report text embeddings or forecast revisions as input.
+The estimates for 2015, for example, are derived from Ridge regression models trained on 2000-2014 data.
+The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension, while $\widehat{CAR}_{rev}$ is estimated using three forecast revision measures under identical procedures.
+Table 4 presents sample summary statistics and regression results.
+The analysis sample is restricted to observations with valid current and previous stock recommendations, earnings forecasts, and target prices.
+Panel B’s Column (1) confirms that recommendation revisions, earnings forecast revisions, and target price revisions all contribute significantly to explaining contemporaneous abnormal stock return variations, consistent with Brav and Lehavy (2003) and Huang et al.
 (2014).
-Column (2) reveals that $\widehat{CAR}_{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2.
-The inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% in column (1) to 14.9% in column (3), underscoring the incremental value of text information.
-The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by forecast revisions.
-Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the 3-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
-In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics.
-The $\widehat{CAR}_{rev}$ model alone achieves an $R^{2}$ of 8.9%, while combined input improves it to 14.7%.
-Economically, a one standard deviation increase in $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively.
+Column (2) shows that $\widehat{CAR}{txt}$ alone achieves an $R^{2}$ of 10.5%.
+When both analyst revision measures and $\widehat{CAR}{txt}$ are included as regressors in column (3), the $R^{2}$ increases to 14.9%.
+In economic terms, one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness corresponds to 3-day abnormal return increases of 10, 40, 90, and 120 basis points, respectively.
+Columns (4) and (5) examine $\widehat{CAR}{rev}$, which aggregates all revision metrics.
+The $\widehat{CAR}{rev}$-only model achieves an $R^{2}$ of 8.9%, while combined inputs reach 14.7%.
 These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
-I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises.
-The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged.
-The stability of the coefficients suggests that market-valued analyst information is distinct from recent earnings news and is not simply a reaction to past returns.
-3.2.2 Information Content around Earnings Announcements
-Some studies have emphasized the relationship between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information sources reinforce each other in sequence.
-I examine the information content of sell-side analyst reports around earnings announcements, the most significant corporate events.
+Columns (6)-(7) introduce $SUE$ and $Prior_{C}AR$ controls to capture earnings surprises and recent news.
+The stability of $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ coefficients suggests that market-valued analyst information is distinct from recent earnings news and does not merely reflect reactions to past returns.
+3.2.2 Earnings Announcements
+Building on research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), I analyze the information content of sell-side analyst reports surrounding earnings announcements.
 Following Chen et al.
-(2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week.
-Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week.
+(2010), I group analyst reports into 13 weekly bins around earnings announcement dates.
+Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for weekly report samples.
 Consistent with Huang et al.
-(2018), a large portion of analyst reports are released immediately after earnings announcements.
-Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17.
-This performance drops sharply in the second week, turns negative in the third and fourth weeks.
-The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports surrounding earnings announcements.
-To account for potential temporal variation, I train Ridge models using separate samples of analyst reports released surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports.
-Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows.
-The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained beyond these periods.
-The $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen, declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week.
-This trend suggests that the market values information from analysts who react swiftly to earnings announcements.
-One potential concern is that analysts may simply restate information from earnings calls, and that market reactions are driven by the earnings announcements themselves rather than analyst reports.
-If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls.
-To address this issue, I use earnings conference call transcripts to differentiate the information content between the two text sources.
-I analyze reports released within one day after earnings announcements, when analysts are most likely to simply restate earnings information.
-In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcript embeddings as input.
-[^10] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C.
-This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study.
-I truncate the text at the length limits to maintain maximum information.
-The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16.
+(2018), analyst reports cluster predominantly in the immediate post-earnings announcement period.
+The information content only exhibits strong significance only in the first week, achieving an $R^{2}_{\mathrm{OOS}}$ of approximately 10% with a t-statistic of 10.17.
+The variation in performance across weekly bins may reflect Ridge models capturing dominant patterns in analyst reports issued around earnings announcements.
+To address potential temporal variation, I train separate Ridge models on analyst reports released during earnings announcement periods and those issued beyond these periods, allowing the models to extract distinct information content based on report timing.
+Panel B in Table 5 compares the model performance within and beyond earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows.
+Models trained on earnings announcement periods consistently outperform those trained on non-announcement periods.
+The $R^{2}_{\mathrm{OOS}}$ and t-statistics exhibit a declining pattern as windows lengthen, decreasing from 11.84% to 7.29% over the earnings announcement week, suggesting markets particularly value swift analyst responses to earnings news.
+To address concerns that analysts merely restate earnings call information and that market reactions primarily reflect earnings announcements rather than analyst insights, I conduct a comparative analysis of information content between these two sources.
+I focus on reports released one day after earnings announcements, when analysts would be most likely to simply repackage earnings information.
+Column (1) of Panel C presents the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using earnings conference call transcript embeddings as input.
+[^10] Due to the length of earnings conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for text embeddings in Panel C, which accommodates 8,192 tokens versus the 4,096-token limit of the LLaMA-2-13B model used elsewhere.
+Texts are truncated at these limits to maximize information retention.
+The $R^{2}_{\mathrm{OOS}}$ reaches 4.20% with a DM t-statistic of 5.16.
 Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods.
-The input comprised the mean of all report text embeddings on the day following the earnings announcement.
-The Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
-Incorporating both analyst report embeddings and earnings conference call embeddings in Ridge regressions resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3).
-The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports.
-These findings collectively refute the idea that analysts merely restate information from earnings announcements.
-Instead, they emphasize the unique and timely value analysts bring to the market.
-3.2.3 What Content in Analyst Reports is Valued by Market Investors?
-After showing that information in report text is significant, a key question arises: what specific content in analyst reports is the most valuable to the market?
-To systematically evaluate topic importance, I employ the Shapley value approach to attribute the text-based $R^{2}_{\mathrm{OOS}}$ to the 17 topics described in Section 2.1.1.
-I use sentence-segmented embeddings as report representation.
-The sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another.
-The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings.
-This reduction roughly quantifies the magnitude of contextual information content across sentences.
-Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations.
-These reductions are then averaged over the possible different permutations into a single importance measure.
-By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring an efficient attribution of explanatory power.
-The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports.
-The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each contributing to 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$.
-Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively.
-The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
-The robustness of these findings is further reinforced by additional analyses.
-As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses.
-Figure A5 reveals consistent topic rankings across industries.
-Figure A6 in the Online Appendix shows how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time.
-A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy.
-Despite the volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable.
-Analyst discussions of income statements consistently provide valuable information, even during market disruptions.
-The relative importance of topics could stem from analysts’ expertise in certain areas (supply effect) or from investors’ varying responses to different topics (demand effect).
-Figure A7 compares Shapley values across analysts of different experience levels and brokerage firm sizes.
-The consistency in topic importance across all four subsamples (experienced vs.
-inexperienced analysts, large vs.
-small brokerages) suggests that topic rankings are not driven by differences in analysts’ topical expertise.
-Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements.
-Drawing inspiration from Huang et al.
-(2018) and Blankespoor et al.
-(2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
-To facilitate the analysis, I categorize Income Statement sentences based on two criteria: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income.
+Using report text embeddings from the day following earnings announcements, this analysis achieves an $R^{2}_{\mathrm{OOS}}$ of 9.72% (column 2), substantially outperforming conference call transcripts alone.
+The combination of analyst report and earnings conference call embeddings yields further improvements, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% (column 3).
+A comparison t-statistic of 5.12 between transcript-only and combined models demonstrates analysts’ substantial incremental contribution.
+These findings collectively refute the notion that analysts merely restate earnings information, instead highlighting their distinct and timely value to the market.
+3.2.3 What Content in Analyst Reports is Valued?
+Having established the significance of report text information, I examine which specific content components most influence market response.
+I employ the Shapley value methodology to decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 topics detailed in Table LABEL:tab:_topic.
+The analysis uses sentence-segmented embeddings as report representations to mitigate cross-topic contextual effects.
+When switching from full-context to sentence-segmented embeddings, the $R^{2}_{\mathrm{OOS}}$ decreases by 2.25%, quantifying the magnitude of inter-sentence contextual information.
+Figure 2 illustrates the Shapley value decomposition, revealing a clear hierarchy in topic importance.
+Income Statement Analyses and Financial Ratio Analyses emerge as the most valuable components, contributing 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$, respectively.
+Investment Thesis and Valuation form a second tier of importance, with Shapley values of 1.76% and 1.51%.
+The remaining topics show minimal or negative Shapley values, indicating limited market reactions.
+The findings’ robustness is supported by several additional analyses.
+Figure A4 shows that normalizing Shapley values by sentence count and token length maintains Income Statement Analyses’ dominance.
+Figure A5 demonstrates consistency in topic rankings across different industries.
+The evolution of text-based $R^{2}_{\mathrm{OOS}}$ is depicted in Figure A6 of the Online Appendix.
+While a notable decline in $R^{2}_{\mathrm{OOS}}$ occurs between 2020 and 2021, likely reflecting COVID-19 pandemic disruptions in financial markets, the scaled Shapley value for Income Statement analysis (shown by the red line) exhibits remarkable stability.
+This persistence suggests that analyst commentary on income statements maintains its informational value even during market turbulence.
+To disentangle whether topic importance stems from analysts’ expertise in certain areas (supply effect) or investors’ response (demand effect), Figure A7 compares Shapley values across analyst experience levels and brokerage firm sizes.
+The consistent topic hierarchy across all four subsamples (experienced/inexperienced analysts and large/small brokerages) indicates that topic rankings are not driven by variations in analyst expertise.
+I proceed to examine analysts’ dual roles in income statement analysis: their discovery role in collecting and generating otherwise unavailable information and their interpretation role in integrating and processing available information into more interpretable signals (Huang et al., 2018; Blankespoor et al., 2020).
+To facilitate the analysis, I categorize Income Statement sentences based on: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income.
 This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
@@
 Sentence to classify: {report_sent}.
-Figure 3 provides a detailed breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, with Income Statement subcategories.
-Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw data.
-The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on analysis of actual results than on forward-looking projections.
-These patterns demonstrate that markets particularly value analysts’ interpretation of realized income, more than their reporting of public information or future expectations.
-This insight is reinforced by the findings in section 3.2.2, which show that analyst reports issued promptly after earnings announcements are the most informative.
-Taken together, these results suggest that the market places the highest value on analysts who can quickly digest financial information from earnings announcements and provide insightful interpretations.
+Figure 3 provides a granular breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, including Income Statement subcategories.
+Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw figures.
+The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on actual results than forward-looking projections.
+These patterns emphasizes that markets particularly value analysts’ ability to interpret realized income, beyond mere information reporting or future expectations.
+These findings align with the results from section 3.2.2, where analyst reports released immediately after earnings announcements demonstrate the highest information content.
+Taken together, these results suggest that the market places the highest value on analysts who can rapidly digest earnings announcement information and provide insightful interpretations.
 3.2.4 Comparison of Revision Reports and Reiteration Reports
-Researchers have documented a robust relationship between market prices and analysts’ forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021).
-Analysts’ forecast revisions provide a significant and timely source of information to the financial market and may constitute a more valuable signal than reports that merely reiterate previous forecasts.
-Therefore, I compare the information content of analyst report text associated with forecast revisions and forecast reiterations.
+Given the established relationship between market prices and analyst forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I examine the differential information content between revision and reiteration reports.
 Following the criteria of Huang et al.
-(2014), I consider a report as a revision report if its release date is within two days of the I/B/E/S forecast announcement date.
-For each forecast target in recommendations, target prices, and earnings forecasts, I categorize the report independently.
-Consequently, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
-Table 5 summarizes the out-of-sample performance of models with numerical and textual inputs using both reiteration and revision samples.
-Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions.
-Earnings forecast revisions are the most frequent, comprising 56.57% of the reports.
-In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively.
-When stock recommendations are revised, the explanatory power of various information sources becomes particularly pronounced.
-The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%.
-The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%.
-The combination of both quantitative and qualitative information further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%.
+(2014), I classify reports as revisions when released within two days of the I/B/E/S forecast announcement date.
+This classification is applied independently across three forecast targets—recommendations, target prices, and earnings.
+Therefore, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
+Table 6 presents out-of-sample model performance across reiteration and revision samples.
+The frequency of revisions varies considerably: earnings forecast revisions are most common (56.57% of reports), followed by target price revisions (31.10%), while recommendation revisions are relatively rare (2.29%).
+Recommendation revisions exhibit particularly strong explanatory power.
+Quantitative forecast revision measures achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99% when accompanying recommendation revisions, while report text yields 16.80%.
+The combination enhances explanatory power to 22.63%.
 This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
-Target price revisions are considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003).
-While reports with target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
-As Kothari et al.
-(2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output.
-For models with text from earnings forecast reiteration reports as input, the $R^{2}_{\mathrm{OOS}}$ is insignificant.
-When focusing on analyst earnings forecast revisions, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant.
-The combination yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
-Collectively, Table 5 shows that report content is more informative when it accompanies forecast revisions rather than reiterations.
-Additionally, analyst reports associated with recommendation revisions demonstrate the highest value.
+Target price revisions, noted for their incremental information value due to the limitations of discrete recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), show marked differences between revisions and reiterations.
+While reiteration reports demonstrate minimal information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58%), revision reports exhibit substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
+Earnings forecast revisions show a similar pattern.
+Reiteration reports yield insignificant $R^{2}_{\mathrm{OOS}}$, while revision reports demonstrate significant explanatory power.
+The combination of numerical and textual inputs for revision reports achieves an $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-statistic = 11.42), highlighting revisions’ superior value over reiterations.
+Overall, Table 6 shows that report content carries greater information content when accompanying forecast revisions rather than reiterations, with recommendation revision reports exhibiting the highest value.
 3.2.5 Comparison with Analyst Report Sentiment
-Huang et al.
-(2014) emphasize the information content of analyst report text by extracting tone using a Naive Bayes machine learning approach.
-I compare report-level tone measures derived from both Naive Bayes and BERT classification.
-Following Huang et al.
-(2014), I manually label 10,000 sentences as training data for both models.
-Each sentence is categorized as positive, neutral, or negative using these pre-trained models, and I average the tone at the report level to create $Tone_{N}B$ and $Tone_{B}ERT$ measures.
-I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
-Table 6 presents a comparison of performance between text tone measures and text embeddings.
-Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures.
-The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al.
-(2014).
-A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, comparable with the 41 basis points increase in two-day abnormal return reported in Huang et al.
-(2014).
-Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics.
-Both coefficients of tone are significantly positive.
-Regression results with firm-fixed and year-fixed effects are consistent, as reported in column (3).
-Columns (4)-(6) summarizes regression results with BERT-based tone measures as explanatory variables.
-The coefficient estimates’ magnitude increases dramatically, along with significance levels.
-Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points.
-When both tone measures are included in an OLS regression in column (7), coefficient estimates of Naive Bayes-based tone measure become insignificant, suggesting that BERT model captures sentiment more accurately and effectively.
-Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of tone measures and embeddings.
-The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows.
-The $R^{2}_{\mathrm{OOS}}$ improves to 9.49% and 12.27% after incorporating forecast revisions and text embeddings.
-The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The difference between the $R^{2}_{\mathrm{OOS}}$ of models with and without tone measures is not statistically distinguishable from zero.
-The empirical results in Table 6 highlight two key advantages of LLMs over conventional NLP methods.
-First, tone measures extracted using LLMs better explain contemporaneous stock returns than those from the Naive Bayes algorithm.
-Second, neither BERT nor Naive Bayes tone measures add incremental value beyond text embeddings, suggesting that LLMs capture richer contextual information than simple tone metrics.
-3.2.6 Information Content with Alternative Machine Learning Models
-In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions.
-Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$.
-Table 7 presents results from PLS regressions.
-PLS effectively extracts common components that correlate with the target variable.
-If text embeddings contain strong common components that predict stock returns, PLS regression should capture this information.
-The PLS regressions achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to those in Ridge models.
-DM tests confirm that the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all years.
-Next, I replicate the analyses with the XGBoost algorithm.
-Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees.
-Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information.
-Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns.
-Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable.
-As summarized in Table 7, the XGBoost algorithm underperforms Ridge regressions with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76).
-The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
-I then proceed with Neural Network models, which form the core of LLMs and the transformer architecture.
-In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain.
-Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text.
-I construct the Neural Network estimations as an ensemble of five models for each NN configuration.
-As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms.
-Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al.
-(2022).
-The experiments with machine learning models yield noteworthy findings.
-First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns.
-Second, the superior performance of deep learning architectures emphasizes the importance of contextual information and word interdependencies in textual analyses.
+The sentiment analysis serves two objectives: to evaluate how LLMs perform compared to traditional NLP methods, and to assess the relative information content of embeddings versus tone measures.
+I compare tone measures derived from both Huang et al.
+(2014)’s Naive Bayes approach and BERT classification.
+I manually label 10,000 sentences as training data for both models.
+After training the models to categorize sentences as positive, neutral, or negative, I average these classifications at the report level to create $Tone\_NB$ and $Tone\_BERT$ measures.
+Table 7 contrasts the performance of text tone measures against text embeddings.
+Panel A presents OLS regression results of three-day $CAR$ on tone measures.
+Column (1) shows a coefficient estimate of 0.022 for $Tone\_NB$, aligning with Huang et al.
+(2014)’s reported 0.0208.
+A one standard deviation increase in text tone favorableness corresponds to a 54 basis point increase in $CAR$, comparable to their reported 41 basis point increase in two-day abnormal returns.
+Column (2) reveals significantly positive coefficients for tone measures for both Income Statement and non-Income Statement topics.
+The results remain consistent under firm-fixed and year-fixed effects in column (3).
+BERT-based tone measures (columns 4-6) demonstrate substantially larger coefficient estimates with higher significance levels.
+A one standard deviation increase in BERT-based tone favorableness yields a 95 basis point increase in three-day $CAR$.
+When both measures are included (column 7), Naive Bayes-based tone coefficients become insignificant, indicating BERT’s superior sentiment capture.
+Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across models.
+Tone inputs alone achieve $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68% in the first four rows.
+Including forecast revisions and text embeddings improves $R^{2}_{\mathrm{OOS}}$ to 9.49% and 12.27%, respectively.
+The panel’s final rows show that adding tone measures to text embedding models yields no statistically significant improvement in $R^{2}_{\mathrm{OOS}}$.
+The results highlight two key advantages of LLMs over conventional NLP methods: stronger explanatory power for stock returns through better sentiment capture, and the ability to extract richer textual information beyond simple sentiment measures.
+3.2.6 Alternative Machine Learning Models
+I examine the information content using three alternative machine learning algorithms, following the same procedure of Ridge regressions. Results are presented in Table 8.
+PLS regressions extract common components that correlate with the target variable.
+If text embeddings contain strong common components that correlates with stock returns, PLS should outperform.
+Using this method, I achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to Ridge models.
+DM tests confirm significance at the 1% level across all years.
+The XGBoost algorithm, which builds sequential trees to combine weak estimations into stronger ones, allows for testing whether text embeddings contain discrete informational components of varying importance.
+Despite its ability to capture nonlinear relationships, XGBoost underperforms Ridge regressions with an $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76).
+This relatively weaker performance suggests that linguistic information is distributed across the embedding space rather than concentrated in specific dimensions.
+Neural Network models, which underpin LLMs and transformer architectures, show superior performance.
+Given their interconnected node structure mirroring human neural networks, these models are well-suited to capture word relationships and contextual information.
+Using ensembles of five models for each configuration, Neural Networks achieve the highest $R^{2}_{\mathrm{OOS}}$ among tested algorithms, ranging from 10.42% to 12.12%.
+Performance improves with network depth, consistent with Kelly et al.
+(2022)’s virtue of complexity principle.
+Two key findings emerge from these experiments.
+First, the significant $R^{2}_{\mathrm{OOS}}$ across models demonstrates meaningful information content in report text, particularly given the low signal-to-noise ratio in stock returns.
+Second, although Neural Networks achieve the highest performance, most of the textual information can be captured by simple linear models like Ridge regressions.
 3.3 Information Value of Analyst Reports
-How to translate statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms?
-In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3.
-This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal, while accounting for price impact.
-In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
-I first summarize the value of information from analyst reports in Table 8, covering the period from 2015 to 2023.
-Both the mean and standard deviation are derived using the delta method.
-For S&P 100 constituent stocks, the average three-day window information value from sell-side analysts is \$0.47 million, with a standard error of \$0.05 million.
-The mean information value of text alone is \$0.38 million, and the mean value for revisions is \$0.34 million.
-Considering that a stock typically has reports released on an average of 15 days each year, the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm is \$6.89 million.
-The narrow standard errors and confidence intervals across different information values indicate consistent estimates.
-The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, indicating that \$1 million dollar order pushes the stock price up by 0.034%.
-These statistics illustrate that information from analysts is economically meaningful, with both textual information and revisions contributing significantly.
-Figure 4 describes the evolution of information value from analyst reports over time.
-The solid line represents the strategic value of combined numerical and textual information, while the dashed line shows the value of textual information alone.
-Both lines exhibit significant fluctuations over the period.
-Despite these fluctuations, there is a general upward trend in information value.
-For most of the period, the combined value is higher than the textual value alone, suggesting that numerical information often complements the textual content in analyst reports.
-The shaded area represents the 95% confidence interval.
-After 2020, there appears to be greater variability in the information value, as indicated by the wider confidence intervals and more pronounced peaks and troughs.
-This period corresponds with the global COVID-19 pandemic, which likely introduced additional uncertainty and volatility in financial markets, impacting the value of information from analyst reports.
-These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions.
-The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
-In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value.
-Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million.
-Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time.
+To translate the statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms, this section quantifies the information value of analyst reports using the methodology established in Section 2.3.
+This measure captures the economic profit generated from reduced price uncertainty, while incorporating price impact considerations.
+The measure’s construction implies that both higher information content and lower price impact contribute positively to the estimated information value.
+Table 9 presents the statistical evidence estimated with the delta method.
+For S&P 100 constituent stocks, sell-side analysts generate an average three-day window information value of \$0.47 million (SE = \$0.05 million).
+This value can be decomposed into its constituent components: textual content contributes \$0.38 million, while numerical revisions account for \$0.34 million.
+Extrapolating from the typical frequency of analyst coverage—approximately 15 report days per year—the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm amounts to \$6.89 million.
+The precision of these estimates is supported by consistently narrow standard errors and confidence intervals across different information value measures.
+The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, suggesting that a \$1 million order pushes the stock price up by 0.034%.
+The statistics demonstrate that analyst information value is economically meaningful, with both textual content and numerical forecasts contributing significantly.
+Figure 4 illustrates the dynamics of analyst report information value.
+The analysis tracks both the strategic value of combined numerical and textual information (solid line) and the isolated value of textual information (dashed line).
+While both measures exhibit substantial temporal variation, they demonstrate an overall upward trend.
+The consistently higher combined value relative to textual value suggests complementarity between numerical and textual components.
+The 95% confidence interval reveals increased variability in information value post-2020, characterized by wider confidence bounds and more pronounced peaks and troughs.
+The heightened volatility coincides with the COVID-19 pandemic period, which introduced substantial market uncertainty.
+The patterns demonstrate that the dynamic nature of analyst information value and its sensitivity to market conditions.
+The Online Appendix provides additional robustness checks.
+Table A7 presents alternative information value estimates using high-frequency realized return volatility measures and varied trade-signing algorithms, yielding estimates ranging from \$0.42 million to \$0.64 million.
+Figure A8 demonstrates that these alternative specifications produce consistent temporal patterns in information value evolution.
 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
-The information value of sell-side analysts likely vary in the cross-section of stocks.
-Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984).
-For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995).
-I plot the scatter of information value and market capitalization by stock in Figure A9.
-There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values.
-However, the relationship is not strictly linear, as there are notable exceptions.
-I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects.
-I also control for the book-to-market ratio in the regressions.
-Table 9 shows the results.
-Stocks with larger market equity have significantly higher information value from analysts.
-Economically, results in column (4) indicate a 1% increase in market equity is associated with a 0.864% increase in analyst information value.
-Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value.
-I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact of large stocks.
+Analyst information value may vary in the cross-section of stocks.
+Prior literature suggests that information processing and dissemination costs vary with firm size (Zeghal, 1984), and analyst opinions may trigger larger market reactions for smaller stocks due to their relatively limited public information environment (Stickel, 1995).
+Figure A9 illustrates this relationship through a scatter plot of information value against stock market capitalization.
+While the plot reveals a positive association between market equity and information value, the relationship exhibits notable nonlinearities.
+To formally examine this relationship, I estimate regressions of information value on stock size, incorporating firm and year fixed effects and the book-to-market ratio control.
+Table 10 shows that stocks with larger market equity have significantly higher information value from analysts.
+The economic magnitude is substantial: estimates from column (4) indicate that a 1% increase in market equity corresponds to a 0.864% increase in analyst information value.
+Columns (5)-(8) reveal that textual information value exhibits greater size sensitivity compared to the combined numerical and textual information value.
+To better understand the mechanisms driving this positive size elasticity, I decompose the information value into two components: explained return volatility and price impact.
 The decomposition can be expressed logarithmically as:
 $\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
-Panel B of Table 9 shows that the higher information value for large stocks stems primarily from their greater liquidity.
-Analyst reports of large stocks do not explain higher contemporaneous return volatility, yet their value increases with stock size due to enhanced liquidity.
+Panel B of Table 10 shows that higher analyst information value in larger stocks stems primarily from their greater liquidity.
+While analyst reports do not better explain return volatility in large stocks, the reduced price impact of liquid stocks enhances their information value.
 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
-Analysts who make bold forecasts incorporate their private information more completely and provide more relevant information to investors than those who herd in their forecasts (Clement and Tse, 2005).
-I examine how information value varies with analysts’ herding behavior.
-To investigate the cross-section of analyst characteristics, I regress the stock-analyst-day information value on an indicator of forecast boldness, controlling for firm-fixed effects, analyst-fixed effects, and year-fixed effects.
-I also control for broker size, as Jacob et al.
-(1999) suggests that larger brokerage firms provide better research.
-The analyst-specific information value measure is defined as
+Prior literature suggests that analysts who make bold forecasts incorporate private information more comprehensively than those who herd (Clement and Tse, 2005).
+This section examines the relationship between information value and analysts’ herding behavior.
+To analyze cross-sectional variation in analyst characteristics, I estimate regressions of stock-analyst-day information value on forecast boldness indicators, incorporating firm, analyst, and year fixed effects.
+Following Jacob et al.
+(1999), who suggests larger brokerage firms provide superior research, I include broker size as a control variable.
+The analyst-specific information value measure is defined as:
 $\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
-Table 10 presents the results.
-In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts.
-On average, information from bold reports is 29% more valuable than that from herding reports.
-This magnitude persists when examining the information value of report text alone.
-The insignificant coefficients for $Brokersize$ suggest that the value of analyst information is not related to brokerage firm size.
-A subsequent research question is whether the relationship between analyst boldness and information value is driven by higher report informativeness or lower price impact.
-In Panel B, I regress the two decompositions of information value on the boldness indicator.
-As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information.
-In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant.
+Table 11 presents the regression results.
+Panel A, columns (1)-(4), reveals significantly positive coefficient estimates on $Bold$, indicating more strategically valuable information when analysts provide bold forecasts.
+Quantitatively, information from bold reports demonstrates 29% higher value compared to herding reports—a magnitude that persists when examining textual content alone.
+The insignificant coefficients for $Brokersize$ suggest that broker size is related to analyst information value.
+To understand the underlying mechanisms, Panel B decomposes the relationship between analyst boldness and information value into informativeness and price impact components.
+As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts issue bold forecasts when possessing relevant private information.
+Columns (5)-(8) reveal no statistically significant relationship between boldness and price impact, suggesting that the enhanced value of bold forecasts stems primarily from their information content rather than market microstructure effects.
 3.3.3 The Value of Analyst Information around Earnings Announcements
 Previous evidence underlines the information content of analyst reports in the first week following earnings announcements.
-Accordingly, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
-Figure 5 illustrates the information value of analyst reports between earnings announcements on a weekly basis, from the first week after the earnings announcement up to the $13^{th}$ week.
-The information value of analyst reports is highest immediately following earnings announcements, particularly in the first week, peaking at a sample mean of \$0.84 million.
-This value sharply declines by the second week and stabilizes thereafter, with minor fluctuations.
-These patterns suggest that the market values analyst reports most highly immediately after earnings announcements, with diminishing importance in subsequent weeks.
-To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$).
-This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010).
-The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al.
-(2024).
-The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect analyst information value.
-Stock and year-fixed effects are included to control for time-invariant stock characteristics and market-wide trends.
-Table 11 presents regression results showing the influence of earnings announcements and trading volume on information value.
-Panel A indicates that the information value is significantly higher for reports released within one week of earnings announcements ($Week$), with positive and significant coefficients across all models.
-The interaction between $Week$ and $TradingVolume$ also shows a significant positive effect, suggesting that the information value increases with higher trading volumes around earnings announcements.
-The rationale is that investors may rely more on analyst opinions when there is greater disagreement in the interpretation of earnings information.
-Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable.
-Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact.
-These findings indicate that analyst reports provide more valuable insights in the first week after earnings announcements, especially during periods of high investor uncertainty.
-Investors value analysts’ ability to promptly interpret complex earnings information and make sense of the latest earnings numbers.
+Accordingly, this section evaluates the information value of analyst reports issued promptly following earnings announcements and beyond that.
+Figure 5 examines different samples of analyst reports categorized by their timing relative to earnings announcements, from reports released in the first week after earnings announcements to those issued in the $13^{th}$ week.
+Reports issued in the first post-earnings week exhibit the highest information value, with a sample mean of \$0.84 million.
+Reports from subsequent weeks show a sharp decline in value.
+This pattern again suggests that markets place premium value on timely post-earnings analyst reports.
+To formally test the observed pattern, I estimate regressions of log information value on an indicator variable $Week$, which identifies reports released within one week after earnings announcements.
+Following Banerjee (2011) and Beckmann et al.
+(2024), I include earnings announcement day trading volume ($TradingVolume$) as a proxy for investor disagreement.
+The model incorporates an interaction term ($Week\times TradingVolume$) to capture the joint effect of post-earnings timing and investor disagreement on analyst information value.
+The specification builds on Chen et al.
+(2010)’s finding that analysts’ interpretative role becomes more critical with increasing financial information complexity.
+The model includes stock and year fixed effects.
+Table 12 presents regression results.
+Panel A documents significantly higher information value for reports issued within the first post-earnings week, as evidenced by positive and significant coefficients on $Week$ across all models.
+The positive and significant interactions between $Week$ and $TradingVolume$ indicate that information value increases with trading volume on earnings announcement days.
+The results are consistent with investors placing greater reliance on analyst interpretations during periods of heightened disagreement.
+Panel B decomposes the earnings announcement effects into return variance and price impact, showing significant increases in both components.
+The coefficient magnitudes indicate that the timing effect is primarily driven by higher explained return volatility, rather than changes in price impact.
+These findings lead to conclude that investors value the most of analysts’ interpretation role during periods of heightened post-earnings uncertainty.
 4 Conclusion
-In this paper, I examine both quantitative and qualitative content of analyst reports using $R^{2}_{\mathrm{OOS}}$.
-The results show that report text contains more market-valued information content than quantifiable forecast revisions economically, especially right following earnings announcements.
-A Shapley value decomposition method reveals that analysts’ interpretation of realized income statements is the most valuable content within written reports.
-The text embedding representation extracts more comprehensive information from analyst reports beyond text sentiment.
-I further translate the information content into dollar terms to provide a sense of economic magnitude.
-Using a measure of explained return volatility relative to price impact, I find that the annualized expected profit from early information acquisition of the analyst reports for an average S&P 100 constituent firm is larger than \$6.89 million.
-While these results are economically meaningful, readers should exercise caution in generalizing beyond the S&P 100 sample.
+This paper examines both the quantitative and qualitative content of analyst reports.
+I find that report text contains more market-valued information content than quantifiable forecast revisions economically, especially shortly following earnings announcements.
+A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable component of written reports.
+I further translate the information content into economic terms.
+Using a measure of explained return volatility relative to price impact, I find that early acquisition of the analyst reports for an average S&P 100 constituent firm yields an annualized profit larger than \$6.89 million.
 Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
 These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets.
-The significant potential profit from early access to analyst reports raises concerns about selective distribution.
-Such practices would violate FINRA Rule 2241, which aims to ensure the objectivity of research and prevent the selective distribution of research reports to preferred clients ahead of others.
-It is crucial to emphasize that this study presents a hypothetical scenario without assessing or endorsing actual market practices.
-Rather, by quantifying the potential profit of privileged access to analyst information, this research emphasizes the importance of regulations in ensuring fair disclosure.
+The substantial potential profits from early access to analyst reports raise concerns about selective distribution practices—a direct violation of FINRA Rule 2241, which mandates research objectivity and prohibits preferential distribution to certain clients.
+This study presents a hypothetical scenario for quantifying information value rather than an assessment of market practices.
+By documenting the magnitude of potential profits from privileged access, my research emphasizes the importance of regulations for fair disclosure.
 A. Additional Figures
@@
 The company operates five distinct segments: Americas (71% of FY15 profit); Europe, Middle East, and Africa (4% of FY15 profit); China and Asia Pacific (12% of FY15 profit); Channel Development (13% of FY15 profit).
-Growth Strategies
-Visa has been especially active on the acquisition front over the last several months.
+Growth Strategies Visa has been especially active on the acquisition front over the last several months.
 Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle.
@@
 Valuation
-Our DCF derives an intrinsic value of ¿\$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC.
+Our DCF derives an intrinsic value of \$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC.
 We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment.
 Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
-Investment Thesis
-The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road.
+Investment Thesis The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road.
 As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high.

### v4

@@
 I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports.
-Using embeddings from state-of-the-art large language models, I show that textual information in analyst reports explains 10.19% of contemporaneous stock returns out-of-sample, a value that is economically more significant than quantitative forecasts.
+Using embeddings from state-of-the-art large language models, I show that qualitative information in analyst reports explains above 10% of contemporaneous stock returns out-of-sample, a value that is more economically significant than quantitative forecasts.
 I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns.
 The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power.
-Expressing these findings in economic terms, I estimate that early acquisition of analysts’ reports can yield significant profits.
-Analysts’ information value peeks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data.
+Expressing these findings in economic terms, I estimate that early acquisition of analyst reports can yield significant profits.
+Analyst information value peaks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data.
 JEL Classification: G11, G14, G24
@@
 Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021).
-This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
-Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data.
-Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018).
+This raises a fundamental question: Do analysts generate value for their clients, or are they merely peddling expensive noise?
+Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of text.
+Prior studies have shed light on the incremental information value of report text through sentiment analysis (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018).
 However, analyst reports are meant to provide contextual information beyond the simple sentiment dimension.
-The context has been perceived at least as important as the quantifiable outputs.
+The context has been perceived to be at least as important as the quantitative outputs.
 According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations.
 Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
-The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024).
-Consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning.
-The question now becomes: do investors disregard obvious facts and figures in favor of details in lines?
-I investigate the question through three general steps: analyst output representation, econometric modeling, and value decomposition.
-Analyst outputs include both quantitative information and qualitative information.
-The quantitative data consists of numerical forecasts like stock recommendations, earnings forecasts, and target prices.
-For qualitative textual information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
+The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (see, for example: Chen et al., 2022; Li et al., 2024).
+Consider the following two statements: “Reported EPS of \$0.94 misses consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning.
+The question now becomes: When presented with both, do investors prioritize the explicit figures or the explanatory narrative?
+I investigate the research question through three general steps: analyst output representation, econometric modeling, and value decomposition.
+For quantitative information, I consider three numerical forecasts, including stock recommendations, earnings forecasts, and target prices.
+For qualitative information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
 With structured representation, I evaluate how well analyst information explains stock returns as evidence of information content.
-A significant challenge in projecting stock returns onto text embeddings is the risk of in-sample overfitting due to high dimensionality.
-To address this issue, I employ Ridge regressions.
-The simple model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market.
-Additionally, I use out-of-sample R-squared to evaluate model performances.
+Due to the high dimensionality of LLM embeddings, projecting stock returns on them bears overfitting risk.
+To address this issue, I employ the Ridge regression.
+The simple linear model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market.
 I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR_{[-1,+1]}$), surpassing the 9.01% explained by numerical information.
 When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using either information type alone.
-These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
-The informativeness of analyst reports varies substantially with forecast changes.
-Analyst reports accompanied by forecast revisions provide significantly more valuable information to financial markets compared to reiterations.
+These findings remain robust across various LLM representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
+The informativeness of analyst reports varies substantially depending on whether they are accompanied by forecast revisions.
+Reports that revise earnings forecasts or recommendations convey significantly more value to investors than those that simply reiterate prior views.
 For reports containing recommendation revisions, the combination of numerical and textual information generates an out-of-sample $R^{2}$ of 22.63%.
-In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$ is not statistically significant, indicating they provide minimal new information to market participants.
-Previous research emphasizes analysts’ role in information processing around corporate earnings announcements( e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017).
-My analysis on samples during and outside earnings announcement periods reveals that analyst reports’ information content peaks in the first week following earnings releases.
-The finding aligns with Huang et al.
-(2018), who argue that prompt responses to corporate disclosures provide valuable insights to clients.
-Using distinct models trained on samples during and outside earnings announcement windows, the out-of-sample $R^{2}$ approximately doubles in the first week following earnings announcement dates compared to other periods.
-The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings information for investors.
-To address the concern that my results merely capture analysts’ “piggybacking” on earnings conference call information or post-earning-announcement drift (PEAD), I control for both earnings surprises and the latest earnings conference call transcripts using text embeddings.
-The results show that analyst reports provide distinct and valuable insights even after accounting for these alternative explanations.
-As raised in the review paper of Bradshaw et al.
-(2017), identifying the qualitative attributes of analyst reports that investors find most salient remains largely unsolved.
-I develop a systematic approach to quantify the importance of different report content.
-By exploiting the additive property of text embeddings and implementing a Shapley value decomposition, I attribute the explanatory power to 17 major topics discussed in analyst reports.
-To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports.
-The results reveal that Income Statement Analyses, particularly interpretations of realized earnings, account for over half of the reports’ explanatory power in terms of out-of-sample $R^{2}$.
-These results reinforce the critical role of analysts in interpreting financial data and constructing the narratives and context essential for investment decisions.
-To translate statistical measures of information content into economic terms, I adopt Kadan and Manela (2024)’s framework to quantify the dollar value of analyst information for strategic investors.
-The multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination.
-The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts.
-Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to price impact, representing the cost of trading on this information.
-The analysis reveals substantial economic value in analyst information.
-For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports amount to \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when combined.
-Based on an average of 15 report days annually, the aggregate annualized information value reaches \$6.89 million—a conservative lower bound according to Kadan and Manela (2024).
-This value increases for large-cap stocks, bold analysts, and reports released promptly following earnings announcements.
+In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$.
+Previous research emphasizes analysts’ role in processing information around corporate earnings announcements (e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017).
+Consistent with this literature, my analysis shows that the information content of analyst reports peaks in the first week following earnings releases.
+Comparing models trained separately on post-earnings and non-earnings periods, I find that the out-of-sample $R^{2}$ approximately doubles during the one-week earnings announcement window.
+The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings news for investors.
+To address the concern that the explanatory power may simply reflect “piggybacking” on earnings conference call information or post-earnings-announcement drift (PEAD), I control for both earnings surprises and the latest conference call transcripts.
+The results confirm that analyst reports provide distinct and valuable insights.
+An unresolved question in the literature, highlighted by Bradshaw et al.
+(2017), is which qualitative aspects of analyst reports investors find most informative.
+This study addresses this gap by developing a systematic approach to quantify the relative importance of various types of report content.
+Specifically, I leverage the additive structure of text embeddings and implement Shapley value decomposition to attribute the explanatory power of analyst reports across 17 prevalent topics.
+The results reveal that Income Statement Analyses, especially interpretations of realized earnings, account for over half of this explanatory power.
+These results reinforce analysts’ critical role in interpreting financial data and constructing the narratives essential for investment decisions.
+To translate the statistical informativeness of analyst reports into economic significance, I adopt the Kadan and Manela (2025) framework to quantify the dollar value of analyst information for strategic investors.
+Their multi-agent model provides a realistic basis for assessing the economic impact of analyst report dissemination.
+The information value represents the total expected profits accruing to strategic investors who either: (1) gain early access to analyst views, or (2) independently derive insights comparable to those from analysts.
+Intuitively, the measure reflects both explainable return variance (the uncertainty reduction from analyst information) and the price impact (the cost of trading on such information).
+The analysis indicates substantial economic value in analyst information.
+For an average S&P 100 stock, the estimated profit over a three-day window for strategic investors with early access (two days prior to public release) is approximately \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when both sources are combined.
+Assuming an average of 15 report days per year, this implies an annualized information value of \$6.89 million, a conservative lower bound according to Kadan and Manela (2025).
+The value is notably higher for large-cap stocks, reports from bold analysts, and those released promptly following earnings announcements.
 My paper relates to three strands of literature.
-It first connects to recent studies of market reactions to security analyst reports (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024).
-Earlier investigations primarily rely on manual text coding and sentiment analysis, which are labor-intensive and may overlook nuanced insights.
-My work demonstrates the superiority of state-of-the-art LLMs in generating contextualized textual representations.
-Compared to Huang et al.
-(2014)’s Naive Bayes tone classification, LLMs offer two main advantages: better capture of overall report sentiment and extraction of richer information beyond tone.
-Moving from simple tone analysis to comprehensive LLM embeddings boosts the $R^{2}$ from 3% to over 10%.
-The substantial improvement in explanatory power highlights LLMs’ potential to transform textual analysis.
-Additionally, the topic decomposition approach provides novel insights into analysts’ value by revealing their role in interpreting financial data.
-Second, this work contributes to the growing field of machine learning and artificial intelligence in finance.
-Recent studies like Chen et al.
-(2022), Li et al.
-(2024), Jha et al.
-(2024), and Beckmann et al.
-(2024) have showcased the potential of implementing LLMs in a financial paradigm.
-However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns.
-By exploiting the additive property of LLM embeddings and developing a Shapley value decomposition approach, my research provides a new framework for explainable AI.
-Third, this paper advances the understanding of analyst tipping - the practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010).
-Tipping allows brokerages to offset research costs by earning commissions from clients who benefit from the privileged information.
-Prior studies, such as those by Green (2006) and Kadan et al.
-(2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%.
-This paper extends by quantifying the economic dollar value of such privileged information.
+First, this paper offers several extensions to prior research that highlight the incremental information in report narratives and the interpretative role of analysts (e.g., Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018).
+Specifically, I show that more contextualized textual representations not only provide significant incremental information but can also surpass the explanatory power of analysts’ traditional quantitative forecasts.
+Additionally, I demonstrate that the informational value derived from the report text extends substantially beyond simple sentiment measures.
+To illustrate, transitioning from basic textual tone measures to comprehensive LLM embeddings increases the out-of-sample $R^{2}$ from approximately 3% to over 10%.
+Furthermore, to my knowledge, this is the first study to systematically identify and quantify the specific qualitative attributes in analyst reports that investors find most salient.
+The research complements the existing literature by providing a more granular understanding of which textual elements drive the value of analyst reports and when that value peaks, such as after earnings announcements or when accompanying forecast revisions.
+Second, this paper makes a distinct methodological contribution to the expanding literature on AI and machine learning in finance.
+While the application of LLMs in financial paradigms is promising (e.g., Chen et al., 2022; Li et al., 2024; Lopez-Lira and Tang, 2023; Jha et al., 2024), their inherent ‘black-box’ nature presents significant hurdles for model validation and reliability.
+To address this critical interpretability gap, this paper implements a Shapley value decomposition framework specifically designed for LLM embeddings plus topic modeling.
+By leveraging the additive structure of these embeddings, this approach systematically attributes an LLM’s performance to constituent textual features, thereby enhancing model transparency and offering a robust pathway toward explainable AI in financial applications.
+Third, this paper offers new insights into analyst tipping.
+It provides a rigorous quantification of the economic dollar value accruing from privileged early access to research, a practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010).
+While prior work (e.g., Green, 2006; Kadan et al., 2018) documents significant abnormal returns for clients with early access, this study further translates the information advantages into tangible monetary terms.
+The dollar value offers a more concrete measure for understanding the incentive of selective disclosure and the scale of wealth transfer involved.
 The rest of the paper is organized as follows.
-Section 2 introduces methodology, measures of analyst information value, and interpretation framework.
+Section 2 introduces the methodology, measures of analyst information value, and the interpretation framework.
 Section 3 describes the data and empirical results.
@@
 2 Methodology
-This section describes the key methodologies: the use of large language models for topic modeling and text embeddings, estimation of information content and strategic information value, and decomposition metrics for evaluating topic-specific information value.
-[^1] The details of machine learning models used to explain contemporaneous stock returns are provided in the Online Appendix C.
-2.1 Large Language Models
-The analysis of analyst reports leverages three advanced generative AI models.
-A fine-tuned BERT model classifies sentences according to predefined criteria, ChatGPT extracts and synthesizes major topics through prompting, and LLaMA generates high-dimensional vector embeddings to represent report content.
-2.1.1 ChatGPT Prompting
-To avoid ad-hoc classification, I employ ChatGPT prompting to extract exclusive and meaningful topics.
-The goal is to identify primary topics covered in analyst reports and assign each sentence to a single topic category.
-In the first step, I use a random sample of 100 analyst reports and request ChatGPT to analyze the content and identify high-level categories with the following prompt:
+This section describes the key methodologies: using large language models for text representation and topic modeling, and the decomposition techniques to evaluate the information value of individual topics.
+2.1 Text Representation
+A core step in analyzing the information value of analyst reports is constructing structured, numerical representations of textual content.
+LLMs offer a significant advancement by providing richer, more contextualized linguistic representations than traditional methods.
+As illustrated in Figure A1 in the online Appendix, the process begins with text tokenization, where text segments are mapped to unique identifiers based on a predefined vocabulary.
+These tokens are then processed through a transformer architecture.
+The transformer’s key output consists of dense, continuous vectors that capture both semantic and syntactic information.
+To generate report-level representations, I compute embeddings as the average across all tokens and transformer layers:
+$y^{emb}=\frac{1}{N_{k}N_{l}}\sum_{l=1}^{N_{l}}\sum_{i=1}^{N_{k}}{e}_{i,l},$ (1)
+where $e_{i,l}$ denotes the contextual embedding of token $i$ generated by layer $l$, $N_{k}$ is the total number of tokens in the report, $N_{l}$ is the number of layers from which embeddings are aggregated.
+The embedding representation provides two salient advantages over traditional sentiment measures.
+First, LLMs position each token within a high-dimensional semantic space.
+This maintains complex relationships both within and across sentences, producing tone measurements that are substantially more nuanced and reliable than those generated by dictionary-based approaches.
+Second, these embeddings encode context-dependent meaning and narrative structure, thereby conveying information that extends well beyond simple positive-negative classifications and conveys far richer linguistic signals.
+For this study, I use MetaAI’s pre-trained LLaMA-2-13B model and its associated tokenizer.
+With an input length limit of 4,096 tokens, the model accommodates almost all analyst reports, which have a median length of 1,393 tokens and a mean of 2,055 tokens.
+The resulting contextual embeddings serve as the foundation for downstream tasks such as topic attribution and explaining stock returns.
+2.2 Topic Modeling
+This section introduces the methodology used to identify prevalent themes in analyst reports and to classify each sentence into a single, primary topic.
+The initial challenge is selecting a topic modeling approach that effectively avoids ad-hoc classifications and ensures economically meaningful interpretations.
+I exclude manual topic construction to minimize researcher subjectivity.
+Given the availability of sentence-level embeddings, clustering these embeddings based on semantic proximity is a straightforward procedure to consider.
+Accordingly, I expriment a standard K-Means algorithm for initial assessment.
+[^1] Topic classification results from the K-Means experimentation are available upon request.
+The unsupervised clustering approach is not adopted as the final methodology for two primary reasons.
+First, due to the industry-sensitive nature of analyst reports, K-Means tended to generate topics heavily aligned with particular sectors (e.g., pharmaceuticals, banking), whereas this study seeks to identify common thematic categories that span across industries.
+Second, as also noted by Aleti and Bollerslev (2025), topics derived from purely data-driven clustering tend to be messy and difficult to interpret from an economically meaningful perspective.
+To provide a set of distinct and interpretable topics, this study uses targeted prompting of the ChatGPT-4o model.
+The objective of this approach is to identify a comprehensive set of primary topics consistently discussed in analyst reports and to assign each sentence to one of these predefined categories.
+In the first step, I use a random sample of 100 analyst reports and request ChatGPT-4o model to analyze the content and identify high-level categories with the following prompt:
 Prompt 1: Please read the provided text file of sell-side analyst reports carefully.
-What are high-level mutually exclusive topics covered in these reports?
+What are high-level, mutually exclusive topics covered in these reports?
 Make sure that each sentence from the text file can be assigned to one of the topics.
@@
 It ends up with 16 categories defined in Table LABEL:tab:_topic.
-A category labeled “None of the Above” is included to accommodate sentences that do not fit into predefined topics.
-To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category.
-Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
-In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences.
-To begin with, I employ the GPT-4-Turbo model to categorize each sentence in a random sample of 17,028 sentences from 100 analyst reports into a single, most relevant topic.
+A category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the 16 topics.
+Table LABEL:tab:_topic_example offers illustrative sentences of each category.
+In the second step, I build a sentence classifier to assign each sentence to one topic.
+To begin with, I employ the GPT-4-turbo model to categorize each sentence in a random sample of 100 analyst reports into a single, most relevant topic.
 Prompt 2: Please read the following sentence from a sell-side analyst report of the company {firm} ({ticker}) carefully.
@@
 Here is the sentence from the analyst report: {sentence}.
-This labeled sample is used to fine-tune a BERT model, which then classifies all sentences in the corpus.
-The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample.
+I then fine-tune a BERT-based classifier using the labeled sample and classify all 6,975,114 sentences in the sample.
 The distribution of topics is relatively stable across years, as illustrated in Figure 1.
-Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively.
+Income Statement Analyses and Financial Ratios are the most frequently discussed topics, accounting for 17.23% and 15.65% of all sentences, respectively.
 These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics.
-Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least frequent topics.
-Several factors contribute to these observed frequencies.
-The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase.
-Executive Summary sections appear infrequently because they are difficult to identify at the sentence level.
-While ESG discussions are minimal overall, they increase notably after 2020, reflecting growing market interest in sustainability issues.
-2.1.2 Text Representations
-I need a structured representation of writtrn reports.
-Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps.
-Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters.
-Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters.
-These tokens are then mapped to unique identifiers based on a predefined vocabulary learned during model training.
-To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”.
-This granular approach allows for nuanced interpretation of complex terms.
-For instance, ‘next-gen workloads’ becomes five separate tokens: ‘next’, ‘-’, ‘gen’, ‘work’, and ‘loads’.
-At the heart of LLMs lies the transformer architecture.
-A core output of the transformer is text embeddings - dense, continuous vectors that capture semantic and syntactic information.
-Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia.
-These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
-For this analysis, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI.
-The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.
-The resulting contextual embeddings capture nuanced report information and are suitable for downstream tasks like sentiment analysis and economic modeling.
-2.2 Information Content
+While ESG discussions are minimal overall (0.23%), they increase notably after 2020, reflecting growing market interest in sustainability issues.
+I validate the topic modeling procedure based on three key evidence.
+First, the fine-tuned BERT classifier demonstrates high performance, achieving 89% accuracy on the test sample.
+Given the challenge of distinguishing among 17 topic categories, it demonstrates a notably strong performance.
+Second, the 16 GPT-generated topics provide comprehensive coverage of the report content.
+Less than 5% of sentences fall into the fallback category “None of the Above”, indicating that the topic set captures nearly all relevant content.
+Third, Figure A2 presents word clouds showing the most frequent terms within each topic.
+The word frequencies align closely with expected topic content.
+For example, “Income Statement Analysis” is strongly associated with terms like revenue, EPS, and sales, confirming the semantic coherence of the classification.
+2.3 Shapley Value Decomposition
+Next, I propose a methodology to decompose the value of different topics within analyst reports.
+As a starting point, a simple approach is aggregating the embeddings of tokens corresponding to each topic and measure their corresponding value independently.
+However, this method has two significant drawbacks.
+First, the self-attention architecture of transformer models creates cross-topic contamination.
+It means, each token’s embedding incorporates contextual information from all other tokens in the report, including those assigned to different topics.
+Second, topics are not independent.
+Their interactions can be a primary driver of value.
+For example, the combination of a valuation discussion and an earnings analysis may be more informative than either alone.
+To mitigate the first concern of cross-topic contamination, I construct sentence-segmented embeddings.
+Specifically, each sentence is processed by the LLM individually.
+The procedure ensures that the resulting token embeddings $\left(e_{i}\right)$ are contextualized only by the words within their own sentence.
+A report-level embedding $\left(y^{emb}\right)$ is then constructed as the token-weighted average of these sentence embeddings:
+$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}S_{i}^{emb},$ (2)
+where $S_{i}^{emb}$ is the average token embeddings of sentence $i$, $n$ is the number of sentences in the report, and $Token_{i}$ is the number of tokens in sentence $i$.
+To account for topic interdependency, I implement the Shapley Additive exPlanations (SHAP) framework.
+Originating in cooperative game theory (Shapley, 1953), the Shapley value fairly allocates a collective payoff among contributing members.
+This concept is widely used in machine learning to ensure fair attribution of model predictions to individual features (Lundberg and Lee, 2017; Chen et al., 2022).
+The SHAP framework is ideal for topic value attribution for two reasons.
+First, it satisfies the additivity (or efficiency) property, ensuring that the total information content of a report is fully distributed among all constituent topics without omissions or double-counting.
+Second, it inherently accounts for topic interdependencies by evaluating the marginal contribution of each topic across all possible topic combinations.
+The Shapley value $\left(\varphi_{p}\right)$ for a specific topic $p$, which quantifies its contribution to the overall out-of-sample explanatory power ($R_{\text{oos }}^{2}$), is defined as:
+$\varphi_{p}(R_{oos}^{2})=\sum_{S\subseteq P\setminus{p}}\frac{|S|!(|P|-|S|-1)!}{|P|!}\left[R_{oos}^{2}(S\cup{p})-R_{oos}^{2}(S)\right],$ (3)
+where $P$ is the set of all topics in a report.
+The summation is over all subsets $S$ of topics excluding $p$.
+The term $R_{\text{oos }}^{2}(S)$ denotes the out-of-sample $R^{2}$ of a model whose input is the report embedding constructed using only sentences belonging to topics in the subset $S$ [^2] I define missing topic embeddings as zero vectors, effectively removing them from the model input.
+For a given subset of topics $S$, the corresponding embedding is constructed using the token-weighted average formula above, but summed only over sentences classified into topics within $S$.
+3 Data and Empirical Results
+This section describes the data and presents empirical results of the analyst information value.
+I begin with an examination of analyst information content and its decomposition across topics.
+I then translate these findings into economic terms by quantifying the strategic dollar value of analyst information.
+3.1 Data
+The initial sample comprises 223,091 sell-side analyst reports of S&P 100 constituent stocks from 2000 to 2023 downloaded from the Mergent Investext database.
+After converting the reports from their original PDF format to plain text using Adobe Acrobat, the extracted text is segmented into individual sentences.
+To focus on stock-relevant discussions, I fine-tune a BERT classifier to identify and remove boilerplate content.
+[^3] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions.
+Following Li et al.
+(2024), I create a training set by sampling one report from each of the top 20 brokerage firms and manually labeling sentences.
+The fine-tuned BERT model achieves 97.31% accuracy on a held-out test set.
+The Investext database provides detailed metadata for analyst reports, including analyst identities and release dates.
+Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window.
+I first match lead analyst names to unique analyst identifiers (AMASKCD) following the procedure in Li et al.
+(2024).
+I then associate each report with corresponding earnings forecasts, target prices, and stock recommendations in I/B/E/S.
+Since multiple reports may be issued between a forecast announcement and revision, I adopt the windowing strategy of Huang et al.
+(2014), linking reports to forecasts issued between two days prior to the I/B/E/S announcement date (ANNDATS) and two days after the revision date (REVDATS).
+[^4] I/B/E/S records both the forecast announcement date (ANNDATS) and the revision date (REVDATS), which marks when the forecast becomes outdated.
+To isolate the distinct information content of analyst reports, I exclude reports released on earnings announcement days.
+The final matched sample consists of 122,252 analyst reports on S&P 100 stocks, authored by 1,305 analysts across 140 brokerage firms.
+Table 2 provides summary statistics of analyst reports across years and the Fama-French 12 industries, including the number of reports, unique brokerage firms, and analysts.
+The table also reports average report length, measured by page and token counts.
+The temporal distribution of reports follows the pattern documented by Bonini et al.
+(2023), with a peak in 2013 followed by a steady decline, likely driven by regulatory changes such as the Dodd-Frank Act and MiFID II (Bradshaw et al., 2017).
+To quantify market reactions, I estimate stock-level price impacts using intraday data from the NYSE TAQ database for the out-of-sample period of January 2, 2015, to December 28, 2023.
+All monetary values are adjusted for inflation using the Consumer Price Index (CPI) and are normalized to December 2020.
+Additionally, I match each analyst report with the latest earnings conference call transcript to control for concurrent events.
+Transcripts for S&P 100 firms are sourced from Seeking Alpha and matched to firm-quarter earnings announcement dates (RDQ) from Compustat, resulting in a sample of 7,909 firm-quarter earnings conference calls from 2005 to 2023.555Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts.
+A comparative analysis is presented in Section 3.2.2.
+I include a set of firm-specific numerical features in the machine learning models, categorized as follows.
+Firm characteristics are sourced from CRSP and Compustat.
+[^6] Detailed variable definitions are provided in Table A2, and their summary statistics are in Table A3.
+Report-Level Measures: $REC_{REV}$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
+$EF_{REV}$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 trading days prior to the report’s release.
+$TP_{REV}$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 trading days prior to release.
+$Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both.
+$SR$ denotes stock recommendations in the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell).
+$ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 trading days prior to release.
+$PriorCAR$ measures cumulative abnormal return over [-10, -2] window preceding the report release.
+Firm-Level Measures: $Size$ is quantified by market equity.
+Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value.
+Earnings surprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value.
+$Miss$ is a binary indicator equal to 1 if $SUE<0$.
+$TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding.
+Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI).
+$Fluidity$ measures firms’ product market competition following Hoberg et al.
+(2014).
+Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the stock’s Fama-French 48 industry experiences negative monthly returns in the bottom quintile.
+Macroeconomic Measures: $TimeTrend$ is a temporal variable measuring the number of years since the start of the sample period.
+3.2 Information Content of Analyst Reports
+I start by investigating market reactions to analyst quantitative and qualitative outputs.
 Following Asquith et al.
-(2005), I consider $R^{2}$ as a proxy for information content.
-To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations annually using an expanding window:
+(2005), I consider $R^{2}$ as a proxy for the information content of analyst reports, which quantifies how well analyst information explains contemporaneous stock returns.
+I estimate Ridge regression models annually using an expanding window.
+The specification is given by:
 $CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (4)
 […]
-where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ for stock $i$ at day $t$.
-The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
-Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ using:
+where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stock $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ of stock $i$ at day $t$.
+Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return.
+For reports issued on non-trading days, $t=0$ is adjusted to the next available trading day on CRSP.
+In-sample $R^{2}$ can be inflated due to overfitting, particularly when employing machine learning techniques and high-dimensional embeddings.
+Following Gu et al.
+(2020), I calculate the out-of-sample $R^{2}$ using:
 $R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum_{(i,t)\in\mathcal{T}}CAR_{[-1,+1],it}^{2}},$ (6)
@@
 $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the proportion of return variance attributable to analyst information.
-To assess the statistical significance of the $R_{\mathrm{oos}}^{2}$, I implement a modified version of the Diebold and Mariano (2002) test for predictive accuracy comparison.
+The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
+I implement a modified version of the Diebold and Mariano (2002) (DM) test to compare the statistical significance of the $R_{\mathrm{oos}}^{2}$.
 Given the strong cross-sectional dependence inherent in the stock-level analysis, I follow Gu et al.
-(2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional averages of prediction errors.
-2.3 Information Value
-Following Kadan and Manela (2024), I evaluate the value of information using a measure that approximates information value in a strategic setting where multiple informed investors possess imperfect signals:
-$\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
-where $\widehat{\Omega}_{it}$ represents the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price on day t-2.
-$\widehat{r}_{ijt}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$.
-Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return.
-$N$ represents the number of analyst reports for stock $i$ on day $t$.
-To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$).
-The order flow is derived from changes in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals, with trades classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991).
-[^2] The algorithms from Ellis et al.
-(2000) and Chakrabarty et al.
-(2007) are also used in robustness tests.
-The regression specification is:
-$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
-where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda).
-In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$.
-The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
-The measure reflects the ratio of return volatility to price impact and is grounded in the theoretical framework of Back et al.
-(2000), which is a continuous-time extension of Kyle (1985) model.
-Under risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of asset price.
-The intuition is illustrated with an extended single-auction Kyle model with incomplete information in Online Appendix D.
-I employ the measure for two key reasons.
-First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information.
-Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
-[^3] Kadan and Manela (2024) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply in the analyst context.
-Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced.
-To address this challenge, I apply the delta method to estimate the mean and variance of information value over sample $s$.
-The derivation of the first-order estimator using the delta method is detailed in Online Appendix E.
-2.4 Shapley Value Decomposition
-To quantify the relative contribution of different topics, I implement the Shapley Additive exPlanations (SHAP) methodology.
-SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves.
-It was later applied for machine learning model interpretation (Lundberg and Lee, 2017 and Chen et al., 2022).
-The approach allows for attributing an analyst report’s overall explanatory power to its constituent topics.
-Leveraging the additive nature of text embeddings, I develop a novel decomposition method to systematically evaluate the relative importance of various topics in analyst reports.
-The report-level embeddings are generated as the average of embeddings across all tokens and layers. The report embeddings can be formalized as:
-$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
-where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ denotes the total token count of the report.
-Given the tokens can be classified into distinct topics, the embeddings can be fully decomposed into topic-specific components:
-$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
-where $P$ denotes the total number of topics, $y_{p}^{emb}$ represents the embedding for topic $p$, $N_{p}$ is the number of tokens classified as topic $p$, and $i_{p}$ indicates the index of tokens within topic $p$.
-A consideration in the transformer architectures is the self-attention mechanism, which allows the model to compute relative importance weights between tokens.
-This architecture enables each token in the input sequence to “talk” with each other.
-Specifically, the embedding of token $i$ incorporates contextual information from token $j$, even when these tokens belong to distinct topics.
-To mitigate this cross-topic attention effect, I process embeddings sentence by sentence to ensure the token embeddings only reflect intra-sentence contextual information.
-I then take the token-weighted mean of sentence embeddings as the report representation, which I designate as sentence-segment embeddings:
-$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
-where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embeddings of the $i$-th sentence.
-The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding quantifies the information content of context across sentences.
-Subsequently, I compute the Shapley value across topics using sentence-segmented embeddings. The Shapley value is defined as:
-$\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
-where the sum extends over all subsets $S\subseteq P\backslash\{p\}$, including the empty set.
-$y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$.
-The $R_{\text{oos}}^{2}$ is calculated for each unique combination of topic embeddings.
-[^4] Shapley value requires a defined input for missing $X$ values.
-I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation.
-This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ across topics, revealing their relative contributions to explaining market reactions.
-3 Data and Empirical Results
-In this section, I present the empirical analysis of analyst report information, beginning with an examination of information content and its decomposition across topics.
-I then quantify the strategic dollar value of analyst information.
-The results emphasize analysts’ income statement analyses following earning announcements.
-3.1 Data
-I refer to the Mergent Investext database for 223,091 sell-side analyst reports for S&P 100 firms spanning the years 2000 to 2023.
-The reports, initially in PDF format, are converted to plain text using Adobe Acrobat.
-The extracted text is then segmented into sentences.
-To ensure a clean dataset focused on firm-relevant discussions, I fine-tune a BERT model to filter out boilerplate content.
-[^5] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions.
-Following Li et al.
-(2024), I construct a training dataset by randomly sampling one report from each of the top 20 brokerage firms by report volume and manually label each sentence.
-The BERT model fine-tuned for boilerplate classification achieves a 97.31% accuracy rate on test data.
-The Mergent Investext Database provides analyst identification and report release dates.
-To connect Investext with the I/B/E/S database, I first match lead analyst names to their corresponding Analyst IDs (AMASKCD) following Li et al.
-(2024).
-I then match individual reports to corresponding EPS forecasts, price targets, and recommendation files in I/B/E/S.
-Given that multiple reports may be issued between a forecast announcement and its subsequent revision, I adopt Huang et al.
-(2014)’s approach and define the matching window as spanning from two days before the announcement date to two days after the revision date.
-[^6] I/B/E/S records two timestamps for each analyst forecast: the announcement date (ANNDATS) when the new forecast is issued and the old one is revised, and the revision date (REVDATS) when the forecast becomes invalid or undergoes revision.
-Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window.
-Additional firm characteristics are sourced from CRSP and Compustat.
-[^7] Detailed variable definitions and summary statistics are provided in Table A2 and Table A3, respectively.
-After the matching procedure and excluding reports released on earnings announcement days, the final sample consists of 122,252 reports covering S&P 100 stocks, produced by 1,305 analysts across 140 distinct brokerage firms.
-Table 2 provides summary statistics of analyst reports, including the number of reports, unique brokerage firms, and analysts across years and the Fama-French 12 industries.
-It also reports average page and token counts per report as measures of report length.
-The temporal distribution of analyst reports aligns with patterns documented in Bonini et al.
-(2023), while the average page count per report corresponds with findings in Huang et al.
-(2014).
-[^8] The data reveals a peak in the number of analyst reports in 2013, followed by a gradual decline.
-This trend is likely influenced by the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and MiFID II in Europe (Bradshaw et al., 2017).
-Stock-level price impacts are estimated using intraday data from the NYSE TAQ database for an out-of-sample period spanning January 2, 2015, to December 28, 2023.
-To ensure temporal comparability, all prices are adjusted for inflation using the Consumer Price Index (CPI), normalized to December 2020 as the base period.
-To isolate the information value of analyst outputs, I control for alternative information sources available to investors.
-In Section 3.2.2, I conduct a comparative analysis between earnings conference call content and subsequent analyst reports.
-Conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts..
-These transcripts provide detailed documentation of executive presentations and manager-analyst interactions during Q&A sessions.
-I sample transcripts of S&P 100 constituent firms from 2005 to 2023, matching them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat.
-The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks, yielding a final sample of 7,909 firm-quarter-level conference calls.
-The set of numerical variables used as input in machine learning models includes:
-Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
-$EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
-$TP\_REV$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 days pre-release.
-$Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both.
-$SR$ denotes stock recommendations following the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell).
-$ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 days pre-release.
-$Prior\_CAR$ measures cumulative abnormal return over [-10, -2].
-Firm-Level Measures: $Size$ is quantified by market equity.
-Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value.
-Following Huang et al.
-(2018), Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value.
-$Miss$ is a binary indicator for positive $SUE$.
-$TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding.
-Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI).
-$Fluidity$ measures firms’ product market fluidity.
-Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the firm’s industry (defined by Fama-French 48 classification) experiences negative monthly returns in the bottom quintile.
-Macroeconomic Measures: $TimeTrend$ is a temporal variable measuring the number of years since the start of the sample period.
-3.2 Information Content of Analyst Reports
-This section investigates market reactions to analysts’ quantitative and qualitative outputs.
-I employ the Shapley value decomposition methodology introduced in Section 2.4 to attribute the information content across distinct topics.
+(2020) and compare the cross-sectional means of squared prediction errors to evaluate the improvement over a benchmark.
 3.2.1 Quantitative and Qualitative Information
-Table 3 compares the information content embedded in numerical and textual measures.
-The analysis employs Ridge regression as specified in Equations 1 and 2, with ${CAR}_{[-1,+1]}$ centered at the report release date serving as the dependent variable.
-For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database.
-The estimation utilizes distinct input configurations with an expanding window approach, where training sets span from 2000 to the year preceding each target year.
-Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
-Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for forecast revision input, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst revisions.
-The incorporation of additional 14 numerical features in column (2) yields a modest improvement to 9.08% $R^{2}_{\mathrm{OOS}}$.
-Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the Diebold-Mariano test t-statistics reported in columns (2) and (4).
-Columns (5) and (6) reveal that Ridge models with text embeddings achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text.
-The $R^{2}_{\mathrm{OOS}}$ remains stable from 2015 to 2023, except during the 2020 pandemic recession.
-A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating substantial information content of report text relative to forecast revisions.
-Addressing the look-ahead bias concern raised by Sarkar and Vafa (2024) regarding pre-trained LLMs, I examine the $R^{2}_{\mathrm{OOS}}$ in 2023.
-Since the employed LLMs were trained on pre-2023 data, this period represents true out-of-sample performance, validating the robustness of the main findings.
-Columns (7) and (8) examine the model performance when integrating both forecast revisions and text embeddings.
-This combined approach achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively.
-The DM t-statistics of 3.95 and 3.77 for comparisons between combined input versus revision input and text input demonstrate that report text and forecast revisions contain distinct, complementary information sets valued by investors.
-The explanatory power might stem from quantitative information embedded in the numbers mentioned within the text.
-To address this concern, I conduct robustness checks using content with numbers removed.
+Table 3 evaluates the information content of numerical versus textual measures.
+Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) 17 numerical measures, forecast revisions plus the report-level, firm-level, industry-level, and macroeconomic indicators detailed in Section 3.1; (3) text embeddings, 5,120-dimensional full-context representations; and (4) a hybrid input combining forecast revisions with report embeddings.
+Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for the forecast revisions, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) that documents the price impact of analyst revisions.
+The incorporation of 14 additional numerical features in column (3) yields a modest improvement to an $R^{2}_{\mathrm{OOS}}$ of 9.08%.
+Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the DM t-statistics reported in columns (2) and (4).
+Columns (5) and (6) show that text embeddings alone achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text.
+The explanatory power of textual inputs remains relatively stable across the 2015–2023 period, with the exception of a dip during the 2020 pandemic recession.
+A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating that report text offers comparable or marginally larger information content relative to quantitative forecasts.
+One potential concern is that the explanatory power might stem from the numbers mentioned in the text.
+To address this concern, I conduct robustness checks using text embeddings generated by reports with all numbers removed.
 Table A4 presents the performance of models using number-free text embeddings.
-The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% significantly exceeds that of quantitative revision measures (t-statistic = 2.40), suggesting that the superior performance of text-based models is not driven by embedded numerical content.
-I examine the scaling law relationship by comparing LLMs with varying parameters, including BERT, OpenAI text-embedding-3-small, and LLaMA-3 models.
-The LLaMA models feature significantly more parameters than the BERT and OpenAI models.
-Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 benchmark.
-The input in Panel B is report text embeddings, comparable to column (3) in Panel A.
-A positive t-statistic indicates superior performance relative to the LLaMA-2 model.
-While smaller models like BERT and OpenAI underperform LLaMA-2, the larger LLaMA-3 model shows only marginally better performance with an $R^{2}_{\mathrm{OOS}}$ of 9.66%, though the improvement is statistically insignificant.
-These results demonstrate the robustness of analyst report text information content across different LLMs and illuminate the relationship between model scale and performance capabilities.
-To validate the robustness of contemporaneous measures, I examine alternative $CAR$ windows, including 2-day and 5-day intervals following Huang et al.
-(2014) and Asquith et al.
-(2005), respectively.
-Table A5 shows consistent $R^{2}_{\mathrm{OOS}}$ magnitudes for $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% during the 2015-2023 out-of-sample period.
-Table A6 presents the cross-industry analysis of analyst report text information content across Fama-French 12 classifications.
-Five industries exhibit $R^{2}_{\mathrm{OOS}}$ exceeding 10%: Shops, Other, Manufacturing, Chemicals, and Durables.
-Conversely, Non-Durables, Telecommunications, Energy, and Utilities sectors show significantly lower $R^{2}_{\mathrm{OOS}}$ and DM t-statistics.
-This pattern, visualized in Figure A3, suggests that industries with complex products and services derive greater value from detailed analyst insights, while those operating in stable or regulated environments benefit less from incremental analyst information.
-To evaluate the economic significance, I estimate OLS regressions specified as follows:
+The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% exceeds that of quantitative revision measures, with a t-statistic of 2.40, suggesting that the superior performance of text-based models is not driven by embedded numerical information.
+To ensure the results aren’t look-ahead biased by the models having seen future information (Sarkar and Vafa, 2024), I evaluate model performance in 2023.
+Since all the language models in this study were trained only on pre-2023 text, this setting ensures a truly out-of-sample test.
+The $R^{2}_{\mathrm{OOS}}$ of 9.30% reinforces the robustness of the main findings and confirms that the explanatory power is not driven by information leakage.
+Columns (7) and (8) report the model performance when combining forecast revisions with text embeddings.
+The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively.
+The DM t-statistics of 3.95 and 3.77 demonstrate that both the report text and forecast revisions contain distinct, complementary information valued by investors.
+To validate that the findings are not contingent on a specific model choice, I test their robustness across both alternative language models used to generate embeddings and the machine learning algorithms used for estimation.
+First, I assess the sensitivity to the choice of LLMs by replicating the analysis with embeddings from several alternatives: BERT, OpenAI’s text-embedding-3-small, and LLaMA-3-8B model.
+As shown in Panel B of Table 3, while smaller models like BERT and OpenAI underperform the LLaMA-2 benchmark, the similar-size LLaMA-3 model achieves a comparable $R^{2}_{\mathrm{OOS}}$ of 9.66%.
+Second, I test the LLaMA-2 embeddings across three conceptually distinct machine learning algorithms: Partial Least Squares (PLS), XGBoost, and Neural Networks (NNs).
+The results, summarized in Table A5, show that the significant explanatory power is successfully identified by all models, confirming that the main conclusions are not an artifact of using Ridge regression.
+While NNs achieve the highest performance, the strong results from linear models like PLS and Ridge indicate that a substantial portion of the information captured in the embeddings is linearly related to stock returns.
+Collectively, the tests demonstrate that the significant, market-moving information contained in analyst report narratives is robustly captured across a range of language models and machine learning algorithms.
+To ensure the findings are not contingent on the specific three-day event window used in the main analysis, I test the robustness of results to alternative return horizons.
+I re-estimate Ridge models using several different dependent variables for the CAR, including a one-day window $\left(AR_{[0]}\right)$, a two-day window $\left(CAR_{[0,+1]}\right)$, and a five-day window $\left(CAR_{[-2,+2]}\right)$.
+As presented in Table A6, the explanatory power of the textual embeddings remains stable and economically significant across all tested windows (out-of-sample $R^{2}_{\mathrm{OOS}}$ ranges from 6.92% to 7.57%), confirming that the central finding is robust to the choice of the contemporaneous return measurement window.
+I next explore whether the information content of analyst reports varies systematically across industries.
+Table A7 presents the performance of text-based models across the Fama-French 12 industry classifications.
+A pronounced heterogeneity emerges.
+The model demonstrates substantial explanatory power (defined as $R^{2}_{\mathrm{OOS}}>10\%$) in five industries: Shops, Manufacturing, Chemicals, Durables, and a diversified “Other” category.
+Conversely, the information content is significantly lower in sectors such as Non-Durables, Telecommunications, Energy, and Utilities.
+This cross-sectional pattern, also visualized in Figure A3, supports the economic intuition that the value of detailed analyst reports is greatest for firms with higher operational complexity or informational uncertainty.
+In contrast, for industries operating in more stable, regulated, or commoditized environments, the incremental value of qualitative analyst insights appears to be more limited.
+To assess the economic significance and directly compare the information content of different sources, I conduct a series of “horse race” regressions.
+The dependent variable is the actual three-day cumulative abnormal return, $CAR_{[-1,+1],it}$, and the primary independent variables are the out-of-sample predictions generated by Ridge regressions.
+The OLS specification is:
 $CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (7)
@@
 The estimates for 2015, for example, are derived from Ridge regression models trained on 2000-2014 data.
-The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension, while $\widehat{CAR}_{rev}$ is estimated using three forecast revision measures under identical procedures.
-Table 4 presents sample summary statistics and regression results.
-The analysis sample is restricted to observations with valid current and previous stock recommendations, earnings forecasts, and target prices.
-Panel B’s Column (1) confirms that recommendation revisions, earnings forecast revisions, and target price revisions all contribute significantly to explaining contemporaneous abnormal stock return variations, consistent with Brav and Lehavy (2003) and Huang et al.
-(2014).
-Column (2) shows that $\widehat{CAR}{txt}$ alone achieves an $R^{2}$ of 10.5%.
-When both analyst revision measures and $\widehat{CAR}{txt}$ are included as regressors in column (3), the $R^{2}$ increases to 14.9%.
-In economic terms, one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness corresponds to 3-day abnormal return increases of 10, 40, 90, and 120 basis points, respectively.
-Columns (4) and (5) examine $\widehat{CAR}{rev}$, which aggregates all revision metrics.
-The $\widehat{CAR}{rev}$-only model achieves an $R^{2}$ of 8.9%, while combined inputs reach 14.7%.
-These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
-Columns (6)-(7) introduce $SUE$ and $Prior_{C}AR$ controls to capture earnings surprises and recent news.
-The stability of $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ coefficients suggests that market-valued analyst information is distinct from recent earnings news and does not merely reflect reactions to past returns.
-3.2.2 Earnings Announcements
-Building on research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), I analyze the information content of sell-side analyst reports surrounding earnings announcements.
+The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single number, while $\widehat{CAR}_{rev}$ aggregates the information from the three forecast revision measures into a single variable.
+Table 4, Panel B, presents the regression results.
+[^7] The sample for this analysis is restricted to observations with valid data for all current and prior forecast revision variables.
+As a baseline, Column (1) confirms that the individual revision measures for recommendations, earnings forecasts, and target prices are all significant predictors of abnormal returns, consistent with prior literature (Brav and Lehavy, 2003; Asquith et al., 2005).
+The subsequent columns conduct the main “horse race”.
+A model with only the estimated CAR from text $\left(\widehat{CAR}_{txt}\right)$ as a regressor (Column 2) achieves an $R^{2}$ of $10.5\%$.
+A model with only the estimated CAR from revisions $\left(\widehat{CAR}_{rev}\right)$ yields an $R^{2}$ of $8.9\%$ (Column 4).
+When both predictors are included in the same regression (Column 5), the adjusted $R^{2}$ rises to $14.7\%$.
+The coefficients on both $\widehat{CAR}_{txt}$ and $\widehat{CAR}_{rev}$ remain highly statistically significant, underscoring that the textual content of analyst reports provides substantial explanatory power over and above that contained in forecast revisions.
+The economic magnitudes of these effects are also significant.
+Based on the coefficients from standardized variables, a one-standard-deviation increase in the predicted CAR from report text $\left(\widehat{CAR}_{txt}\right)$ is associated with a 120 basis point increase in the actual 3-day abnormal return.
+The equivalent impacts for recommendation, earnings, and target price revisions are 10, 40, and 90 basis points, respectively.
+Finally, Columns (6) and (7) test the robustness of these findings by including controls of the latest earnings surprises ($SUE$) and recent stock performance ($PriorCAR$).
+The stability of the coefficients on both $\widehat{CAR}_{txt}$ and $\widehat{CAR}_{rev}$ indicates that the information captured by analyst reports is distinct from and incremental to publicly available information about recent earnings and past returns.
+3.2.2 The Timing of Analyst Reports around Earnings Announcements
+Building on prior research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010), this section investigates the information content of analyst reports around earnings announcements.
 Following Chen et al.
-(2010), I group analyst reports into 13 weekly bins around earnings announcement dates.
-Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for weekly report samples.
-Consistent with Huang et al.
-(2018), analyst reports cluster predominantly in the immediate post-earnings announcement period.
-The information content only exhibits strong significance only in the first week, achieving an $R^{2}_{\mathrm{OOS}}$ of approximately 10% with a t-statistic of 10.17.
-The variation in performance across weekly bins may reflect Ridge models capturing dominant patterns in analyst reports issued around earnings announcements.
-To address potential temporal variation, I train separate Ridge models on analyst reports released during earnings announcement periods and those issued beyond these periods, allowing the models to extract distinct information content based on report timing.
-Panel B in Table 5 compares the model performance within and beyond earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows.
-Models trained on earnings announcement periods consistently outperform those trained on non-announcement periods.
-The $R^{2}_{\mathrm{OOS}}$ and t-statistics exhibit a declining pattern as windows lengthen, decreasing from 11.84% to 7.29% over the earnings announcement week, suggesting markets particularly value swift analyst responses to earnings news.
-To address concerns that analysts merely restate earnings call information and that market reactions primarily reflect earnings announcements rather than analyst insights, I conduct a comparative analysis of information content between these two sources.
-I focus on reports released one day after earnings announcements, when analysts would be most likely to simply repackage earnings information.
-Column (1) of Panel C presents the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using earnings conference call transcript embeddings as input.
-[^10] Due to the length of earnings conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for text embeddings in Panel C, which accommodates 8,192 tokens versus the 4,096-token limit of the LLaMA-2-13B model used elsewhere.
-Texts are truncated at these limits to maximize information retention.
-The $R^{2}_{\mathrm{OOS}}$ reaches 4.20% with a DM t-statistic of 5.16.
-Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods.
-Using report text embeddings from the day following earnings announcements, this analysis achieves an $R^{2}_{\mathrm{OOS}}$ of 9.72% (column 2), substantially outperforming conference call transcripts alone.
-The combination of analyst report and earnings conference call embeddings yields further improvements, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% (column 3).
-A comparison t-statistic of 5.12 between transcript-only and combined models demonstrates analysts’ substantial incremental contribution.
-These findings collectively refute the notion that analysts merely restate earnings information, instead highlighting their distinct and timely value to the market.
+(2010), I group analyst reports into 13 weekly bins relative to the earnings announcement date.
+Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics for each bin.
+Report issuance peaks in the week immediately following earnings announcements.
+This is also the period when reports are most informative.
+The $R_{\mathrm{OOS}}^{2}$ reaches nearly $10\%$ (DM t-statistic $=10.17$ ), indicating significant market reaction to analysts’ timely interpretations of earnings news.
+To account for the possibility that analyst information differs depending on its timing relative to an EA, I estimate separate Ridge models for reports issued within different event windows.
+This approach allows the models to learn distinct relationships between text and returns for reports published in information-rich versus information-sparse periods.
+Panel B of Table 5 compares the performance of models trained on reports from varying windows ($1,2,3$, and $7$ days post-EA).
+The models trained on reports issued most promptly after earnings consistently outperform those trained on other periods.
+Notably, the $R_{\mathrm{OOS}}^{2}$ declines from $11.84\%$ to $7.29\%$ as the training sample window expands from one to seven days, suggesting that the market places the highest value on immediate, timely analysis following corporate disclosures.
+A critical question is whether analysts provide unique insights or merely act as conduits, reiterating information already disclosed in earnings conference calls.
+To test this, I conduct a direct comparison of the information content of analyst reports and the latest earnings conference call transcripts.
+[^8] Given the significant length of conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for this analysis, as its 8,192-token context window is larger than the 4,096-token limit of the LLaMA-2-13B model used elsewhere.
+Texts are truncated at these limits.
+I deliberately focus on reports released one day after an EA, creating a “hard test” for the analyst’s contribution, as this is when both analyst reports and contemporaneous returns are more likely to be influenced by earnings conference call information.
+I estimate Ridge regressions that combine both information sources:
+$CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{{\text{AI+EA}}}+\epsilon_{ijt},$
+[…]
+where $y_{ijt}^{\mathrm{AI}+\mathrm{EA}}$ is created by concatenating the embeddings from analyst report $j$ for stock $i$ on day $t$ with the embeddings from that stock’s most recent earnings conference call.
+Panel C of Table 5 presents the results.
+A model using only the text embeddings of earnings conference call transcripts achieves an $R^{2}_{\mathrm{OOS}}$ of $4.20\%$ (Column 1).
+The transcript data starts in 2005.
+For a fair comparison, I retrain analyst report models over the same 2005-2023 sample period.
+This model achieves an $R_{\mathrm{OOS}}^{2}$ of $9.72\%$ (Column 2), substantially outperforming the explanatory power of the earnings conference call transcripts alone.
+Finally, a model that combines embeddings from both analyst reports and earnings conference calls yields an $R_{\mathrm{OOS}}^{2}$ of $11.96\%$ (Column 3).
+The highly significant improvement over the transcript-only model (DM t-statistic = 5.12) demonstrates the substantial, incremental information content provided by analyst reports.
+These findings collectively refute the notion that analysts merely restate earnings information.
+Instead, they emphasize analysts’ role in distilling and contextualizing corporate disclosures in ways that are uniquely informative to the market.
 3.2.3 What Content in Analyst Reports is Valued?
-Having established the significance of report text information, I examine which specific content components most influence market response.
-I employ the Shapley value methodology to decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 topics detailed in Table LABEL:tab:_topic.
-The analysis uses sentence-segmented embeddings as report representations to mitigate cross-topic contextual effects.
-When switching from full-context to sentence-segmented embeddings, the $R^{2}_{\mathrm{OOS}}$ decreases by 2.25%, quantifying the magnitude of inter-sentence contextual information.
+Having established that analyst report text contains significant value-relevant information, this section investigates which specific content categories drive this informativeness.
+Using the Shapley value methodology, I decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 report topics pre-defined in Table LABEL:tab:_topic.
+To minimize the confounding effects of cross-topic attention, the analysis uses sentence-segmented embeddings rather than full-context embeddings.
+This adjustment results in a 2.25 percentage point drop in $R^{2}_{\mathrm{OOS}}$, a figure that quantifies the value of inter-sentence contextuality.
 Figure 2 illustrates the Shapley value decomposition, revealing a clear hierarchy in topic importance.
-Income Statement Analyses and Financial Ratio Analyses emerge as the most valuable components, contributing 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$, respectively.
-Investment Thesis and Valuation form a second tier of importance, with Shapley values of 1.76% and 1.51%.
-The remaining topics show minimal or negative Shapley values, indicating limited market reactions.
-The findings’ robustness is supported by several additional analyses.
-Figure A4 shows that normalizing Shapley values by sentence count and token length maintains Income Statement Analyses’ dominance.
-Figure A5 demonstrates consistency in topic rankings across different industries.
-The evolution of text-based $R^{2}_{\mathrm{OOS}}$ is depicted in Figure A6 of the Online Appendix.
-While a notable decline in $R^{2}_{\mathrm{OOS}}$ occurs between 2020 and 2021, likely reflecting COVID-19 pandemic disruptions in financial markets, the scaled Shapley value for Income Statement analysis (shown by the red line) exhibits remarkable stability.
-This persistence suggests that analyst commentary on income statements maintains its informational value even during market turbulence.
-To disentangle whether topic importance stems from analysts’ expertise in certain areas (supply effect) or investors’ response (demand effect), Figure A7 compares Shapley values across analyst experience levels and brokerage firm sizes.
-The consistent topic hierarchy across all four subsamples (experienced/inexperienced analysts and large/small brokerages) indicates that topic rankings are not driven by variations in analyst expertise.
-I proceed to examine analysts’ dual roles in income statement analysis: their discovery role in collecting and generating otherwise unavailable information and their interpretation role in integrating and processing available information into more interpretable signals (Huang et al., 2018; Blankespoor et al., 2020).
-To facilitate the analysis, I categorize Income Statement sentences based on: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income.
-This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
+Income Statement Analysis emerges as the single most valuable component.
+The Shapley value indicates it accounts for a relative contribution of 67% to the total $R^{2}_{\mathrm{OOS}}$, followed by Financial Ratio Analysis as the second most crucial topic.
+A second tier of importance includes Investment Thesis and Valuation, with Shapley values of 1.76% and 1.51%.
+The remaining topics exhibit minimal or even negative Shapley values, suggesting little contribution to reports’ market-moving power.
+The robustness of this topic hierarchy is confirmed through several additional tests.
+Normalizing the Shapley values by sentence count and token length (Figure A4) confirms that the dominance of Income Statement Analysis is not merely an artifact of size effect.
+Furthermore, this ranking remains remarkably consistent across all Fama-French 12 industries (Figure A5).
+The Online Appendix (Figure A6) shows that while the absolute predictive power of text fluctuated over time, dipping during the COVID-19 pandemic, the relative contribution of Income Statement Analysis remained stable, underscoring its enduring information value.
+To disentangle whether the topic importance stems from analyst expertise in certain areas (a supply-side effect) or from consistent investor demand for specific information (a demand-side effect), I examine how topic importance varies across different analyst and report characteristics.
+Figure A7 presents the Shapley value decompositions for six subsamples.
+The topic hierarchy proves to be remarkably consistent across all comparison groups: experienced versus inexperienced analysts; analysts at large versus small brokerage firms; and reports that primarily use DCF-based versus multiple-based valuation models.
+This pervasive consistency suggests that the dominance of certain topics, such as Income Statement Analysis, does not merely reflect analysts’ differing abilities, professional experience, or chosen valuation methodologies.
+Rather, the evidence indicates that the hierarchy is driven by the market’s broad and persistent demand for these specific types of fundamental analysis.
+Given its paramount importance, I conduct a deeper analysis of the “Income Statement Analysis” topic.
+I examine analysts’ dual roles: a discovery role, where they generate or report otherwise unavailable information, and an interpretation role, where they process and contextualize available information into more digestible signals (Huang et al., 2018; Blankespoor et al., 2020).
+I sub-classify each sentence in this topic along two dimensions: (1) Information Acquisition vs.
+Interpretation and (2) Realized vs.
+Expected Income, using the classification framework detailed in Prompt 3.
 Prompt 3: Please read the following sentence from a sell-side analyst report for the company {firm} ({ticker}) and classify it based on two criteria:
@@
 Realized Income (0): Sentences referring to concrete, historical results from completed periods. Example: ‘The company’s revenue for the fiscal year 2022 was \$500 million.’
-Expected Income (1): Sentences containing subjective predictions, or expectations for future periods.
+Expected Income (1): Sentences containing subjective predictions or expectations for future periods.
 Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
 Output your classification as two comma-separated numbers: the first for Information Type (0 for Acquisition, 1 for Integration) and the second for Time Reference (0 for Actual, 1 for Forecast).
-Output format: InfoType, TimeRef
 Sentence to classify: {report_sent}.
-Figure 3 provides a granular breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, including Income Statement subcategories.
-Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw figures.
-The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on actual results than forward-looking projections.
-These patterns emphasizes that markets particularly value analysts’ ability to interpret realized income, beyond mere information reporting or future expectations.
-These findings align with the results from section 3.2.2, where analyst reports released immediately after earnings announcements demonstrate the highest information content.
-Taken together, these results suggest that the market places the highest value on analysts who can rapidly digest earnings announcement information and provide insightful interpretations.
-3.2.4 Comparison of Revision Reports and Reiteration Reports
-Given the established relationship between market prices and analyst forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I examine the differential information content between revision and reiteration reports.
-Following the criteria of Huang et al.
-(2014), I classify reports as revisions when released within two days of the I/B/E/S forecast announcement date.
-This classification is applied independently across three forecast targets—recommendations, target prices, and earnings.
-Therefore, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
-Table 6 presents out-of-sample model performance across reiteration and revision samples.
-The frequency of revisions varies considerably: earnings forecast revisions are most common (56.57% of reports), followed by target price revisions (31.10%), while recommendation revisions are relatively rare (2.29%).
-Recommendation revisions exhibit particularly strong explanatory power.
-Quantitative forecast revision measures achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99% when accompanying recommendation revisions, while report text yields 16.80%.
-The combination enhances explanatory power to 22.63%.
-This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
-Target price revisions, noted for their incremental information value due to the limitations of discrete recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), show marked differences between revisions and reiterations.
-While reiteration reports demonstrate minimal information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58%), revision reports exhibit substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
-Earnings forecast revisions show a similar pattern.
-Reiteration reports yield insignificant $R^{2}_{\mathrm{OOS}}$, while revision reports demonstrate significant explanatory power.
-The combination of numerical and textual inputs for revision reports achieves an $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-statistic = 11.42), highlighting revisions’ superior value over reiterations.
-Overall, Table 6 shows that report content carries greater information content when accompanying forecast revisions rather than reiterations, with recommendation revision reports exhibiting the highest value.
-3.2.5 Comparison with Analyst Report Sentiment
-The sentiment analysis serves two objectives: to evaluate how LLMs perform compared to traditional NLP methods, and to assess the relative information content of embeddings versus tone measures.
-I compare tone measures derived from both Huang et al.
-(2014)’s Naive Bayes approach and BERT classification.
-I manually label 10,000 sentences as training data for both models.
-After training the models to categorize sentences as positive, neutral, or negative, I average these classifications at the report level to create $Tone\_NB$ and $Tone\_BERT$ measures.
-Table 7 contrasts the performance of text tone measures against text embeddings.
-Panel A presents OLS regression results of three-day $CAR$ on tone measures.
-Column (1) shows a coefficient estimate of 0.022 for $Tone\_NB$, aligning with Huang et al.
-(2014)’s reported 0.0208.
-A one standard deviation increase in text tone favorableness corresponds to a 54 basis point increase in $CAR$, comparable to their reported 41 basis point increase in two-day abnormal returns.
-Column (2) reveals significantly positive coefficients for tone measures for both Income Statement and non-Income Statement topics.
-The results remain consistent under firm-fixed and year-fixed effects in column (3).
-BERT-based tone measures (columns 4-6) demonstrate substantially larger coefficient estimates with higher significance levels.
-A one standard deviation increase in BERT-based tone favorableness yields a 95 basis point increase in three-day $CAR$.
-When both measures are included (column 7), Naive Bayes-based tone coefficients become insignificant, indicating BERT’s superior sentiment capture.
-Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across models.
-Tone inputs alone achieve $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68% in the first four rows.
-Including forecast revisions and text embeddings improves $R^{2}_{\mathrm{OOS}}$ to 9.49% and 12.27%, respectively.
-The panel’s final rows show that adding tone measures to text embedding models yields no statistically significant improvement in $R^{2}_{\mathrm{OOS}}$.
-The results highlight two key advantages of LLMs over conventional NLP methods: stronger explanatory power for stock returns through better sentiment capture, and the ability to extract richer textual information beyond simple sentiment measures.
-3.2.6 Alternative Machine Learning Models
-I examine the information content using three alternative machine learning algorithms, following the same procedure of Ridge regressions. Results are presented in Table 8.
-PLS regressions extract common components that correlate with the target variable.
-If text embeddings contain strong common components that correlates with stock returns, PLS should outperform.
-Using this method, I achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to Ridge models.
-DM tests confirm significance at the 1% level across all years.
-The XGBoost algorithm, which builds sequential trees to combine weak estimations into stronger ones, allows for testing whether text embeddings contain discrete informational components of varying importance.
-Despite its ability to capture nonlinear relationships, XGBoost underperforms Ridge regressions with an $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76).
-This relatively weaker performance suggests that linguistic information is distributed across the embedding space rather than concentrated in specific dimensions.
-Neural Network models, which underpin LLMs and transformer architectures, show superior performance.
-Given their interconnected node structure mirroring human neural networks, these models are well-suited to capture word relationships and contextual information.
-Using ensembles of five models for each configuration, Neural Networks achieve the highest $R^{2}_{\mathrm{OOS}}$ among tested algorithms, ranging from 10.42% to 12.12%.
-Performance improves with network depth, consistent with Kelly et al.
-(2022)’s virtue of complexity principle.
-Two key findings emerge from these experiments.
-First, the significant $R^{2}_{\mathrm{OOS}}$ across models demonstrates meaningful information content in report text, particularly given the low signal-to-noise ratio in stock returns.
-Second, although Neural Networks achieve the highest performance, most of the textual information can be captured by simple linear models like Ridge regressions.
+Figure 3 presents the results of this granular decomposition.
+Analysts’ interpretation of income statement data accounts for nearly three times the explanatory power of their acquisition and reporting of raw data.
+This suggests that investors value analytical context and insight far more than mere financial numbers.
+Similarly, discussions of realized (historical) income contribute three times as much to the $R^{2}_{\mathrm{OOS}}$ as discussions of expected (future) income, indicating the market’s strong preference for analysis grounds in actual financial outcomes over forward-looking speculation.
+These results reinforce the findings from Section 3.2.2 that the market reacts most strongly to reports issued immediately following earnings announcements.
+This section reveals that the most valued content within those reports is the interpretation of realized performance.
+Taken together, the evidence suggests that the primary value-add of sell-side analysts lies in their ability to rapidly digest complex, newly-released corporate disclosures and provide timely, insightful interpretations of realized income.
+3.2.4 Comparison of Revision and Reiteration Reports
+This section investigates the differential informativeness of analyst reports conditional on whether they contain forecast revisions or simple reiterations.
+While the market impact of forecast revisions is well-documented (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I test whether the overall information content of analyst reports is significantly greater in revision versus reiteration events.
+Following the classification methodology of Huang et al.
+(2014), I identify a report as a revision report if it is released within two days that a new forecast is recorded in the I/B/E/S database.
+This classification is applied independently to recommendations, target prices, and earnings forecasts, acknowledging that a single report may contain a revision for one forecasting target (e.g., target price) while reiterating another (e.g., recommendation).
+Table 6 presents the out-of-sample performance, estimated separately for revision and reiteration report subsamples.
+Descriptive statistics reveal that revisions are relatively common for earnings forecasts (56.6%) and target prices (31.1%) but are rare for stock recommendations (2.3%).
+The subsequent analysis demonstrates that the market response is overwhelmingly concentrated in revision reports.
+Among the three forecasting targets, recommendation revisions are the most potent information events.
+For reports containing upgrades or downgrades, the combined numerical and textual information achieves an exceptionally high $R_{\mathrm{OOS}}^{2}$ of $22.63\%$.
+This finding aligns with prior literature documenting the significant market impact of recommendation changes (Loh and Stulz, 2011; Bradley et al., 2014).
+Reports accompanied by target price revisions also show strong explanatory power.
+While reiteration reports produce a modest combined $R^{2}_{\mathrm{OOS}}$ of 3.58%, revision reports attain 20.88% (t-stat = 11.18).
+This stark contrast underscores the incremental information value of specific target price changes beyond the coarser recommendations (Bradshaw, 2002; Brav and Lehavy, 2003).
+Earnings forecast revisions follow a similar pattern.
+Reiteration reports offer little explanatory power, whereas revision reports achieve a combined $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-stat = 11.42), confirming that the market responds specifically to forecast revision events.
+In summary, the evidence in Table 6 consistently demonstrates that the informational value of analyst reports is heavily concentrated in those that revise explicit quantitative forecasts.
+Among these, the rarest recommendation changes, appear to convey the strongest signal to the market.
+3.2.5 Comparison of Analyst Report Sentiment Measures
+This section conducts a sentiment analysis to serve two primary purposes: first, to benchmark the performance of an LLM-based sentiment classifier (BERT) against a Naive Bayes algorithm used in Huang et al.
+(2014); and second, to assess the incremental information content of high-dimensional text embeddings relative to a single sentiment dimension.
+Specifically, I examine tone derived from a Naive Bayes classifier (${Tone}_{{NB}}$) and a BERT-based sentiment classifier (${Tone}_{{BERT}}$).
+Both models are trained on a manually labeled dataset of 10,000 sentences categorized as positive, neutral, or negative.
+I calculate report-level sentiment scores for each report as the number of positive sentences minus the number of negative sentences, scaled by the total number of sentences.
+Panel A of Table 7 presents OLS regressions of the CAR on these tone measures.
+In Column (1), I replicate the original analysis of Huang et al.
+(2014).
+The coefficient on ${Tone}_{{NB}}$ is 0.022, closely matching their estimate of 0.021.
+A one-standard-deviation increase in this tone measure is associated with a 54 basis point increase in CAR, comparable to the 41 basis point impact they document over a two-day window.
+Column (2) adds granularity by separating tone into Income Statement and non-Income Statement content, both yielding statistically significant effects.
+Column (3) confirms the robustness of these findings with firm and year fixed effects.
+Columns (4)-(6) show that the BERT-based tone measure has substantially stronger explanatory power.
+A one-standard-deviation increase in ${Tone}_{BERT}$ corresponds to a 95 basis point increase in CAR.
+In a direct “horse race” regression including both measures (Column 7), the coefficient on ${Tone}_{{BERT}}$ remains large and highly significant, while the coefficient on ${Tone}_{NB}$ becomes insignificant.
+This indicates that the more advanced, context-aware BERT model captures more value-relavant text sentiment than the Naive Bayes classifier.
+Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across distinct inputs.
+Tone measures alone offer modest explanatory power, with $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68%.
+Combining tone with forecast revision measures increases performance to 9.49%.
+Further integrating text embeddings lifts it to 12.27%.
+Notably, adding tone measures to a model already incorporating text embeddings yields no significant incremental improvement.
+Together, the findings highlight two key advantages of LLMs.
+First, for the specific task of classifying sentiment, LLMs like BERT more accurately capture nuanced tone and generate stronger signals than traditional NLP methods.
+Second, LLM embeddings are informationally superior to single sentiment scores.
+The fact that tone adds no value once embeddings are included suggests that the high-dimensional embeddings already capture the sentiment dimension, in addition to the complex, context-rich information that simple sentiment indicators miss.
 3.3 Information Value of Analyst Reports
-To translate the statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms, this section quantifies the information value of analyst reports using the methodology established in Section 2.3.
-This measure captures the economic profit generated from reduced price uncertainty, while incorporating price impact considerations.
-The measure’s construction implies that both higher information content and lower price impact contribute positively to the estimated information value.
-Table 9 presents the statistical evidence estimated with the delta method.
-For S&P 100 constituent stocks, sell-side analysts generate an average three-day window information value of \$0.47 million (SE = \$0.05 million).
-This value can be decomposed into its constituent components: textual content contributes \$0.38 million, while numerical revisions account for \$0.34 million.
-Extrapolating from the typical frequency of analyst coverage—approximately 15 report days per year—the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm amounts to \$6.89 million.
-The precision of these estimates is supported by consistently narrow standard errors and confidence intervals across different information value measures.
-The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, suggesting that a \$1 million order pushes the stock price up by 0.034%.
-The statistics demonstrate that analyst information value is economically meaningful, with both textual content and numerical forecasts contributing significantly.
-Figure 4 illustrates the dynamics of analyst report information value.
-The analysis tracks both the strategic value of combined numerical and textual information (solid line) and the isolated value of textual information (dashed line).
-While both measures exhibit substantial temporal variation, they demonstrate an overall upward trend.
-The consistently higher combined value relative to textual value suggests complementarity between numerical and textual components.
-The 95% confidence interval reveals increased variability in information value post-2020, characterized by wider confidence bounds and more pronounced peaks and troughs.
-The heightened volatility coincides with the COVID-19 pandemic period, which introduced substantial market uncertainty.
-The patterns demonstrate that the dynamic nature of analyst information value and its sensitivity to market conditions.
-The Online Appendix provides additional robustness checks.
-Table A7 presents alternative information value estimates using high-frequency realized return volatility measures and varied trade-signing algorithms, yielding estimates ranging from \$0.42 million to \$0.64 million.
-Figure A8 demonstrates that these alternative specifications produce consistent temporal patterns in information value evolution.
+To translate the statistical measure of $R^{2}_{\mathrm{OOS}}$ into economic terms, I follow the framework of Kadan and Manela (2025).
+The framework develops an empirical measure for quantifying the value of insider information base on the theoretical model of Back et al.
+(2000).
+I employ the measure for two primary reasons.
+First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of multiple investors with asymmetric yet overlapping access to analyst information.
+Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
+The economic value of information ($\widehat{\Omega}_{it}$) for stock $i$ on day $t$ is defined as the ratio of the explainable component of return variance to the standardized price impact of trading:
+$\widehat{\Omega}_{it}=\frac{\text{Explainable Return Variance}}{\text{Standardized Price Impact}}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{it}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}}.$ (8)
+Here, $\widehat{r}_{it}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ for stock $i$ on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$ (where $\hat{r}_{it}=\frac{1}{N}\sum_{j=1}^{N}\hat{r}_{ijt}$ if multiple reports exist).
+The numerator thus captures the portion of return variance explained by the analyst reports.
+The denominator consists of the estimated price impact, $\widehat{\lambda}_{it}$, scaled by the stock’s closing price at day $t-2$, $p_{it\_}$.
+This measure provides an estimate of the potential trading profits available to strategic investors who can trade on the information contained in analyst reports.
+By construction, the value of information ($\widehat{\Omega}_{it}$) increases with the model’s explanatory power (a higher numerator) and decreases with market liquidity (a lower price impact, $\widehat{\lambda}_{it}$, in the denominator), linking statistical informativeness with economic relevance.
+To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns $r_{itk}$ on contemporaneous order flow $y_{itk}$:
+$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (9)
+where $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ is the log price change, and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$ denotes signed order flow over interval $[\tau_{k-1},\tau_{k}]$.
+$\widehat{\lambda}_{it}$ (Kyle’s lambda) represents how sensitive the price is to trading volume.
+Buy and sell trades are classified using the algorithm of Lee and Ready (1991).
+Robustness checks also consider methods from Ellis et al.
+(2000) and Chakrabarty et al.
+(2007).
+The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
+Given that $\widehat{\Omega}_{it}$ is a ratio of two random variables measured with error, the interpretation of the information value estimate requires care.
+To address this challenge, I apply the delta method to estimate the mean and variance of the information value.
+The derivation of first-order estimations using the delta method is detailed in the Online Appendix E.
+Table 8 reports the average analyst information value estimates.
+The average economic value of analyst reports for an S&P 100 stock is \$0.47 million over a three-day event window (SE = \$0.05 million).
+This figure represents the potential profits available to strategic investors who can trade on the combined numerical and textual information.
+To understand the sources of this value, I estimate the information value derived from textual and numerical content separately.
+A model based solely on text embeddings yields information value of \$0.38 million, while a model based only on numerical forecast revisions yields \$0.34 million.
+The fact that the combined model’s value (\$0.47M) is less than the sum of the individual components suggests a significant overlap or interaction between these information sources.
+Extrapolating from an average of 15 report days per year, the annualized lower-bound estimate of investor profits from early access to analyst information is \$6.89 million for a typical S&P 100 stock.
+A key input to this calculation is the price impact of trading.
+The average standardized price impact ($\widehat{\lambda}_{it}/p_{it\_}$) is 0.34, which implies that a \$1 million trade moves the stock’s price by approximately 3.4 basis points.
+The modest price impact of these highly liquid, large-cap stocks helps explain the substantial economic value.
+Figure 4 plots the time series of information value estimates from 2015Q1 to 2023Q4.
+The solid line, representing the combined value from text and numbers, exhibits a mild upward trend, suggesting that the economic value of analyst reports has, on average, increased over the past decade.
+The persistent gap between the combined value and the value from text alone confirms the complementary nature of quantitative and qualitative information.
+Notably, the shaded 95% confidence band widens post-2020, signaling greater uncertainty and variation in information value during and after the COVID-19 pandemic.
+The fluctuations demonstrate the dynamic nature of analyst information value and its sensitivity to market conditions.
+Further robustness checks are detailed in the Online Appendix.
+Table A8 shows that using alternative high-frequency volatility measures and different trade-signing algorithms (e.g., Ellis et al., 2000; Chakrabarty et al., 2007) yields consistent estimates, with average information values ranging from \$0.42 million to \$0.64 million.
+Furthermore, Figure A8 demonstrates that the temporal trends in information value are consistent across these alternative specifications.
 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
-Analyst information value may vary in the cross-section of stocks.
-Prior literature suggests that information processing and dissemination costs vary with firm size (Zeghal, 1984), and analyst opinions may trigger larger market reactions for smaller stocks due to their relatively limited public information environment (Stickel, 1995).
+The economic value of analyst information may vary across the cross-section of stocks.
+Prior literature offers competing hypotheses regarding firm size.
+On the one hand, information acquisition and analysis are typically more costly for smaller, less-followed firms, suggesting analysts might focus on larger firms (Zeghal, 1984).
+On the other hand, the relative scarcity of public information for smaller firms could make analyst opinions more influential and thus more valuable to the market (Stickel, 1995).
+This creates an empirical question as to whether information value is positively or negatively related to firm size.
 Figure A9 illustrates this relationship through a scatter plot of information value against stock market capitalization.
 While the plot reveals a positive association between market equity and information value, the relationship exhibits notable nonlinearities.
-To formally examine this relationship, I estimate regressions of information value on stock size, incorporating firm and year fixed effects and the book-to-market ratio control.
-Table 10 shows that stocks with larger market equity have significantly higher information value from analysts.
-The economic magnitude is substantial: estimates from column (4) indicate that a 1% increase in market equity corresponds to a 0.864% increase in analyst information value.
-Columns (5)-(8) reveal that textual information value exhibits greater size sensitivity compared to the combined numerical and textual information value.
-To better understand the mechanisms driving this positive size elasticity, I decompose the information value into two components: explained return volatility and price impact.
-The decomposition can be expressed logarithmically as:
-$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
-Panel B of Table 10 shows that higher analyst information value in larger stocks stems primarily from their greater liquidity.
-While analyst reports do not better explain return volatility in large stocks, the reduced price impact of liquid stocks enhances their information value.
+To formally test this relationship, I estimate regressions of the information value measure on firm size (log market equity), and controlling for the book-to-market ratio, firm, and year fixed effects.
+The results in Table 9 confirm a strong positive relationship.
+The coefficient estimates in Column (4) imply an elasticity of 0.864, meaning a 1% increase in market equity is associated with a 0.864% increase in analyst information value.
+Notably, Columns (5) through (8) show that this elasticity is even stronger for text information value, suggesting that the interpretive role of analysts is particularly important for larger, often more complex firms.
+To better understand the mechanisms behind this relationship, I decompose information value into two components: the explanatory power of the report and the price impact of trading.
+Expressed in logarithmic form, the decomposition is given by:
+$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\hat{r}_{it}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (10)
+This allows for testing whether the higher information value for large firms stems from an information channel (analysts providing better or more impactful information) or a liquidity channel (trading on the information is cheaper due to lower price impact).
+Panel B of Table 9 shows that the higher information value for large-cap stocks arises mainly from lower price impact, not greater explanatory power.
+That is, while analyst reports do not better explain return variance for larger firms, their improved liquidity reduces transaction costs, thereby enhancing the value of analyst information from a trading perspective.
+Panel B of Table 9 presents results of the decomposition analysis.
+The findings are striking: the positive relationship between firm size and information value arises mainly from the liquidity channel.
+While analyst reports do not exhibit significantly greater explanatory power for larger firms, the price impact of trading is substantially lower for these high-liquidity stocks.
+This reduced transaction cost enhances the value of analyst information.
+As my sample is composed of large stocks, it is crucial to recognize that the economic value of analyst information likely decreases for smaller, less liquid firms.
 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
-Prior literature suggests that analysts who make bold forecasts incorporate private information more comprehensively than those who herd (Clement and Tse, 2005).
-This section examines the relationship between information value and analysts’ herding behavior.
-To analyze cross-sectional variation in analyst characteristics, I estimate regressions of stock-analyst-day information value on forecast boldness indicators, incorporating firm, analyst, and year fixed effects.
-Following Jacob et al.
-(1999), who suggests larger brokerage firms provide superior research, I include broker size as a control variable.
+This section explores how the economic value of analyst reports varies with analyst characteristics, focusing on bold and herding forecast behavior.
+Prior research suggests that bold forecasts are more likely to be based on an analyst’s private information, whereas herding behavior often reflects a reliance on public signals or consensus views (Clement and Tse, 2005).
+I therefore test the hypothesis that bold reports generate greater strategic information value.
+To examine this cross-sectional variation, I regress the stock-analyst-day level information value on an indicator for forecast boldness, controlling for brokerage size as well as firm, analyst, and year fixed effects.
 The analyst-specific information value measure is defined as:
-$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
-Table 11 presents the regression results.
-Panel A, columns (1)-(4), reveals significantly positive coefficient estimates on $Bold$, indicating more strategically valuable information when analysts provide bold forecasts.
-Quantitatively, information from bold reports demonstrates 29% higher value compared to herding reports—a magnitude that persists when examining textual content alone.
-The insignificant coefficients for $Brokersize$ suggest that broker size is related to analyst information value.
-To understand the underlying mechanisms, Panel B decomposes the relationship between analyst boldness and information value into informativeness and price impact components.
-As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts issue bold forecasts when possessing relevant private information.
-Columns (5)-(8) reveal no statistically significant relationship between boldness and price impact, suggesting that the enhanced value of bold forecasts stems primarily from their information content rather than market microstructure effects.
+$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}},$ (11)
+where $\hat{r}_{ijt}$ is the out-of-sample CAR estimation for report $j$ of stock $i$ on day $t$.
+The regression results are presented in Table 10.
+Panel A shows that the coefficient on the boldness indicator (Bold) is positive and highly significant across all specifications.
+The estimated magnitude implies that, on average, bold analyst reports are 29% more valuable than their herding counterparts.
+This pattern holds when analyzing the value derived from textual content alone.
+In contrast, the coefficient on brokerage size is consistently insignificant, suggesting that after controlling for other factors, the size of brokerage firms does not materially influence the economic value of analysts’ reports.
+To understand the channel through which boldness creates value, Panel B decomposes the information value into its two components: explained return variance (the information channel) and price impact (the liquidity channel).
+The results clearly indicate that the effect is driven by the information channel.
+Columns (1)–(4) show that bold reports are associated with significantly larger explained return variance, reinforcing the view that they contain meaningful signals.
+Conversely, Columns (5)–(8) reveal no systematic relationship between boldness and price impact, going against the possibility that the higher value of bold reports stems from differential market liquidity or trading frictions.
 3.3.3 The Value of Analyst Information around Earnings Announcements
-Previous evidence underlines the information content of analyst reports in the first week following earnings announcements.
-Accordingly, this section evaluates the information value of analyst reports issued promptly following earnings announcements and beyond that.
-Figure 5 examines different samples of analyst reports categorized by their timing relative to earnings announcements, from reports released in the first week after earnings announcements to those issued in the $13^{th}$ week.
-Reports issued in the first post-earnings week exhibit the highest information value, with a sample mean of \$0.84 million.
-Reports from subsequent weeks show a sharp decline in value.
-This pattern again suggests that markets place premium value on timely post-earnings analyst reports.
-To formally test the observed pattern, I estimate regressions of log information value on an indicator variable $Week$, which identifies reports released within one week after earnings announcements.
-Following Banerjee (2011) and Beckmann et al.
-(2024), I include earnings announcement day trading volume ($TradingVolume$) as a proxy for investor disagreement.
-The model incorporates an interaction term ($Week\times TradingVolume$) to capture the joint effect of post-earnings timing and investor disagreement on analyst information value.
-The specification builds on Chen et al.
-(2010)’s finding that analysts’ interpretative role becomes more critical with increasing financial information complexity.
-The model includes stock and year fixed effects.
-Table 12 presents regression results.
+Building on earlier findings that the statistical information content of analyst reports peaks shortly after earnings announcements, this section examines whether the economic value of these reports follows a similar pattern.
+I investigate how the value of analyst information varies based on its timeliness relative to corporate earnings disclosures.
+Figure 5 plots the average information value of analyst report subsamples across 13 weekly bins following earnings announcements.
+Reports issued in the first week show the highest average value, at approximately \$0.84 million.
+This value declines sharply in subsequent weeks, suggesting that the market places a significant premium on timely analyst reports that digest and interpret corporate news.
+To formally test this observation, I estimate panel regressions of log information value on a post-earnings week indicator ($Week$), which equals one if the report is issued within one week of the earnings announcement.
+Following prior literature that suggests analyst interpretation is most valuable amidst uncertainty (Chen et al., 2010), I also test if this effect is amplified during periods of high investor disagreement.
+I proxy for disagreement using trading volume on the earnings announcement day ($TradingVolume$) and include an interaction term ($Week\times TradingVolume$), following Banerjee (2011) and Beckmann et al.
+(2024).
+All specifications include stock and year fixed effects.
+The regression results are presented in Table 11.
 Panel A documents significantly higher information value for reports issued within the first post-earnings week, as evidenced by positive and significant coefficients on $Week$ across all models.
-The positive and significant interactions between $Week$ and $TradingVolume$ indicate that information value increases with trading volume on earnings announcement days.
-The results are consistent with investors placing greater reliance on analyst interpretations during periods of heightened disagreement.
-Panel B decomposes the earnings announcement effects into return variance and price impact, showing significant increases in both components.
-The coefficient magnitudes indicate that the timing effect is primarily driven by higher explained return volatility, rather than changes in price impact.
-These findings lead to conclude that investors value the most of analysts’ interpretation role during periods of heightened post-earnings uncertainty.
+Furthermore, the positive and significant coefficient on the interaction term indicates that this timeliness premium is even greater when there is high investor disagreement surrounding earnings announcements.
+This is consistent with investors placing the greatest reliance on analyst sense-making during periods of heightened uncertainty.
+Panel B decomposes this finding to identify the channel through which timeliness creates value.
+By regressing the two components of the information value measure on the same set of variables, I can distinguish between an information channel and a liquidity channel.
+The results show that the timeliness effect is driven primarily by the information channel: reports issued shortly in the post-earnings announcement window are associated with a significant increase in explained return variance, while changes in price impact are relatively modest.
 4 Conclusion
-This paper examines both the quantitative and qualitative content of analyst reports.
-I find that report text contains more market-valued information content than quantifiable forecast revisions economically, especially shortly following earnings announcements.
-A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable component of written reports.
-I further translate the information content into economic terms.
-Using a measure of explained return volatility relative to price impact, I find that early acquisition of the analyst reports for an average S&P 100 constituent firm yields an annualized profit larger than \$6.89 million.
-Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
-These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets.
-The substantial potential profits from early access to analyst reports raise concerns about selective distribution practices—a direct violation of FINRA Rule 2241, which mandates research objectivity and prohibits preferential distribution to certain clients.
-This study presents a hypothetical scenario for quantifying information value rather than an assessment of market practices.
-By documenting the magnitude of potential profits from privileged access, my research emphasizes the importance of regulations for fair disclosure.
+This paper examines both the quantitative and qualitative information content of analyst reports.
+I find that the textual content of analyst reports carries more economically significant, market-valued information than quantifiable forecast revisions.
+This effect is especially pronounced in the period immediately following earnings announcements.
+A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable topic within written reports.
+To translate these findings into economic terms, I adopt a measure of information value accounting for both the explained return variance and the price impact.
+The results suggest that early acquisition of analyst reports for an average S&P 100 constituent stock yields annualized profits exceeding \$6.89 million.
+The value of analyst information is higher for large-cap firms, bold analyst forecasts, and when investors’ disagreement is high.
+These findings have important implications for regulators.
+By quantifying the substantial economic value of privileged access to analyst reports, I reveal strong incentives for selective distribution, a practice prohibited by FINRA Rule 2241.
+It is important to note that the analysis provides a hypothetical framework for understanding the potential value of such information, not an empirical assessment of market practices.
+Nevertheless, the magnitude of these potential profits reinforces the fundamental principle that equal access to value-relevant information is essential for market integrity.
 A. Additional Figures
+Figure A1: Large Language Model Architecture
+This figure demonstrates how a large language model generates context-aware representations using a sample sentence from an analyst report on NVIDIA.
+The model first tokenizes the sentence and generates two types of embeddings: token embeddings and positional embeddings.
+These are then passed through a stack of transformer layers, the core of the architecture.
+The output of each transformer layer is a sequence of context-aware vectors, one for each token, capturing its meaning concerning the full context.
 (a) (1) Executive Summary
@@
 (d) (4) Competitive Landscape
-Figure A1: Word Clouds of Topics
-This figure presents word clouds for 16 key topics frequently discussed in analyst reports.
-Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency.
+Figure A2: Word Clouds of Analyst Report Topics
+This figure displays word clouds for 16 key topics commonly discussed in analyst reports.
+Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency.
 The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures.
@@
 (d) (8) Financial Ratios
-Figure A1: Word Clouds of Topics
-This figure presents word clouds for 16 key topics frequently discussed in analyst reports.
-Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency.
+Figure A2: Word Clouds of Analyst Report Topics
+This figure displays word clouds for 16 key topics commonly discussed in analyst reports.
+Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency.
 The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures.
@@
 (d) (12) Management and Governance
-Figure A1: Word Clouds of Topics
-This figure presents word clouds for 16 key topics frequently discussed in analyst reports.
-Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency.
+Figure A2: Word Clouds of Analyst Report Topics
+This figure displays word clouds for 16 key topics commonly discussed in analyst reports.
+Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency.
 The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures.
@@
 (d) (16) Appendices and Disclosures
-Figure A1: Word Clouds of Topics
-This figure presents word clouds for 16 key topics frequently discussed in analyst reports.
-Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency.
+Figure A2: Word Clouds of Analyst Report Topics
+This figure displays word clouds for 16 key topics commonly discussed in analyst reports.
+Each word cloud highlights the most frequently occurring terms within each topic, with word size proportional to term frequency.
 The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures.
 The “None of the Above” category is excluded from the visualization.
-Figure A2: Large Language Model Architecture
-This figure shows an example of how large language models process a sentence from an analyst report on Nvidia.
-The model breaks this sentence into tokens, and then generates two types of embeddings: token embeddings and position embeddings.
-These embeddings then pass through the core of the model - a stack of transformer layers.
-The output of each transformer layer is a series of vectors representing the tokens in their full context.
 Figure A3: Analyst Information Value by Industry
-This figure shows the estimated information value of analyst reports for Fama-French 12 industries from 2015Q1 to 2023Q4, quantified in millions of dollars.
-The solid green line marks the sample mean value of information, with green bars showing the 95% confidence intervals.
-The horizontal dashed red line marks the sample mean of stocks from all industries.
-Information value estimates are derived using the delta method and adjusted to 2020 levels using the Consumer Price Index (CPI).
-Figure A4: Shapley Values Scaled by Length
-This figure presents the SHAP values of topics scaled by the total number of sentences (Panel A) and the total number of tokens (Panel B) for topics with more than 100,000 sentences.
-The scaling process involves two steps: First, normalize the Shapley value of each topic by dividing it by the number of sentences in that topic.
-Then, adjust the normalized values to ensure the total sum of SHAP values remains consistent.
-This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
-Figure A5: Shapley Values for Topic Importance across Industries
-This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value.
-For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings.
-The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
-Figure A6: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
-This figure shows the $R^{2}_{\mathrm{OOS}}$ of sentence-segmented report embeddings from 2015 to 2023.
-The color gradient indicates the magnitude of $R^{2}_{\mathrm{OOS}}$.
-The red line represents the SHAP value of the Income Statements Analyses topic across the years.
-The shaded area marks the pandemic recession in 2020.
-Figure A7: Shapley Values for Topic Importance by Analyst Characteristics
-This figure illustrates how 17 different topics contribute to the predictive power of analyst reports across four subsamples, measured using SHAP values.
-The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), and large (Panel C) versus small brokerage firms (Panel D), with subsamples determined by yearly median.
-For each subsample, the bars show the SHAP values of individual topics, whose sum equals the out-of-sample R-squared ($R^{2}_{\mathrm{OOS}}$) from the sentence-segmented report embeddings.
-The red line indicates each topic’s relative importance, calculated as the topic’s SHAP value divided by the total $R^{2}_{\mathrm{OOS}}$.
-Figure A8: Alternative Analyst Information Value Estimations Over Time
-This figure presents the alternative estimations of analyst information value from 2015Q1 to 2024Q4, reported in millions of dollars and adjusted to 2020 dollars.
-Quarterly mean approximations are calculated using the delta method.
-TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns.
-EMO and CLNV refer to trade-signing algorithms developed by Ellis et al.
+This figure shows the estimated information value of analyst reports across Fama-French 12 industries from 2015Q1 to 2023Q4, with values reported in millions of 2020 dollars and estimated via the delta method.
+The solid green line indicates the mean information value for each industry, while the green bars represent 95% confidence intervals.
+The horizontal dashed red line denotes the cross-industry average.
+Figure A4: Shapley Values Scaled by Topic Length
+This figure adjusts the topic Shapley values to measure their per-sentence or per-token contribution rather than their aggregate importance, focusing on topics with over 100,000 sentences.
+Panel A presents the values scaled by the number of sentences in each topic, while Panel B presents them scaled by the number of tokens.
+This two-step scaling process involves first normalizing each topic’s total Shapley value by its length (sentences or tokens) and then rescaling the results to maintain the original sum of all values, thereby preserving the model’s total explanatory power.
+Figure A5: Topic Importance via Shapley Value across Industries
+This figure decomposes the information content ($R^{2}_{\mathrm{OOS}}$) of analyst reports into contributions from 17 topics for each of the Fama-French 12 industries.
+Within each industry’s panel, the height of a bar represents the Shapley value for a specific topic, indicating its contribution to the total out-of-sample $R^{2}$ for reports in that sector.
+The red line shows each topic’s SHAP value as a proportion of the aggregated $R^{2}_{\mathrm{OOS}}$.
+Figure A6: Information Content of Sentence-Segmented Embeddings by Year
+This figure shows the evolution of the total out-of-sample explanatory power ($R^{2}_{\mathrm{OOS}}$) of sentence-segmented analyst reports from 2015 to 2023.
+Specifically, the red line isolates the contribution of the ‘Income Statement Analyses’ topic, as measured by its annual Shapley value.
+The shaded area denotes the 2020 pandemic recession.
+Figure A7: Topic Importance via Shapley Value by Analyst Characteristics
+This figure illustrates how 17 different topics contribute to the analyst report information content across six subsamples.
+The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), large (Panel C) versus small brokerage firms (Panel D), and the primary valuation methodology used (Panel E: Multiple-Based vs.
+Panel F: DCF-Based).
+The splits for experience and brokerage size are based on the yearly median.
+For each group, bars represent the SHAP values of individual topics, with the red line indicating each topic’s relative importance scaled by total $R^{2}_{\mathrm{OOS}}$.
+Figure A8: Alternative Estimations of Analyst Information Value Over Time
+This figure presents several alternative estimations of the quarterly analyst information value from 2015Q1 to 2024Q4, serving as a robustness check to the primary measure.
+These alternatives modify key components of the main calculation.
+The ‘TAQ volatility’ series uses realized volatility from one-minute log returns ($\sigma_{v}^{2}$) to calculate explained return volatility as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$.
+The ‘EMO’ and ‘CLNV’ series use alternative trade-signing algorithms for the price impact calculation developed by Ellis et al.
 (2000) and Lee and Ready (1991), respectively.
-Figure A9: Analyst Information Value by Stock
-This figure shows the estimated information value of analyst reports for individual stocks from 2015Q1 to 2023Q4, quantified in millions of dollars.
-Estimates are derived using the delta method.
-Market equity, averaged daily at the time of each report release, is reported in billions of dollars.
-All values are adjusted to 2020 levels using the Consumer Price Index (CPI).
+All estimates are reported in millions of 2020 dollars, with quarterly means approximated via the delta method.
+Figure A9: Analyst Information Value and Stock Market Equity
+This figure plots the estimated information value of analyst reports against the market equity for each individual stock, covering the period from 2015Q1 to 2023Q4.
+The information value (vertical axis) for each stock is estimated using the delta method and reported in millions of 2020 dollars, adjusted using the CPI.
+The market equity (horizontal axis) is the firm’s average daily market capitalization at the time of each report release, reported in billions of dollars.
 B. Additional Tables
@@
 Numerical Measures Definition and/or sources
-Panel A: Firm Level Measures
+Panel A: Firm-Level Measures
 Size The market value equity of the firm ($CSHOQ*PRCCQ$) at the end of the quarter prior to report release.
-BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of quarter prior the report release.
-Prior_CAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
+BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of the quarter prior to the report release.
+PriorCAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
 SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement.
-Consensus EPS is the median value of 1-year EPS forecast within 90-days window of all analysts following the firm.
+Consensus EPS is the median value of 1-year EPS forecast within a 90-day window of all analysts following the firm.
 The unexpected earnings is scaled by price per share at the fiscal quarter end.
@@
 TradingVolume Trading volume at earnings announcement day (or the first trading day post earnings announcement), calculated as $VOL/SHROUT$.
-Distance to default The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
+DD The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
 Fluidity A measure of firms’ product market competition introduced by Hoberg et al. (2014). The data is compiled from the Hoberg and Phillips database.
-Panel B: Industry Level Measures
+Panel B: Industry-Level Measures
 IndustryRecession An indicator variable that equals one if the FF-48 industry return is negative and in the bottom quintile of FF-48 industry returns and zero otherwise.
 Panel C: Macroeconomic Measures
-Time trend (ttr) The number of years elapsed from the beginning of the sample.
-Panel D: Report Level Measures
-$REC\_REV$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
-$EF\_REV$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
-$TP\_REV$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
+TimeTrend The number of years elapsed from the beginning of the sample.
+Panel D: Report-Level Measures
+$REC_{REV}$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
+$EF_{REV}$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
+$TP_{REV}$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
 Boldness An indicator variable that takes the value of 1 if the EPS forecast revision is above both the analyst’s own prior forecast and the consensus forecast, or else below both, and zero otherwise.
@@
 Table A3: Sumamry Statistics for Numerical Measures
-This table reports summary statistics of numerical measures.
-Mean Std p25 P50 P75 N
+This table reports summary statistics of numerical measures. See Table A2 for detailed variable definitions.
+Mean Std. Dev. p25 P50 P75 N
 Miss 0.35 0.48 0.00 0.00 1.00 102776
 TradingVolume 18.19 21.98 7.24 11.46 20.29 99908
-Prior_CAR 0.00 0.05 -0.02 0.00 0.02 122251
+PriorCAR 0.00 0.05 -0.02 0.00 0.02 122251
 Size 2.39 0.11 2.33 2.39 2.46 117534
 BtoM 0.46 0.43 0.18 0.32 0.61 117534
-Distance to default 0.00 0.01 0.00 0.00 0.00 86693
+DD 0.00 0.01 0.00 0.00 0.00 86693
 Fluidity 7.17 3.66 4.40 6.57 9.44 114942
@@
 Table A4: Information Content of Analyst Reports with Numbers Removed
-This table presents the $R^{2}_{\mathrm{OOS}}$ using analyst reports with numbers removed.
-The input sets include revision numerical measures, text embeddings from number-free report text, and a combination of both.
-The estimation targets are the three-day $CAR$ centered around the analyst report release date.
-The machine learning model employed is Ridge regression, with training samples expanding yearly.
-The $t$-statistic for the $R^{2}_{\mathrm{OOS}}$ is calculated following Gu et al.
+This table presents an analysis of analyst reports where all numerical values have been removed from the text, designed to isolate the contribution of pure qualitative information.
+I report the $R^{2}_{\mathrm{OOS}}$ from Ridge regressions predicting three-day cumulative abnormal returns ($CAR[-1,+1]$) using several input configurations: numerical forecast revisions only, embeddings from the number-free text, and a combination of both.
+The table also includes pairwise tests to assess whether the number-free text contains more information than the numerical revisions alone (Column (3)-(1)), and whether those revisions provide incremental value when added to the number-free text (Column (5)-(3)).
+T-statistics for the $R^{2}_{\mathrm{OOS}}$ and their differences are calculated following the procedure outlined by Gu et al.
 (2020).
-In columns (2), (4), and (6), the benchmark estimation is set to zero.
-Columns (3)-(1) and (5)-(3) present pairwise tests comparing the performance of models using revision-only input versus text-only input, and text-only input versus combined text and revision input, respectively.
-Year Revision only t-stat Text only t-stat Revision plus text t-stat t-stat t-stat
+Year Rev only t-stat Text only t-stat Rev + text t-stat t-stat t-stat
 (1) (2) (3) (4) (5) (6) (3)- (1) (5)- (3)
@@
 Overall 8.93% 8.51 10.95% 7.87 12.51% 8.83 2.40 3.00
+Table A5: Information Content of Analyst Reports using Alternative Machine Learning Models
+This table compares the performance ($R^{2}_{\mathrm{OOS}}$) of several alternative machine learning models in estimating the information content of analyst report text.
+The models tested include Partial Least Squares (PLS), Extreme Gradient Boosting (XGBoost), and a series of Neural Networks (NN1-NN5) with an increasing number of hidden layers.
+The out-of-sample $R^{2}$ for each model is calculated annually using an expanding training window that begins in 2000.
+The ‘Overall’ row summarizes the performance across the full 2015–2023 sample period.
+T-statistics for the $R^{2}_{\mathrm{OOS}}$ are calculated against a zero benchmark following the procedure in Gu et al.
+(2020).
+PLS XGBoost NN1 NN2 NN3 NN4 NN5
+year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
+2015 12.23% 10.95 8.87% 11.25 15.81% 9.72 13.84% 14.53 15.30% 14.10 11.58% 12.32 16.61% 16.95
+2016 10.27% 15.59 6.66% 5.63 12.59% 24.64 11.78% 13.02 10.99% 9.10 12.03% 14.75 12.46% 17.41
+2017 6.60% 6.94 4.92% 5.26 9.08% 4.64 10.01% 4.64 9.58% 4.76 8.57% 4.51 10.03% 5.49
+2018 8.08% 13.37 5.15% 8.07 10.73% 26.65 10.54% 16.42 10.41% 11.43 9.86% 17.44 11.98% 13.97
+2019 11.54% 38.70 6.32% 16.40 13.43% 17.37 15.18% 14.73 15.59% 18.06 15.89% 10.48 15.37% 16.94
+2020 3.21% 2.16 3.02% 3.64 4.45% 2.92 4.73% 2.78 3.85% 2.27 4.25% 3.08 5.51% 4.06
+2021 2.14% 1.76 4.30% 3.34 4.07% 6.44 5.88% 7.13 9.28% 13.59 8.47% 9.44 8.83% 7.67
+2022 11.42% 7.40 10.04% 7.80 13.78% 8.31 15.67% 6.38 16.44% 6.64 17.63% 8.55 17.56% 7.12
+2023 8.14% 9.26 4.23% 3.82 9.54% 8.50 12.16% 10.76 11.06% 8.59 12.18% 11.59 11.70% 6.78
+Overall 8.26% 6.66 5.97% 6.76 10.42% 8.11 11.04% 7.54 11.18% 7.29 11.05% 6.62 12.12% 7.93
 Table A6: Cumulative Abnormal Returns with Alternative Windows
-This table reports the $R^{2}_{\mathrm{OOS}}$ of using analyst report text embeddings to estimate $CAR$ with alternative windows.
-The $R^{2}_{\mathrm{OOS}}$ is calculated annually using a training sample from 2000 to the preceding year.
-$CAR$ measures are centered on the release day of analyst reports ($T_{0}$).
-Panel A shows results for the full sample, while Panels B and C focus on periods within and beyond 1-day earnings announcement windows, respectively.
+This table reports the $R^{2}_{\mathrm{OOS}}$ from Ridge regression models that use analyst report text embeddings to estimate CAR across alternative event windows.
+The target variables are abnormal returns centered on the report release date ($T_{0}$): same-day abnormal return (AR[0]), two-day cumulative return (CAR[0,+1]), and five-day cumulative return (CAR[–2,+2]).
+All $R^{2}_{\mathrm{OOS}}$ values are computed annually using an expanding training sample starting in 2000.
+Panel A shows results for the full sample, while Panels B and C restrict the sample to reports released within one day (excluding same day) of an earnings announcement and those outside this window, respectively.
 The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023.
@@
 year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
-Panel A: Full sample period
+Panel A: Full sample
 2015 8.03% 8.53 8.24% 8.36 10.87% 9.40
@@
 Overall 7.55% 8.07 6.92% 6.77 7.57% 6.79
-Panel B: Earnings announcement period
+Panel B: Reports issued within earnings announcement window
 Overall 9.78% 5.58 7.87% 5.10 13.14% 8.02
-Panel C: Non-earnings announcement period
+Panel C: Reports issued beyond earnings announcement window
 Overall 5.77% 5.78 5.62% 5.64 4.96% 4.91
 Table A7: Information Content of Analyst Reports across Industries
-This table reports the $R^{2}_{\mathrm{OOS}}$ of Fama-French 12 industries, using the definition from Kenneth French’s website.
-The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023.
-The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al.
+This table reports information content of analyst report text for each of the Fama-French 12 (FF12) industries for the period of 2015–2023.
+Industry definitions are based on Kenneth French’s data library.
+T-statistics for all $R^{2}_{\mathrm{OOS}}$ values are calculated against a zero benchmark following the procedure in Gu et al.
 (2020).
-year Shops Other Manuf Chems
-$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
+Shops Other Manuf Chems
+year $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
 2015 13.27% 6.03 657 17.69% 5.03 585 10.84% 3.19 808 15.05% 1.74 147
@@
 Overall 13.88% 6.03 4003 11.27% 3.69 4374 11.39% 5.57 5151 11.35% 3.55 926
-year Durbl BusEq Hlth NoDur
-$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
+Durbl BusEq Hlth NoDur
+year $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
 2015 16.83% 5.77 216 15.66% 6.41 1378 5.75% 2.72 1242 0.68% 0.31 241
@@
 Overall 11.02% 3.24 1757 9.68% 6.77 9488 7.53% 5.71 8525 7.50% 1.38 2167
-year Money Telcm Enrgy Utils
-$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
+Money Telcm Enrgy Utils
+year $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
 2015 18.13% 4.41 1153 1.68% 0.00 403 -4.21% -1.45 303 18.47% 3.69 401
@@
 Overall 5.18% 3.31 7990 3.32% 1.39 2482 1.90% 0.31 2738 -1.97% -0.88 1954
-Table A7: Summary Statistics for Alternative Analyst Information Value Estimations
-The analyst information value is measured as the explained return variance divided by the price impact, with the results reported in millions of dollars.
-All dollar values are adjusted for inflation to reflect 2020 values using the Consumer Price Index (CPI).
-The sample comprises analyst reports targeting common constituent firms of the S&P 100 index from 2015Q1 to 2023Q4.
-Both mean and standard deviation are computed using the delta method.TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns.
-Intraday denotes that $\sigma_{v}^{2}$ does not include overnight return volatility.
-EMO and CLNV refer to trade-signing algorithms developed by Ellis et al.
-(2000) and Lee and Ready (1991), respectively.
+Table A8: Summary Statistics of Alternative Analyst Information Value Estimations
+This table reports summary statistics for several alternative estimations of analyst information value for S&P 100 stocks from 2015Q1 to 2023Q4.
+These alternatives modify different components of the main information value calculation in Table 8 (defined as explained return variance divided by price impact).
+The ‘TAQ volatility’ specification calculates explained variance as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$ using one-minute log returns, while the ‘Intraday’ version excludes overnight returns from this calculation.
+The ‘EMO’ and ‘CLNV’ specifications use alternative trade-signing algorithms from Ellis et al.
+(2000) and Lee and Ready (1991) to calculate price impact.
+For each specification, the table reports the mean, standard error (SE), and 95% and 99% confidence intervals, with all estimates computed with the delta method and expressed in millions of 2020 dollars.
 Mean SE 95%CI 99%CI N
@@
 I extend the model by considering a case where $\phi$ percentage of variance in $\tilde{v}$ is explained by the informed trader’s information.
-A Single Auction Model
+Single Auction Model
 There are two periods, $t_{0}$ and $t_{1}$.
@@
 $X(\tilde{s})=\alpha+\beta\tilde{s}.$
-The profit of the insider can be written as:
+The profit of the insider can be written as
 $E[\tilde{\pi}(X,P)\mid\tilde{s}=s]$ $=E[(\tilde{v}-\mu-\lambda(\tilde{u}+x))x\mid\tilde{s}=s]$
@@
 Traders take into account that her order flow will move the price against her, which serves to restrain her position size.
-Solving for optimal profit, I get:
+Solving for optimal profit, I get
 $x^{*}=\frac{P_{0}+s-\mu}{2\lambda}=\alpha+\beta v.$
-Hence, I can express $\alpha$ and $\beta$ :
+Hence, I can express $\alpha$ and $\beta$ as
 $\begin{gathered}\beta=\frac{1}{2\lambda},\\ \alpha=\frac{P_{0}-\mu}{2\lambda}=(P_{0}-\mu)\beta.\end{gathered}$
 When the market maker puts a higher weight on the order flow in setting the price, the trader puts a lower weight on her information.
-I now look at the price-setting rule:
+I now look at the price-setting rule as
 $\mu+\lambda y=E\{\tilde{v}\mid\alpha+\beta\tilde{s}+\tilde{u}=y\}.$
@@
 $=p_{0}+\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}}(y-\alpha).$
-Hence, I can express $\lambda$ and $\mu$ :
+Hence, I can express $\lambda$ and $\mu$ as
 $\lambda=\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}},$
@@
 $g(X,Y)\approx g(\mu_{X},\mu_{Y})+\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}(X-\mu_{X})+\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}(Y-\mu_{Y}).$
-The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are:
+The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are
 $\frac{\partial g}{\partial X}=\frac{1}{Y},\quad\frac{\partial g}{\partial Y}=-\frac{X}{Y^{2}}.$
-Evaluating these at $(\mu_{X},\mu_{Y})$, I get:
+Evaluating these at $(\mu_{X},\mu_{Y})$, I get
 $\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}=\frac{1}{\mu_{Y}},\quad\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}=-\frac{\mu_{X}}{\mu_{Y}^{2}}.$
-Substituting the partial derivatives into the Taylor expansion, I obtain:
+Substituting the partial derivatives into the Taylor expansion, I obtain
 $\frac{X}{Y}\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}(X-\mu_{X})-\frac{\mu_{X}}{\mu_{Y}^{2}}(Y-\mu_{Y}).$
-Taking the expectation on both sides:
+Taking expectations on both sides gives
 $E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}E[X-\mu_{X}]-\frac{\mu_{X}}{\mu_{Y}^{2}}E[Y-\mu_{Y}].$
-Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation simplifies to:
+Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation can be simplified to
 $E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}.$
-Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by:
+Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by
 $\operatorname{Var}(Z)\approx\left(\frac{\partial g}{\partial X}\right)^{2}\operatorname{Var}(X)+\left(\frac{\partial g}{\partial Y}\right)^{2}\operatorname{Var}(Y)+2\left(\frac{\partial g}{\partial X}\right)\left(\frac{\partial g}{\partial Y}\right)\operatorname{Cov}(X,Y).$
-Substituting the partial derivatives:
+Substituting the partial derivatives yields
 $\operatorname{Var}\left(\frac{X}{Y}\right)\approx\frac{\operatorname{Var}(X)}{\mu_{Y}^{2}}+\frac{\mu_{X}^{2}\operatorname{Var}(Y)}{\mu_{Y}^{4}}-2\frac{\mu_{X}\operatorname{Cov}(X,Y)}{\mu_{Y}^{3}}.$
@@
 Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$.
-Define the mean and variance over subsample $s$ as:
+Define the mean and variance over subsample $s$ as
 $\mu_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right),$
-and the mean price impact per dollar over subsample $s$ as:
+and the mean price impact per dollar over subsample $s$ as
 $\mu_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}.$
-Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by:
+Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by
 $E\widehat{\Omega}_{s}\approx\frac{\mu_{\nu s}}{\mu_{\lambda s}}.$
-The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as:
+The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as
 $\text{Var}\left(\widehat{\Omega}_{s}\right)\approx\frac{1}{\mu_{\lambda s}^{2}}\left(\Sigma_{\nu s}+\frac{\mu_{\nu s}^{2}}{\mu_{\lambda s}^{2}}\Sigma_{\lambda s}-2\frac{\mu_{\nu s}}{\mu_{\lambda s}}\Sigma_{\nu\lambda s}\right),$
-where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as:
+where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as
 $\Sigma_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(\widehat{\sigma}_{it}^{2}-\mu_{\nu s}\right)^{2},$